Jefferson Lab HALL A Logbook for October 10 - Full Listing

  • Main INDEX

    Sunday

  • 337894 10/31/10 23:23 a-onl Start_Run_8188,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337893 10/31/10 23:21 a-onl End_of_Run_8187
  • 337892 10/31/10 23:19 A. Tobias Other ROCs for run 8187
  • 337891 10/31/10 23:18 A. Tobias RunControl and Scalars for run 8187
  • 337890 10/31/10 23:17 A. Camsonne Moved clear of VME ADCs in DVCS crates to prestart
  • 337889 10/31/10 23:17 A. Tobias ROCs for run 8187
  • 337888 10/31/10 22:39 a-onl Start_Run_8187,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337887 10/31/10 22:37 a-onl End_of_Run_8186
  • 337886 10/31/10 22:34 A. Tobias Other ROCs for run 8186
  • 337885 10/31/10 22:33 A. Tobias RunControl and Scalars for run 8186
  • 337884 10/31/10 22:32 A. Tobias ROCs for run 8186
  • 337883 10/31/10 22:02 Abdurahim Target screen & Strip charts
  • 337882 10/31/10 21:55 a-onl Start_Run_8186,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LED Clock ON: Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337881 10/31/10 21:53 Abdurahim Compton Vertical Scan
  • 337880 10/31/10 21:53 A. Tobias RuncControl for Run 8185
  • 337879 10/31/10 21:53 a-onl End_of_Run_8185
  • 337878 10/31/10 21:53 A. Camsonne Clock with a large prescale for data taking
  • 337877 10/31/10 21:53 A. Tobias ROCs and Scalars for run 8185
  • 337876 10/31/10 21:39 a-onl Start_Run_8185,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LED Clock ON: Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337875 10/31/10 21:35 A. Camsonne Random trigger on for DVCS at about 11 Hz
  • 337874 10/31/10 21:29 Ramesh DVCS calo anode current
  • 337873 10/31/10 21:28 a-onl End_of_Run_8184
  • 337872 10/31/10 21:25 A. Tobias Other ROCs for run 8184
  • 337871 10/31/10 21:25 A. Tobias RunControl and Scalars for run 8184
  • 337870 10/31/10 21:23 A. Tobias RunControl and Scalars for run 8184
  • 337869 10/31/10 21:22 A. Tobias ROCs for run 8184
  • 337868 10/31/10 20:42 a-onl Start_Run_8184,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337867 10/31/10 20:40 a-onl End_of_Run_8183
  • 337866 10/31/10 20:38 A. Tobias Other ROCs for run 8183
  • 337865 10/31/10 20:36 A. Tobias RunControl and Scalars for run 8183
  • 337864 10/31/10 20:35 A. Tobias ROCs for run 8183
  • 337863 10/31/10 19:54 auto_snap Radiation and Ozone snapshot
  • 337862 10/31/10 19:54 auto_snap Miscellaneous snapshot
  • 337861 10/31/10 19:54 auto_snap Hall A Tools snapshot
  • 337860 10/31/10 19:54 auto_snap Compton Polarimeter snapshot
  • 337859 10/31/10 19:53 auto_snap Beam Position and Current snapshot
  • 337858 10/31/10 19:51 a-onl Start_Run_8183,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337857 10/31/10 19:50 a-onl End_of_Run_8181
  • 337856 10/31/10 19:46 A. Tobias Other ROCs for 8182
  • 337855 10/31/10 19:45 A. Tobias RunControl and Scalars for run 8182
  • 337854 10/31/10 19:44 A. Tobias ROCs for run 8182
  • 337853 10/31/10 19:12 Abdurahim Target screen & Strip charts
  • 337852 10/31/10 19:04 a-onl Start_Run_8182,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337851 10/31/10 19:02 a-onl End_of_Run_8180
  • 337850 10/31/10 19:01 A. Tobias Other ROCs for run 8181
  • 337849 10/31/10 19:00 A. Tobias RunControl & Scalars for run 8181
  • 337848 10/31/10 18:59 auto_snap Radiation and Ozone snapshot
  • 337847 10/31/10 18:59 A. Tobias ROCs for run 8181
  • 337846 10/31/10 18:59 auto_snap Miscellaneous snapshot
  • 337845 10/31/10 18:59 auto_snap Hall A Tools snapshot
  • 337844 10/31/10 18:59 auto_snap Compton Polarimeter snapshot
  • 337843 10/31/10 18:59 auto_snap Beam Position and Current snapshot
  • 337842 10/31/10 18:17 a-onl Start_Run_8181,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337841 10/31/10 18:07 A. Tobias Other ROCs for run 8180
  • 337840 10/31/10 18:06 A. Tobias RunControl & Scalars for run 8180
  • 337839 10/31/10 18:05 A. Tobias ROCs for run 8180
  • 337838 10/31/10 17:23 a-onl Start_Run_8180,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337837 10/31/10 17:23 Abdurahim Compton cavity lock script has been changed
  • 337836 10/31/10 17:10 Abdurahim Target -> LH2, Loop2, 15 cm
  • 337835 10/31/10 16:56 A. Tobias RunControl for run 8179
  • 337834 10/31/10 16:54 a-onl End_of_Run_8179
  • 337833 10/31/10 16:53 A. Tobias ROCs for run 8179
  • 337832 10/31/10 16:50 Ramesh Target moved to empty one
  • 337831 10/31/10 16:46 A. Tobias TS11 still has ERROR in predicted helicity. You are losing data !!
  • 337830 10/31/10 16:34 a-onl Start_Run_8179,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337829 10/31/10 16:22 a-onl End_of_Run_8176
  • 337828 10/31/10 16:21 a-onl End_of_Run_8175
  • 337827 10/31/10 16:21 A. Tobias Other ROCs for run 8176
  • 337826 10/31/10 16:19 A. Tobias RunControl & Scalars for Run 8176
  • 337825 10/31/10 16:18 A. Tobias ROCs for run 8176
  • 337824 10/31/10 16:17 A. Tobias TS11 Errors run 8176
  • 337823 10/31/10 16:12 a-onl Start_Run_8176,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337822 10/31/10 16:10 a-onl End_of_Run_8172
  • 337821 10/31/10 16:08 A. Tobias Other ROCs for run 8175
  • 337820 10/31/10 16:07 A. Tobias RunControl & Scalars for Run 8175
  • 337819 10/31/10 16:06 A. Tobias ROCs for run 8175
  • 337818 10/31/10 16:01 adaq Beam Time Accounting
  • 337817 10/31/10 15:59 bogdanw shift summary
  • 337816 10/31/10 15:21 a-onl Start_Run_8175,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337815 10/31/10 15:20 a-onl End_of_Run_8174
  • 337814 10/31/10 14:27 a-onl Start_Run_8174,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337813 10/31/10 14:26 a-onl End_of_Run_8173
  • 337812 10/31/10 14:19 chench ROCs for run8173
  • 337811 10/31/10 14:18 chench DAQ and scalers for run8173
  • 337810 10/31/10 13:39 a-onl Start_Run_8173,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337809 10/31/10 13:38 a-onl End_of_Run_8171
  • 337808 10/31/10 13:35 chench ROCs for run 8172
  • 337807 10/31/10 13:34 chench DAQ and scalers for run8172
  • 337806 10/31/10 12:55 a-onl Start_Run_8172,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337805 10/31/10 12:51 a-onl End_of_Run_8163
  • 337804 10/31/10 12:45 chench DAQ and scalers for run8171
  • 337803 10/31/10 12:44 chench ROCs for run8171
  • 337802 10/31/10 12:03 a-onl Start_Run_8171,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337801 10/31/10 12:01 a-onl End_of_Run_8170
  • 337800 10/31/10 11:59 chench DAQ and scalers for run8170
  • 337799 10/31/10 11:57 A. Camsonne Compton rates flatlined at 0 - rebooting the Compton DAQ
  • 337798 10/31/10 11:56 chench Rocs for run8170
  • 337797 10/31/10 11:53 auto_snap Radiation and Ozone snapshot
  • 337796 10/31/10 11:53 auto_snap Miscellaneous snapshot
  • 337795 10/31/10 11:53 auto_snap Hall A Tools snapshot
  • 337794 10/31/10 11:53 auto_snap Compton Polarimeter snapshot
  • 337793 10/31/10 11:53 auto_snap Beam Position and Current snapshot
  • 337792 10/31/10 11:11 a-onl Start_Run_8170,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337791 10/31/10 11:10 a-onl End_of_Run_8169
  • 337790 10/31/10 10:59 auto_snap Radiation and Ozone snapshot
  • 337789 10/31/10 10:59 auto_snap Miscellaneous snapshot
  • 337788 10/31/10 10:59 auto_snap Hall A Tools snapshot
  • 337787 10/31/10 10:58 auto_snap Compton Polarimeter snapshot
  • 337786 10/31/10 10:58 auto_snap Beam Position and Current snapshot
  • 337785 10/31/10 10:53 pzhu PbF2 Calorimeter ADC Readback
  • 337784 10/31/10 10:50 chench DAQ and scalers for run8169
  • 337783 10/31/10 10:47 chench ROCs for run(8169)
  • 337782 10/31/10 10:02 a-onl Start_Run_8169,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337781 10/31/10 10:01 a-onl End_of_Run_8168
  • 337780 10/31/10 09:50 chench DAQ and scalers (run8618)
  • 337779 10/31/10 09:49 chench Rocs for run 8168
  • 337778 10/31/10 09:11 a-onl Start_Run_8168,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337777 10/31/10 08:55 a-onl End_of_Run_8167
  • 337776 10/31/10 08:35 a-onl Start_Run_8167,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337775 10/31/10 08:34 a-onl End_of_Run_8166
  • 337774 10/31/10 08:13 adaq Beam Time Accounting
  • 337773 10/31/10 08:09 hyde Cerenkov efficiency
  • 337772 10/31/10 07:48 hyde Owl Shift Summary
  • 337771 10/31/10 07:42 a-onl Start_Run_8166,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337770 10/31/10 07:41 a-onl End_of_Run_8165
  • 337769 10/31/10 07:40 D.Wang rocs (8165)
  • 337768 10/31/10 07:39 D.Wang coda and scalers (8165)
  • 337767 10/31/10 06:55 a-onl Start_Run_8165,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337766 10/31/10 06:54 a-onl End_of_Run_8164
  • 337765 10/31/10 06:52 D.Wang rocs (8164)
  • 337764 10/31/10 06:52 hyde Normalization check
  • 337763 10/31/10 06:51 D.Wang coda and scalers(8164)
  • 337762 10/31/10 06:07 a-onl Start_Run_8164,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337761 10/31/10 05:59 lselvy Target Update
  • 337760 10/31/10 05:22 a-onl Start_Run_8163,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337759 10/31/10 05:20 a-onl End_of_Run_8162
  • 337758 10/31/10 05:19 D.Wang rocs (8162)
  • 337757 10/31/10 05:18 D.Wang coda and scalers (8162)
  • 337756 10/31/10 04:48 hyde coda crash
  • 337755 10/31/10 04:45 hyde Normalization Runs
  • 337754 10/31/10 04:34 a-onl Start_Run_8162,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 LeftHRS ps5=1
  • 337753 10/31/10 04:26 a-onl End_of_Run_8161
  • 337752 10/31/10 04:08 a-onl End_of_Run_8160
  • 337751 10/31/10 04:02 lselvy Helicity Error
  • 337750 10/31/10 03:56 a-onl Start_Run_8160,Run_type=Check out,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Normalization 4 pass Kin 3 LeftHRS ps5=1
  • 337749 10/31/10 03:53 auto_snap Radiation and Ozone snapshot
  • 337748 10/31/10 03:53 auto_snap Miscellaneous snapshot
  • 337747 10/31/10 03:53 auto_snap Hall A Tools snapshot
  • 337746 10/31/10 03:53 auto_snap Compton Polarimeter snapshot
  • 337745 10/31/10 03:53 auto_snap Beam Position and Current snapshot
  • 337744 10/31/10 03:52 a-onl End_of_Run_8159
  • 337743 10/31/10 03:51 D.Wang rocs (8159)
  • 337742 10/31/10 03:51 D.Wang coda and scalers (8159)
  • 337741 10/31/10 03:47 hyde calo dc anode current
  • 337740 10/31/10 03:45 hyde beam strip chart
  • 337739 10/31/10 03:41 hyde Cerenkov timing in coincidence
  • 337738 10/31/10 03:34 hyde Hall A Tools Snapshot
  • 337737 10/31/10 03:06 a-onl Start_Run_8159,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff
  • 337736 10/31/10 03:06 a-onl End_of_Run_8158
  • 337735 10/31/10 03:05 D.Wang rocs (8158)
  • 337734 10/31/10 03:04 D.Wang coda and scalers(8158)
  • 337733 10/31/10 02:58 auto_snap Radiation and Ozone snapshot
  • 337732 10/31/10 02:58 auto_snap Miscellaneous snapshot
  • 337731 10/31/10 02:58 auto_snap Hall A Tools snapshot
  • 337730 10/31/10 02:58 auto_snap Compton Polarimeter snapshot
  • 337729 10/31/10 02:58 auto_snap Beam Position and Current snapshot
  • 337728 10/31/10 02:26 a-onl Start_Run_8158,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff
  • 337727 10/31/10 02:25 a-onl End_of_Run_8157
  • 337726 10/31/10 02:16 Carlos Production in buffering mode
  • 337725 10/31/10 02:12 D.Wang rocs (8157)
  • 337724 10/31/10 02:12 D.Wang coda and scalers (8157)
  • 337723 10/31/10 02:07 HYDE DVCS CALO Anode current
  • 337722 10/31/10 01:42 a-onl Start_Run_8157,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff
  • 337721 10/31/10 01:39 a-onl End_of_Run_8155
  • 337720 10/31/10 01:39 a-onl End_of_Run_8156
  • 337719 10/31/10 01:37 a-onl Start_Run_8156,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff
  • 337718 10/31/10 01:37 a-onl Start_Run_8154,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff
  • 337717 10/31/10 01:35 a-onl Start_Run_8155,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCS
  • 337716 10/31/10 01:26 a-onl Start_Run_8153,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff
  • 337715 10/31/10 01:26 a-onl End_of_Run_8154
  • 337714 10/31/10 01:19 a-onl Start_Run_8152,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff
  • 337713 10/31/10 01:07 a-onl Start_Run_8149,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff
  • 337712 10/31/10 01:06 a-onl End_of_Run_8148
  • 337711 10/31/10 00:58 hyde raster
  • 337710 10/31/10 00:50 a-onl Start_Run_8148,Run_type=Check out,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=10,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff
  • 337709 10/31/10 00:42 hyde calo HV on
  • 337708 10/31/10 00:33 a-onl End_of_Run_8147
  • 337707 10/31/10 00:30 a-onl Start_Run_8147,Run_type=Check out,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=10,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 with buffering 10uA
  • 337706 10/31/10 00:28 hyde calo HV off
  • 337705 10/31/10 00:23 a-onl End_of_Run_8146
  • 337704 10/31/10 00:14 a-onl Start_Run_8146,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=4,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 with buffering 4uA
  • 337703 10/31/10 00:10 a-onl End_of_Run_8145
  • 337702 10/31/10 00:04 a-onl Start_Run_8145,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 with buffering 3uA
  • 337701 10/31/10 00:02 a-onl End_of_Run_8144
  • 337700 10/31/10 00:01 A. Camsonne End of run 8144
  • 337699 10/31/10 00:00 adaq Beam Time Accounting
  • Saturday

  • 337698 10/30/10 23:58 A. Tobias Target Status: Loop 2 LH2
  • 337697 10/30/10 23:54 A. Tobias Target Stripcharts
  • 337696 10/30/10 23:53 Ramesh Shift Summary
  • 337695 10/30/10 23:51 a-onl Start_Run_8144,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 with buffering
  • 337694 10/30/10 23:51 Abdurahim ROCs and Scaler screen snapshot
  • 337693 10/30/10 23:50 a-onl End_of_Run_8143
  • 337692 10/30/10 23:43 A. Tobias Loop 2 (LH2) Fan set to 19% = 46Hz
  • 337691 10/30/10 23:31 a-onl Start_Run_8143,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 with buffering
  • 337690 10/30/10 23:30 A. Camsonne Run with buffering 8143
  • 337689 10/30/10 23:28 a-onl End_of_Run_8142
  • 337688 10/30/10 23:17 a-onl Start_Run_8142,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=ps9=1, others are zero, 10 min run with lhrs config
  • 337687 10/30/10 23:17 Ramesh ps9=1 run for 10 min
  • 337686 10/30/10 23:15 a-onl End_of_Run_8141
  • 337685 10/30/10 23:14 a-onl End_of_Run_8140
  • 337684 10/30/10 23:14 Abdurahim ROCs and Scaler run8140
  • 337683 10/30/10 23:13 A. Camsonne Compton HV on - started run in backroom -
  • 337682 10/30/10 23:12 a-onl End_of_Run_8135
  • 337681 10/30/10 22:18 Abdurahim TS11 is fine now
  • 337680 10/30/10 22:16 a-onl Start_Run_8140,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware
  • 337679 10/30/10 22:15 Abdurahim Error massage in TS11
  • 337678 10/30/10 22:08 a-onl End_of_Run_8139
  • 337677 10/30/10 22:06 a-onl Start_Run_8139,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware
  • 337676 10/30/10 22:06 a-onl End_of_Run_8138
  • 337675 10/30/10 21:38 R. Michaels re: number of events in halog end-of-run
  • 337674 10/30/10 21:25 Abdurahim ROCs and Scalers snapshot run8137
  • 337673 10/30/10 21:23 a-onl Start_Run_8138,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware
  • 337672 10/30/10 21:22 a-onl End_of_Run_8137
  • 337671 10/30/10 20:36 Abdurahim ROC s and Scalers before the end of run8136
  • 337670 10/30/10 20:35 a-onl Start_Run_8137,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware
  • 337669 10/30/10 20:34 a-onl End_of_Run_8136
  • 337668 10/30/10 20:19 A. Tobias Target Screenshots
  • 337667 10/30/10 19:53 auto_snap Radiation and Ozone snapshot
  • 337666 10/30/10 19:53 auto_snap Miscellaneous snapshot
  • 337665 10/30/10 19:53 auto_snap Hall A Tools snapshot
  • 337664 10/30/10 19:53 auto_snap Compton Polarimeter snapshot
  • 337663 10/30/10 19:53 auto_snap Beam Position and Current snapshot
  • 337662 10/30/10 19:28 a-onl Start_Run_8136,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware
  • 337661 10/30/10 19:27 Abdurahim ROCs, Scalers and CODA screen snapshots before the end of 8135
  • 337660 10/30/10 19:27 a-onl End_of_Run_8132
  • 337659 10/30/10 19:21 Ramesh Discrepancy between Coda_event_number and end_of_run_event_number
  • 337658 10/30/10 19:11 Abdurahim CALO TDC histo mismacth run8133
  • 337657 10/30/10 19:09 Abdurahim BPMAx histo mismatch run8133
  • 337656 10/30/10 19:05 A. Tobias Target: Loop 2 Fan settings
  • 337655 10/30/10 19:00 A. Tobias Target: Loop 2 Fan Settings
  • 337654 10/30/10 19:00 Ramesh Calorimeter current
  • 337653 10/30/10 18:59 A. Tobias Loop 1 Fan speed recovers to 24Hz
  • 337652 10/30/10 18:58 auto_snap Radiation and Ozone snapshot
  • 337651 10/30/10 18:58 auto_snap Miscellaneous snapshot
  • 337650 10/30/10 18:58 auto_snap Hall A Tools snapshot
  • 337649 10/30/10 18:58 auto_snap Compton Polarimeter snapshot
  • 337648 10/30/10 18:58 auto_snap Beam Position and Current snapshot
  • 337647 10/30/10 18:55 Abdurahim mismatched histos in CALO TDC run8134
  • 337646 10/30/10 18:53 Abdurahim mismatched histo in BPM run8134
  • 337645 10/30/10 18:48 Abdurahim Strange Histo in Left arm S2 ADC
  • 337644 10/30/10 18:37 Ramesh Halog down intermittently
  • 337643 10/30/10 18:37 A. Tobias CryoTarget Fans
  • 337642 10/30/10 18:23 Ramesh Flaky Q3 liquid level!
  • 337641 10/30/10 18:23 Abdurahim ROCs, Scalers and CODA screens before the end of run8134
  • 337640 10/30/10 18:22 a-onl Start_Run_8135,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware
  • 337639 10/30/10 18:21 a-onl End_of_Run_8134
  • 337638 10/30/10 17:28 Abdurahim ROCS and CODA screen snapshot before the end of 8133
  • 337637 10/30/10 17:27 a-onl Start_Run_8134,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware
  • 337636 10/30/10 17:26 a-onl End_of_Run_8133
  • 337635 10/30/10 16:30 A. Tobias Target moved to Loop2 H2 15cm
  • 337634 10/30/10 16:26 a-onl Start_Run_8133,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware
  • 337633 10/30/10 16:14 A. Camsonne Added tsUtil.o to trigger_left.boot
  • 337632 10/30/10 16:11 a-onl Start_Run_8132,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337631 10/30/10 16:03 bogdanw shift summary
  • 337630 10/30/10 16:02 adaq Beam Time Accounting
  • 337629 10/30/10 15:19 A. Camsonne Sync ADC work
  • 337628 10/30/10 15:19 R. Michaels neg. deadtimes fixed
  • 337627 10/30/10 14:59 a-onl Start_Run_8131,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337626 10/30/10 14:56 a-onl End_of_Run_8130
  • 337625 10/30/10 14:54 a-onl Start_Run_8130,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337624 10/30/10 14:54 a-onl End_of_Run_8129
  • 337623 10/30/10 14:50 a-onl Start_Run_8129,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337622 10/30/10 14:49 a-onl End_of_Run_8125
  • 337621 10/30/10 14:46 a-onl Start_Run_8128,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337620 10/30/10 14:45 a-onl End_of_Run_8127
  • 337619 10/30/10 14:42 a-onl Start_Run_8127,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337618 10/30/10 14:42 a-onl End_of_Run_8126
  • 337617 10/30/10 14:38 a-onl Start_Run_8126,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337616 10/30/10 14:37 a-onl End_of_Run_8124
  • 337615 10/30/10 14:35 a-onl Start_Run_8125,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337614 10/30/10 14:34 a-onl End_of_Run_8122
  • 337613 10/30/10 14:32 a-onl Start_Run_8124,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337612 10/30/10 14:31 a-onl End_of_Run_8123
  • 337611 10/30/10 14:28 a-onl Start_Run_8123,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337610 10/30/10 14:27 a-onl End_of_Run_8119
  • 337609 10/30/10 14:25 a-onl Start_Run_8122,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337608 10/30/10 14:25 a-onl End_of_Run_8121
  • 337607 10/30/10 14:23 a-onl Start_Run_8121,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337606 10/30/10 14:23 a-onl End_of_Run_8120
  • 337605 10/30/10 14:20 a-onl Start_Run_8120,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337604 10/30/10 14:19 a-onl End_of_Run_8116
  • 337603 10/30/10 14:13 a-onl Start_Run_8119,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337602 10/30/10 14:11 a-onl End_of_Run_8118
  • 337601 10/30/10 14:10 a-onl Start_Run_8118,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC 14 ROC17 flipped
  • 337600 10/30/10 14:09 a-onl End_of_Run_8117
  • 337599 10/30/10 14:05 a-onl Start_Run_8117,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work - ROC17 flipped
  • 337598 10/30/10 14:03 a-onl End_of_Run_8115
  • 337597 10/30/10 14:00 a-onl Start_Run_8116,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337596 10/30/10 13:55 a-onl Start_Run_8115,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337595 10/30/10 13:53 a-onl End_of_Run_8113
  • 337594 10/30/10 13:48 a-onl Start_Run_8113,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337593 10/30/10 13:47 a-onl End_of_Run_8112
  • 337592 10/30/10 13:41 a-onl Start_Run_8112,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337591 10/30/10 13:40 a-onl End_of_Run_8107
  • 337590 10/30/10 13:37 a-onl Start_Run_8111,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337589 10/30/10 13:37 a-onl End_of_Run_8110
  • 337588 10/30/10 13:28 a-onl Start_Run_8110,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337587 10/30/10 13:27 a-onl End_of_Run_8109
  • 337586 10/30/10 13:23 a-onl Start_Run_8109,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337585 10/30/10 13:22 a-onl End_of_Run_8108
  • 337584 10/30/10 13:19 a-onl Start_Run_8108,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337583 10/30/10 13:19 a-onl End_of_Run_8105
  • 337582 10/30/10 13:16 a-onl Start_Run_8107,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337581 10/30/10 13:15 a-onl End_of_Run_8106
  • 337580 10/30/10 13:13 a-onl Start_Run_8106,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337579 10/30/10 13:11 a-onl End_of_Run_8103
  • 337578 10/30/10 13:09 a-onl Start_Run_8105,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337577 10/30/10 13:08 a-onl End_of_Run_8104
  • 337576 10/30/10 13:02 a-onl Start_Run_8104,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337575 10/30/10 13:01 a-onl End_of_Run_8102
  • 337574 10/30/10 12:59 a-onl Start_Run_8103,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337573 10/30/10 12:59 a-onl End_of_Run_8100
  • 337572 10/30/10 12:53 a-onl Start_Run_8102,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337571 10/30/10 12:52 a-onl End_of_Run_8101
  • 337570 10/30/10 12:48 a-onl Start_Run_8101,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337569 10/30/10 12:48 a-onl End_of_Run_8097
  • 337568 10/30/10 12:42 a-onl Start_Run_8100,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337567 10/30/10 12:41 a-onl End_of_Run_8099
  • 337566 10/30/10 12:31 a-onl Start_Run_8099,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337565 10/30/10 12:31 a-onl End_of_Run_8098
  • 337564 10/30/10 12:24 a-onl Start_Run_8098,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337563 10/30/10 12:24 a-onl End_of_Run_8095
  • 337562 10/30/10 12:22 a-onl Start_Run_8097,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337561 10/30/10 12:22 a-onl End_of_Run_8096
  • 337560 10/30/10 12:15 a-onl Start_Run_8096,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337559 10/30/10 12:08 a-onl Start_Run_8095,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware - Sync check ADC work
  • 337558 10/30/10 12:03 a-onl End_of_Run_8094
  • 337557 10/30/10 12:01 adaq hall status
  • 337556 10/30/10 11:53 auto_snap Radiation and Ozone snapshot
  • 337555 10/30/10 11:52 auto_snap Miscellaneous snapshot
  • 337554 10/30/10 11:52 auto_snap Hall A Tools snapshot
  • 337553 10/30/10 11:52 auto_snap Compton Polarimeter snapshot
  • 337552 10/30/10 11:52 auto_snap Beam Position and Current snapshot
  • 337551 10/30/10 11:37 a-onl Start_Run_8094,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware
  • 337550 10/30/10 11:34 a-onl Start_Run_8093,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware
  • 337549 10/30/10 11:33 a-onl End_of_Run_8092
  • 337548 10/30/10 11:32 a-onl End_of_Run_8088
  • 337547 10/30/10 11:29 A. Camsonne DVCS Trigger firmware loaded Run 8091
  • 337546 10/30/10 11:24 A. Camsonne Empty FIFO messages in ROC14 and ROC18
  • 337545 10/30/10 11:20 A. Camsonne DVCS Trigger module pedestals
  • 337544 10/30/10 11:06 a-onl Start_Run_8092,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware
  • 337543 10/30/10 11:00 a-onl End_of_Run_8091
  • 337542 10/30/10 10:58 auto_snap Radiation and Ozone snapshot
  • 337540 10/30/10 10:58 auto_snap Miscellaneous snapshot
  • 337539 10/30/10 10:58 auto_snap Hall A Tools snapshot
  • 337538 10/30/10 10:58 auto_snap Compton Polarimeter snapshot
  • 337537 10/30/10 10:58 a-onl Start_Run_8091,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - new Trigger firmware
  • 337536 10/30/10 10:58 auto_snap Beam Position and Current snapshot
  • 337535 10/30/10 10:19 flay Run 8089: S2m ADCs
  • 337534 10/30/10 10:15 a-onl End_of_Run_8090
  • 337533 10/30/10 10:15 flay Run 8090 Stopped Early
  • 337532 10/30/10 10:05 a-onl Start_Run_8090,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337531 10/30/10 10:04 a-onl End_of_Run_8089
  • 337530 10/30/10 10:03 flay Run 8089 Trigger/ROC Snapshots
  • 337529 10/30/10 09:15 flay Run 8088: Lumis (BCM Normalized) Rates vs. Event # Has Odd Structure
  • 337528 10/30/10 09:03 a-onl Start_Run_8089,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337527 10/30/10 09:02 a-onl End_of_Run_8084
  • 337526 10/30/10 09:01 flay Run 8088 Trigger/ROC Snapshots
  • 337525 10/30/10 08:26 Carlos Re: Compton history plot
  • 337524 10/30/10 08:23 adaq Compton CODA
  • 337522 10/30/10 08:19 efuchey 10-30-10 Owl Shift summary
  • 337523 10/30/10 08:18 adaq Beam Time Accounting
  • 337521 10/30/10 08:02 a-onl Start_Run_8088,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337520 10/30/10 08:01 efuchey Compton data taking restarted
  • 337519 10/30/10 08:00 a-onl End_of_Run_8087
  • 337518 10/30/10 07:51 D.Wang rocs (8087)
  • 337517 10/30/10 07:50 D.Wang coda and scalers (8087)
  • 337516 10/30/10 07:12 a-onl Start_Run_8087,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337515 10/30/10 07:10 a-onl End_of_Run_8086
  • 337514 10/30/10 07:07 D.Wang rocs (8086)
  • 337513 10/30/10 07:07 D.Wang coda and scalers (8086)
  • 337512 10/30/10 06:46 efuchey Compton Rates 10-31
  • 337511 10/30/10 06:27 efuchey Rates going wild
  • 337510 10/30/10 06:09 a-onl Start_Run_8086,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337509 10/30/10 06:09 a-onl End_of_Run_8085
  • 337508 10/30/10 06:08 lselvy Target Update
  • 337507 10/30/10 06:05 D.Wang rocs (8085)
  • 337506 10/30/10 06:04 D.Wang coda and scalers (8085)
  • 337505 10/30/10 05:21 a-onl Start_Run_8085,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337504 10/30/10 05:13 D.Wang rocs (8084)
  • 337503 10/30/10 05:13 D.Wang coda and scalers (8084)
  • 337502 10/30/10 04:33 a-onl Start_Run_8084,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337501 10/30/10 04:24 a-onl End_of_Run_8083
  • 337500 10/30/10 04:23 a-onl End_of_Run_8082
  • 337499 10/30/10 04:22 a-onl End_of_Run_8081
  • 337498 10/30/10 04:19 D.Wang rocs (8082)
  • 337497 10/30/10 04:18 D.Wang coda and scalers (8082)
  • 337496 10/30/10 03:52 auto_snap Radiation and Ozone snapshot
  • 337495 10/30/10 03:52 auto_snap Miscellaneous snapshot
  • 337494 10/30/10 03:52 auto_snap Hall A Tools snapshot
  • 337493 10/30/10 03:52 auto_snap Compton Polarimeter snapshot
  • 337492 10/30/10 03:52 auto_snap Beam Position and Current snapshot
  • 337491 10/30/10 03:35 a-onl Start_Run_8082,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337490 10/30/10 03:35 a-onl End_of_Run_8070
  • 337489 10/30/10 03:29 D.Wang rocs (8081)
  • 337488 10/30/10 03:28 D.Wang coda and scalers (8081)
  • 337487 10/30/10 03:01 efuchey Update on the asymmetry history
  • 337486 10/30/10 02:57 auto_snap Radiation and Ozone snapshot
  • 337485 10/30/10 02:57 auto_snap Miscellaneous snapshot
  • 337484 10/30/10 02:57 auto_snap Hall A Tools snapshot
  • 337483 10/30/10 02:57 auto_snap Compton Polarimeter snapshot
  • 337482 10/30/10 02:57 auto_snap Beam Position and Current snapshot
  • 337481 10/30/10 02:42 efuchey First asymmetry history
  • 337480 10/30/10 02:41 a-onl Start_Run_8081,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337479 10/30/10 02:41 a-onl End_of_Run_8080
  • 337478 10/30/10 02:34 D.Wang rocs (8080)
  • 337477 10/30/10 02:33 D.Wang coda and scalers (8080)
  • 337476 10/30/10 01:56 efuchey Junky Compton "production"
  • 337475 10/30/10 01:52 a-onl Start_Run_8080,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337474 10/30/10 01:51 a-onl End_of_Run_8079
  • 337473 10/30/10 01:50 D.Wang rocs (8079)
  • 337472 10/30/10 01:49 D.Wang coda and scalers (8079)
  • 337471 10/30/10 01:03 a-onl Start_Run_8079,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337470 10/30/10 01:02 a-onl End_of_Run_8078
  • 337469 10/30/10 00:59 D.Wang rocs (8078)
  • 337468 10/30/10 00:59 D.Wang coda and scalers (8078)
  • 337467 10/30/10 00:21 efuchey Hall A tools screenshot.
  • 337466 10/30/10 00:19 efuchey Calo base currents
  • 337465 10/30/10 00:11 a-onl Start_Run_8078,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337464 10/30/10 00:11 a-onl End_of_Run_8077
  • 337463 10/30/10 00:10 D.Wang rocs (run 8077)
  • 337462 10/30/10 00:09 D.Wang coda and scalers
  • 337460 10/30/10 00:00 GOLGE swing shift summary
  • 337461 10/30/10 00:00 adaq Beam Time Accounting
  • Friday

  • 337459 10/29/10 23:48 efuchey Compton rates after beam trip
  • 337458 10/29/10 23:24 Golge ARC magnet trip, beam off
  • 337457 10/29/10 23:24 a-onl Start_Run_8077,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337456 10/29/10 22:21 a-onl End_of_Run_8076
  • 337455 10/29/10 22:20 a-onl Start_Run_8076,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337454 10/29/10 22:18 a-onl End_of_Run_8075
  • 337453 10/29/10 22:11 chench screenshot of roc18 for run8075
  • 337452 10/29/10 22:10 chench screenshot for ROC14 and 17 for run8075
  • 337451 10/29/10 22:10 chench screenshot of DAQ window and scalers for run8075
  • 337450 10/29/10 21:19 a-onl Start_Run_8075,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337449 10/29/10 21:18 a-onl End_of_Run_8074
  • 337448 10/29/10 21:17 chench correction: the run muber is 8074
  • 337447 10/29/10 21:16 chench corrections: the run number is 8074
  • 337446 10/29/10 21:14 chench screenshot of roc18 for run8074
  • 337445 10/29/10 21:13 chench screenshot of ROC14 and 17 for run8075
  • 337444 10/29/10 21:13 chench screenshot of DAQ window and scalers for run8075
  • 337443 10/29/10 20:27 a-onl Start_Run_8074,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337442 10/29/10 20:26 a-onl End_of_Run_8073
  • 337441 10/29/10 20:22 chench screenshot of ROC18 for run8073
  • 337440 10/29/10 20:21 chench screenshot of ROC14 and ROC17
  • 337439 10/29/10 20:21 chench Screenshot of DAQ window and scalers for run8073
  • 337438 10/29/10 19:52 auto_snap Radiation and Ozone snapshot
  • 337437 10/29/10 19:52 auto_snap Miscellaneous snapshot
  • 337436 10/29/10 19:52 auto_snap Hall A Tools snapshot
  • 337435 10/29/10 19:52 auto_snap Compton Polarimeter snapshot
  • 337434 10/29/10 19:52 auto_snap Beam Position and Current snapshot
  • 337433 10/29/10 19:38 a-onl Start_Run_8073,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337432 10/29/10 19:35 a-onl End_of_Run_8072
  • 337431 10/29/10 19:30 chench screenshot of roc18 for run8072
  • 337430 10/29/10 19:30 chench screenshot of ROC14 and 17 for run8072
  • 337429 10/29/10 19:29 chench screenshot of DAQ window and scalers for run8072
  • 337428 10/29/10 18:57 auto_snap Radiation and Ozone snapshot
  • 337427 10/29/10 18:57 auto_snap Miscellaneous snapshot
  • 337426 10/29/10 18:57 auto_snap Hall A Tools snapshot
  • 337425 10/29/10 18:57 auto_snap Compton Polarimeter snapshot
  • 337424 10/29/10 18:57 auto_snap Beam Position and Current snapshot
  • 337423 10/29/10 18:44 a-onl Start_Run_8072,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337422 10/29/10 18:42 a-onl End_of_Run_8071
  • 337421 10/29/10 18:41 A. Camsonne Added clear of SIS scalers in DVCS crates
  • 337420 10/29/10 18:32 chench screenshot for roc18 for run8071
  • 337419 10/29/10 18:32 chench screenshot of ROC14 and ROC17 for run8071
  • 337418 10/29/10 18:31 chench Screenshot of DAQ window and scalers for run8071
  • 337417 10/29/10 17:52 a-onl Start_Run_8071,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337416 10/29/10 17:41 chench Screenshot of roc18 for run8070
  • 337415 10/29/10 17:41 chench Sreenshot of ROC14 and ROC17 for run8070
  • 337414 10/29/10 17:40 chench Screenshot of DAQ window and scalers for run8070
  • 337413 10/29/10 17:32 hyde clock consistency check
  • 337412 10/29/10 17:02 a-onl Start_Run_8070,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337411 10/29/10 17:01 a-onl End_of_Run_8069
  • 337410 10/29/10 16:54 chench screenshot of ROC18 for run8069
  • 337409 10/29/10 16:53 chench screenshot for ROC14 and ROC17 for run8069
  • 337408 10/29/10 16:52 chench Screenshot of DAQ and scalers for run 8069
  • 337407 10/29/10 16:22 efuchey Compton rates for 10-29-10
  • 337406 10/29/10 16:12 a-onl Start_Run_8069,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337405 10/29/10 16:08 a-onl End_of_Run_8068
  • 337404 10/29/10 16:03 A. Camsonne DAQ Sync work
  • 337403 10/29/10 16:02 Saha Shift Summary
  • 337402 10/29/10 16:01 adaq Beam Time Accounting
  • 337401 10/29/10 16:00 a-onl Start_Run_8068,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337400 10/29/10 15:52 a-onl End_of_Run_8067
  • 337399 10/29/10 15:52 R. Michaels target to LH2, 15 cm
  • 337398 10/29/10 15:45 Ole Hansen Analyzer updated
  • 337397 10/29/10 15:38 R. Michaels moving to empty target
  • 337396 10/29/10 15:34 a-onl Start_Run_8067,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337395 10/29/10 15:25 a-onl End_of_Run_8065
  • 337394 10/29/10 15:23 a-onl Start_Run_8066,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337393 10/29/10 15:23 a-onl End_of_Run_8063
  • 337392 10/29/10 15:20 a-onl Start_Run_8065,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337391 10/29/10 15:19 a-onl End_of_Run_8064
  • 337390 10/29/10 15:16 a-onl Start_Run_8064,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337389 10/29/10 15:16 a-onl End_of_Run_8060
  • 337388 10/29/10 15:12 a-onl Start_Run_8063,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337387 10/29/10 15:12 a-onl End_of_Run_8062
  • 337386 10/29/10 15:06 a-onl Start_Run_8062,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337385 10/29/10 15:05 a-onl End_of_Run_8061
  • 337384 10/29/10 15:00 a-onl Start_Run_8061,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337383 10/29/10 14:50 a-onl Start_Run_8060,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337382 10/29/10 14:48 a-onl Start_Run_8059,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337381 10/29/10 14:46 a-onl Start_Run_8058,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337380 10/29/10 14:45 a-onl End_of_Run_8057
  • 337379 10/29/10 14:39 a-onl Start_Run_8057,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337378 10/29/10 14:38 a-onl End_of_Run_8055
  • 337377 10/29/10 14:35 a-onl Start_Run_8056,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337376 10/29/10 14:34 a-onl End_of_Run_8049
  • 337375 10/29/10 14:31 a-onl Start_Run_8055,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337374 10/29/10 14:30 a-onl End_of_Run_8054
  • 337373 10/29/10 14:28 a-onl Start_Run_8054,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337372 10/29/10 14:27 efuchey QWP scan for 10/29/10
  • 337371 10/29/10 14:27 a-onl End_of_Run_8053
  • 337370 10/29/10 14:25 a-onl Start_Run_8053,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337369 10/29/10 14:24 a-onl End_of_Run_8052
  • 337368 10/29/10 14:21 a-onl Start_Run_8052,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337367 10/29/10 14:20 a-onl End_of_Run_8051
  • 337366 10/29/10 14:16 a-onl Start_Run_8051,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337365 10/29/10 14:15 a-onl End_of_Run_8050
  • 337364 10/29/10 14:09 a-onl Start_Run_8050,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337363 10/29/10 14:09 a-onl End_of_Run_8048
  • 337362 10/29/10 14:01 a-onl Start_Run_8049,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337361 10/29/10 13:55 a-onl Start_Run_8048,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337360 10/29/10 13:55 hyde DVCS trig scaler
  • 337359 10/29/10 13:55 a-onl End_of_Run_8046
  • 337358 10/29/10 13:55 a-onl End_of_Run_8047
  • 337357 10/29/10 13:50 a-onl Start_Run_8047,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337356 10/29/10 13:50 a-onl End_of_Run_8045
  • 337355 10/29/10 13:45 a-onl Start_Run_8046,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337354 10/29/10 13:39 a-onl Start_Run_8045,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337353 10/29/10 13:38 a-onl End_of_Run_8044
  • 337352 10/29/10 13:37 a-onl End_of_Run_8043
  • 337351 10/29/10 13:31 a-onl Start_Run_8044,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337350 10/29/10 13:30 a-onl End_of_Run_8042
  • 337349 10/29/10 13:24 a-onl Start_Run_8043,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337348 10/29/10 13:23 a-onl End_of_Run_8041
  • 337347 10/29/10 13:18 a-onl Start_Run_8042,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337346 10/29/10 13:17 a-onl End_of_Run_8039
  • 337345 10/29/10 13:15 a-onl Start_Run_8041,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337344 10/29/10 13:14 a-onl End_of_Run_8040
  • 337343 10/29/10 13:09 a-onl Start_Run_8040,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337342 10/29/10 13:04 a-onl Start_Run_8039,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass - sync DAQ work
  • 337341 10/29/10 12:51 R. Michaels helicity stable for 15 min
  • 337340 10/29/10 12:39 R. Michaels occassional "bad" helicity
  • 337339 10/29/10 12:35 a-onl End_of_Run_8037
  • 337338 10/29/10 12:28 R. Michaels re: last scaler in datastream
  • 337337 10/29/10 12:28 A. Camsonne Sync ADC ROC3 and ROC14
  • 337336 10/29/10 12:25 A. Camsonne Sync ADC for ROC17 and ROC4 look ok - need to check ROC3 and ROC14
  • 337335 10/29/10 11:55 a-onl Start_Run_8037,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337334 10/29/10 11:53 a-onl End_of_Run_8036
  • 337333 10/29/10 11:52 auto_snap Radiation and Ozone snapshot
  • 337332 10/29/10 11:52 auto_snap Miscellaneous snapshot
  • 337331 10/29/10 11:52 auto_snap Hall A Tools snapshot
  • 337330 10/29/10 11:51 auto_snap Compton Polarimeter snapshot
  • 337329 10/29/10 11:51 auto_snap Beam Position and Current snapshot
  • 337328 10/29/10 11:51 kelleher warming up Moller
  • 337327 10/29/10 11:35 Carlos Re: evtyebits problems
  • 337326 10/29/10 11:13 R. Michaels evtyebits problems.
  • 337325 10/29/10 11:03 a-onl Start_Run_8036,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337324 10/29/10 11:02 a-onl Start_Run_8034,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337323 10/29/10 11:00 Punjabi run 8035 ROC screens
  • 337322 10/29/10 10:59 Punjabi RC and scaler screens
  • 337321 10/29/10 10:57 auto_snap Radiation and Ozone snapshot
  • 337320 10/29/10 10:57 auto_snap Miscellaneous snapshot
  • 337319 10/29/10 10:57 auto_snap Hall A Tools snapshot
  • 337318 10/29/10 10:57 auto_snap Compton Polarimeter snapshot
  • 337317 10/29/10 10:57 auto_snap Beam Position and Current snapshot
  • 337316 10/29/10 10:55 jpchen target is fine
  • 337315 10/29/10 10:26 R. Michaels updated halla_scaler_process (halog scalers)
  • 337314 10/29/10 10:17 Carlos Deadtime study special runs
  • 337313 10/29/10 10:15 a-onl Start_Run_8035,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337312 10/29/10 10:12 a-onl End_of_Run_8034
  • 337311 10/29/10 10:09 a-onl End_of_Run_8033
  • 337310 10/29/10 10:06 a-onl Start_Run_8033,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass (S1 OFF)
  • 337309 10/29/10 10:05 a-onl End_of_Run_8032
  • 337308 10/29/10 10:01 a-onl Start_Run_8032,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass (S1 OFF)
  • 337307 10/29/10 09:59 a-onl End_of_Run_8031
  • 337306 10/29/10 09:59 a-onl End_of_Run_8027
  • 337305 10/29/10 09:42 a-onl Start_Run_8031,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337304 10/29/10 09:41 a-onl End_of_Run_8030
  • 337303 10/29/10 09:39 a-onl Start_Run_8030,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337302 10/29/10 09:37 a-onl End_of_Run_8029
  • 337301 10/29/10 09:06 a-onl Start_Run_8029,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 3, 4 pass
  • 337300 10/29/10 09:05 Punjabi End of run ROC screens
  • 337299 10/29/10 09:04 a-onl End_of_Run_8028
  • 337298 10/29/10 08:22 R. Michaels target instabilities
  • 337297 10/29/10 08:12 a-onl Start_Run_8028,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337296 10/29/10 08:10 a-onl End_of_Run_8025
  • 337295 10/29/10 08:05 jroche owl shift summary
  • 337294 10/29/10 08:01 adaq Beam Time Accounting
  • 337293 10/29/10 08:00 flay Run 8026 Trigger/ROC Snapshots
  • 337292 10/29/10 07:10 jroche "new" results for the BCM calibration
  • 337291 10/29/10 07:08 a-onl Start_Run_8027,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337290 10/29/10 07:07 a-onl End_of_Run_8026
  • 337289 10/29/10 07:06 flay Run 8026 Trigger/ROC Snapshots
  • 337288 10/29/10 06:24 jroche Counting room somewhat cold since an hour or so
  • 337287 10/29/10 06:18 jroche End of run halog not showing up or very slow to show up
  • 337286 10/29/10 06:11 a-onl Start_Run_8026,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337285 10/29/10 06:10 a-onl End_of_Run_8024
  • 337284 10/29/10 06:09 a-onl End_of_Run_8025
  • 337283 10/29/10 06:09 flay Run 8025 Trigger/ROC Snapshots
  • 337282 10/29/10 06:05 lselvy Target Update
  • 337281 10/29/10 05:15 a-onl Start_Run_8025,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337280 10/29/10 05:11 flay Run 8024 Trigger/ROC Snapshots
  • 337278 10/29/10 04:14 a-onl Start_Run_8024,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337277 10/29/10 03:51 auto_snap Radiation and Ozone snapshot
  • 337276 10/29/10 03:51 auto_snap Miscellaneous snapshot
  • 337275 10/29/10 03:51 auto_snap Hall A Tools snapshot
  • 337274 10/29/10 03:51 auto_snap Compton Polarimeter snapshot
  • 337273 10/29/10 03:51 auto_snap Beam Position and Current snapshot
  • 337272 10/29/10 03:47 a-onl Start_Run_8023,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337271 10/29/10 03:41 a-onl Start_Run_8022,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337270 10/29/10 03:41 a-onl End_of_Run_8021
  • 337269 10/29/10 03:31 flay Run 8021 Trigger/ROC Snapshots
  • 337268 10/29/10 02:57 auto_snap Radiation and Ozone snapshot
  • 337267 10/29/10 02:56 auto_snap Miscellaneous snapshot
  • 337266 10/29/10 02:56 auto_snap Hall A Tools snapshot
  • 337265 10/29/10 02:56 auto_snap Compton Polarimeter snapshot
  • 337264 10/29/10 02:56 auto_snap Beam Position and Current snapshot
  • 337263 10/29/10 02:48 jroche propagated pion rejector map to online display golden run
  • 337262 10/29/10 02:39 a-onl Start_Run_8021,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337261 10/29/10 02:38 a-onl End_of_Run_8020
  • 337260 10/29/10 02:22 flay Run 8020 Trigger/ROC Snapshots
  • 337259 10/29/10 02:05 jroche Calorimeter DC current
  • 337258 10/29/10 01:35 a-onl Start_Run_8020,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337257 10/29/10 01:34 a-onl End_of_Run_8012
  • 337256 10/29/10 00:42 jroche Helicity flip rate restored to 943 Hz
  • 337255 10/29/10 00:39 a-onl Start_Run_8019,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337254 10/29/10 00:34 a-onl End_of_Run_8018
  • 337253 10/29/10 00:15 a-onl Start_Run_8018,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337252 10/29/10 00:14 a-onl End_of_Run_8017
  • 337251 10/29/10 00:13 flay Run 8017 Trigger/ROC Snapshots
  • 337250 10/29/10 00:00 adaq Beam Time Accounting
  • 337249 10/29/10 00:00 M Friend Helicity board set to 120Hz, Compton DAQ set to 1kHz
  • Thursday

  • 337248 10/28/10 23:59 efuchey 10/28/10 Swing shift summary
  • 337247 10/28/10 23:38 cshloka Pion rejector mapping fixed
  • 337246 10/28/10 23:37 LeRose Target Screens
  • 337245 10/28/10 23:12 a-onl Start_Run_8017,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337244 10/28/10 23:10 a-onl End_of_Run_8016
  • 337243 10/28/10 23:08 itard entries end of run
  • 337242 10/28/10 22:52 efuchey Problem with Compton online analyzer
  • 337241 10/28/10 22:11 a-onl Start_Run_8016,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337240 10/28/10 22:09 a-onl End_of_Run_8015
  • 337239 10/28/10 22:08 itard entries end of run
  • 337238 10/28/10 21:08 a-onl Start_Run_8015,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337237 10/28/10 21:06 LeRose Target Screens
  • 337236 10/28/10 21:06 itard entries end of run
  • 337235 10/28/10 21:05 a-onl End_of_Run_8014
  • 337234 10/28/10 20:40 Carlos Deadtime in buffered mode
  • 337233 10/28/10 20:15 a-onl Start_Run_8014,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production
  • 337232 10/28/10 20:07 a-onl End_of_Run_8013
  • 337231 10/28/10 20:06 itard dead time study 4.5 uA
  • 337230 10/28/10 19:57 a-onl Start_Run_8013,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=1,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=dead time study 4.5uA
  • 337229 10/28/10 19:51 a-onl End_of_Run_8007
  • 337228 10/28/10 19:51 auto_snap Radiation and Ozone snapshot
  • 337227 10/28/10 19:51 auto_snap Miscellaneous snapshot
  • 337226 10/28/10 19:51 auto_snap Hall A Tools snapshot
  • 337225 10/28/10 19:51 auto_snap Compton Polarimeter snapshot
  • 337224 10/28/10 19:51 auto_snap Beam Position and Current snapshot
  • 337223 10/28/10 19:50 itard dead time study 3uA
  • 337222 10/28/10 19:41 a-onl Start_Run_8012,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=1,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=dead time study 3uA
  • 337221 10/28/10 19:36 a-onl End_of_Run_8011
  • 337220 10/28/10 19:27 a-onl Start_Run_8011,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=1,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=dead time study 3uA
  • 337219 10/28/10 19:22 a-onl End_of_Run_8010
  • 337218 10/28/10 19:21 itard dead time study 2 uA
  • 337217 10/28/10 19:12 efuchey current change
  • 337216 10/28/10 19:08 a-onl Start_Run_8010,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=1,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=dead time study 2uA
  • 337215 10/28/10 19:08 a-onl End_of_Run_8009
  • 337214 10/28/10 18:56 a-onl Start_Run_8009,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=1,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=dead time study 2uA
  • 337213 10/28/10 18:56 auto_snap Radiation and Ozone snapshot
  • 337212 10/28/10 18:56 auto_snap Miscellaneous snapshot
  • 337211 10/28/10 18:56 auto_snap Hall A Tools snapshot
  • 337210 10/28/10 18:56 auto_snap Compton Polarimeter snapshot
  • 337209 10/28/10 18:56 auto_snap Beam Position and Current snapshot
  • 337208 10/28/10 18:53 LeRose Target Screens
  • 337207 10/28/10 18:53 a-onl End_of_Run_8008
  • 337206 10/28/10 18:53 itard dead time study
  • 337205 10/28/10 18:42 a-onl Start_Run_8008,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=1,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=dead time study 1uA
  • 337204 10/28/10 18:41 a-onl End_of_Run_8003
  • 337203 10/28/10 18:38 efuchey Cavity scan attempt
  • 337202 10/28/10 18:36 a-onl Start_Run_8007,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=1,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=dead time study 1uA
  • 337201 10/28/10 18:27 itard logs of end of run
  • 337200 10/28/10 18:17 a-onl End_of_Run_8006
  • 337199 10/28/10 17:53 Carlos Dead time study with buffering
  • 337198 10/28/10 17:46 a-onl Start_Run_8006,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test sync scalers
  • 337197 10/28/10 17:45 a-onl End_of_Run_8005
  • 337196 10/28/10 17:41 efuchey Compton Rates for 10-28-10
  • 337195 10/28/10 17:29 Carlos Pedestals fluctuations
  • 337194 10/28/10 17:26 a-onl Start_Run_8005,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test sync scalers
  • 337193 10/28/10 17:25 a-onl End_of_Run_8004
  • 337192 10/28/10 17:22 efuchey Hall A tools screenshot
  • 337191 10/28/10 17:20 efuchey Calo base currents
  • 337190 10/28/10 17:17 a-onl Start_Run_8004,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test sync scalers
  • 337189 10/28/10 17:11 a-onl Start_Run_8003,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test sync scalers
  • 337188 10/28/10 17:10 LeRose Target to LH2
  • 337187 10/28/10 16:59 a-onl End_of_Run_8002
  • 337186 10/28/10 16:56 efuchey Harp scan
  • 337185 10/28/10 16:51 efuchey Beam is back
  • 337184 10/28/10 16:34 efuchey Hall in beam permit
  • 337183 10/28/10 16:31 adaq Beam Time Accounting
  • 337182 10/28/10 16:30 A. Camsonne DAQ Sync work
  • 337181 10/28/10 16:15 a-onl Start_Run_8002,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test sync scalers
  • 337180 10/28/10 16:02 a-onl Start_Run_8001,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test buffering E12
  • 337179 10/28/10 16:00 a-onl End_of_Run_8000
  • 337178 10/28/10 15:56 a-onl Start_Run_8000,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test buffering E12
  • 337177 10/28/10 15:54 a-onl End_of_Run_7999
  • 337176 10/28/10 15:44 a-onl Start_Run_7999,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test buffering E12
  • 337175 10/28/10 15:14 a-onl End_of_Run_7998
  • 337174 10/28/10 15:13 a-onl Start_Run_7998,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Sync check work
  • 337173 10/28/10 15:12 a-onl End_of_Run_7996
  • 337172 10/28/10 15:11 a-onl Start_Run_7997,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Sync check work
  • 337171 10/28/10 15:10 a-onl End_of_Run_7995
  • 337170 10/28/10 15:01 A. Camsonne ARS firmware version E12 loaded
  • 337169 10/28/10 15:00 a-onl Start_Run_7996,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Sync check work
  • 337168 10/28/10 15:00 Carlos Pedestal fluctuations
  • 337167 10/28/10 15:00 a-onl End_of_Run_7992
  • 337166 10/28/10 14:56 a-onl Start_Run_7995,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Sync check work
  • 337165 10/28/10 14:54 a-onl End_of_Run_7994
  • 337164 10/28/10 14:41 a-onl Start_Run_7994,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Sync check work
  • 337163 10/28/10 14:41 a-onl End_of_Run_7993
  • 337162 10/28/10 14:38 a-onl Start_Run_7993,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Sync check work
  • 337161 10/28/10 14:27 a-onl Start_Run_7992,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Sync check work
  • 337160 10/28/10 14:24 a-onl End_of_Run_7991
  • 337159 10/28/10 14:23 a-onl Start_Run_7991,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Sync check work
  • 337158 10/28/10 14:17 Fansler Moeller
  • 337157 10/28/10 13:49 adaq Beam Time Accounting
  • 337156 10/28/10 13:46 cshloka target screens snapshot
  • 337155 10/28/10 13:19 Jack Pion Rejector cabling
  • 337154 10/28/10 11:51 auto_snap Radiation and Ozone snapshot
  • 337153 10/28/10 11:51 auto_snap Miscellaneous snapshot
  • 337152 10/28/10 11:51 auto_snap Hall A Tools snapshot
  • 337151 10/28/10 11:51 auto_snap Compton Polarimeter snapshot
  • 337150 10/28/10 11:51 auto_snap Beam Position and Current snapshot
  • 337149 10/28/10 11:50 efuchey QWP scan for 10-28-10
  • 337148 10/28/10 11:49 a-onl End_of_Run_7990
  • 337147 10/28/10 11:15 O.Glamazdin Beam polarization sign
  • 337146 10/28/10 10:57 Fansler/Ewin RQ2 04/16 Valves
  • 337145 10/28/10 10:56 auto_snap Radiation and Ozone snapshot
  • 337144 10/28/10 10:56 auto_snap Miscellaneous snapshot
  • 337143 10/28/10 10:56 auto_snap Hall A Tools snapshot
  • 337142 10/28/10 10:56 auto_snap Compton Polarimeter snapshot
  • 337141 10/28/10 10:56 auto_snap Beam Position and Current snapshot
  • 337140 10/28/10 10:43 A. Camsonne Fastbus ADC vs V792 ADC in DVCS crate
  • 337139 10/28/10 10:19 jroche testing the end of run scripts
  • 337138 10/28/10 10:14 a-onl Start_Run_7990,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Sync check work
  • 337137 10/28/10 10:12 a-onl End_of_Run_7988
  • 337136 10/28/10 10:12 a-onl End_of_Run_7989
  • 337135 10/28/10 10:07 A. Camsonne /adaql4/work1 full removing PREX rootfiles
  • 337134 10/28/10 10:03 a-onl Start_Run_7989,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Sync check work
  • 337133 10/28/10 10:03 a-onl End_of_Run_7988
  • 337132 10/28/10 09:52 a-onl Start_Run_7988,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Sync check work
  • 337131 10/28/10 09:40 cshloka target status and L2 cernox fluctuations
  • 337130 10/28/10 09:39 a-onl End_of_Run_7987
  • 337129 10/28/10 09:38 a-onl End_of_Run_7987
  • 337128 10/28/10 09:37 a-onl End_of_Run_7987
  • 337127 10/28/10 07:25 lselvy Target Move
  • 337126 10/28/10 07:23 Hisham Controlled access
  • 337125 10/28/10 07:15 Hisham Owl shift summary
  • 337124 10/28/10 07:14 flay Run 7987 is a Junk Run
  • 337123 10/28/10 07:10 flay Run 7986 is a Junk Run
  • 337122 10/28/10 07:07 flay Run 7986 is a Junk Run
  • 337121 10/28/10 07:07 a-onl Start_Run_7987,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin III - 4 pass
  • 337120 10/28/10 07:04 a-onl Start_Run_7986,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin III - 4 pass
  • 337119 10/28/10 07:03 a-onl End_of_Run_7984
  • 337118 10/28/10 07:02 flay Run 7985 Trigger/ROC Snapshots
  • 337117 10/28/10 06:05 lselvy Target Update
  • 337116 10/28/10 06:02 a-onl Start_Run_7985,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin III - 4 pass
  • 337115 10/28/10 06:02 a-onl End_of_Run_7983
  • 337114 10/28/10 06:01 flay Run 7984 Trigger/ROC Snapshot
  • 337113 10/28/10 05:03 flay Run 7984 Trigger/ROC Snapshots
  • 337112 10/28/10 05:02 a-onl Start_Run_7984,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin III - 4 pass
  • 337111 10/28/10 05:01 a-onl End_of_Run_7975
  • 337110 10/28/10 04:01 flay Run 7983 Trigger/ROC Snapshots
  • 337109 10/28/10 04:00 a-onl Start_Run_7983,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin III - 4 pass
  • 337108 10/28/10 03:59 a-onl End_of_Run_7982
  • 337107 10/28/10 03:50 auto_snap Radiation and Ozone snapshot
  • 337106 10/28/10 03:50 auto_snap Miscellaneous snapshot
  • 337105 10/28/10 03:50 auto_snap Hall A Tools snapshot
  • 337104 10/28/10 03:50 auto_snap Compton Polarimeter snapshot
  • 337103 10/28/10 03:50 auto_snap Beam Position and Current snapshot
  • 337102 10/28/10 03:02 flay Run 7982 Trigger/ROC Snapshots
  • 337101 10/28/10 03:02 a-onl Start_Run_7982,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin III - 4 pass
  • 337100 10/28/10 03:01 a-onl End_of_Run_7981
  • 337099 10/28/10 02:56 auto_snap Radiation and Ozone snapshot
  • 337098 10/28/10 02:56 auto_snap Miscellaneous snapshot
  • 337097 10/28/10 02:56 auto_snap Hall A Tools snapshot
  • 337096 10/28/10 02:55 auto_snap Compton Polarimeter snapshot
  • 337095 10/28/10 02:55 auto_snap Beam Position and Current snapshot
  • 337094 10/28/10 02:04 a-onl Start_Run_7981,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin III - 4 pass
  • 337093 10/28/10 02:03 flay Run 7981 Trigger/ROC Snapshots
  • 337092 10/28/10 02:02 a-onl End_of_Run_7980
  • 337091 10/28/10 01:52 hisham Hall A General Tools
  • 337090 10/28/10 01:50 hisham PbF2 Calo ADC Readback
  • 337089 10/28/10 01:46 Hisham dvcstlab1.2006 Alarm
  • 337088 10/28/10 01:02 flay Run 7980 Trigger/ROC Snapshots
  • 337087 10/28/10 01:01 a-onl Start_Run_7980,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin III - 4 pass
  • 337086 10/28/10 01:00 a-onl End_of_Run_7979
  • 337085 10/28/10 00:01 adaq Beam Time Accounting
  • Wednesday

  • 337084 10/27/10 23:59 a-onl Start_Run_7979,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin III - 4 pass
  • 337083 10/27/10 23:59 A. Camsonne Readout of ROC18 and trigger library updated
  • 337082 10/27/10 23:59 Yi Qiang Swing Shift Summary
  • 337081 10/27/10 23:56 a-onl End_of_Run_7978
  • 337080 10/27/10 23:40 mcanan swing_target_screens
  • 337079 10/27/10 23:37 a-onl Start_Run_7978,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin III - 4 pass
  • 337078 10/27/10 23:36 a-onl End_of_Run_7977
  • 337077 10/27/10 23:34 Punjabi roc screens
  • 337076 10/27/10 23:32 Punjabi scaler, RC screens
  • 337075 10/27/10 23:16 Yi Qiang Calorimeter current
  • 337074 10/27/10 23:16 efuchey Rates became too high
  • 337073 10/27/10 22:51 a-onl Start_Run_7977,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin III - 4 pass
  • 337072 10/27/10 22:39 Yi Qiang Problem solved, back to production
  • 337071 10/27/10 22:24 Yi Qiang Beam will be off for a while
  • 337070 10/27/10 22:19 Yi Qiang Beam energy EPICS value missing at the end of run 7975
  • 337069 10/27/10 22:12 a-onl End_of_Run_7976
  • 337068 10/27/10 22:09 a-onl Start_Run_7976,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin III - 4 pass
  • 337067 10/27/10 22:07 a-onl End_of_Run_7971
  • 337066 10/27/10 22:00 A. Camsonne Left HRS efficiency run Run 7975
  • 337065 10/27/10 21:59 a-onl Start_Run_7975,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Left HRS single run - Efficiency check run
  • 337064 10/27/10 21:57 a-onl End_of_Run_7974
  • 337063 10/27/10 21:55 Punjabi roc screens
  • 337062 10/27/10 21:54 Punjabi scaler and RC screens
  • 337061 10/27/10 21:19 efuchey Compton CODA crash
  • 337060 10/27/10 21:13 a-onl Start_Run_7974,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3, 4 pass beam
  • 337059 10/27/10 21:12 a-onl End_of_Run_7973
  • 337058 10/27/10 21:11 Punjabi roc screens
  • 337057 10/27/10 21:11 Punjabi RC, Scaler screens
  • 337056 10/27/10 20:31 a-onl Start_Run_7973,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3, 4 pass beam
  • 337055 10/27/10 20:27 a-onl End_of_Run_7972
  • 337054 10/27/10 20:26 Punjabi Roc screens
  • 337053 10/27/10 20:25 Punjabi rc, scaler screens
  • 337052 10/27/10 19:50 auto_snap Radiation and Ozone snapshot
  • 337051 10/27/10 19:50 auto_snap Miscellaneous snapshot
  • 337050 10/27/10 19:50 auto_snap Hall A Tools snapshot
  • 337049 10/27/10 19:50 auto_snap Compton Polarimeter snapshot
  • 337048 10/27/10 19:50 auto_snap Beam Position and Current snapshot
  • 337047 10/27/10 19:46 a-onl Start_Run_7972,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3, 4 pass beam
  • 337046 10/27/10 19:41 a-onl End_of_Run_7970
  • 337045 10/27/10 19:21 a-onl Start_Run_7971,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3, 4 pass beam
  • 337044 10/27/10 19:21 a-onl Start_Run_7969,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - version E10 on 12 boards
  • 337043 10/27/10 19:18 Punjabi roc screens
  • 337042 10/27/10 19:16 Punjabi roc 14 and roc 17 screens
  • 337041 10/27/10 19:14 Punjabi RC and scaler screens
  • 337040 10/27/10 19:01 efuchey Compton rates for 10-27-10
  • 337039 10/27/10 18:55 auto_snap Radiation and Ozone snapshot
  • 337038 10/27/10 18:55 auto_snap Miscellaneous snapshot
  • 337037 10/27/10 18:55 auto_snap Hall A Tools snapshot
  • 337036 10/27/10 18:55 auto_snap Compton Polarimeter snapshot
  • 337035 10/27/10 18:55 auto_snap Beam Position and Current snapshot
  • 337034 10/27/10 18:31 a-onl Start_Run_7970,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3, 4 pass beam
  • 337033 10/27/10 18:27 mcanan Target Motion
  • 337032 10/27/10 18:00 mcanan Target Motion
  • 337031 10/27/10 17:16 Yi Qiang No beam for a while due to magnet trips
  • 337030 10/27/10 17:00 A. Camsonne ARS firmware version E10 downloaded in 12 ARS boards
  • 337029 10/27/10 17:00 a-onl End_of_Run_7969
  • 337028 10/27/10 16:44 Yi Qiang Beam back to the Hall, continue production
  • 337027 10/27/10 16:38 mcanan Target Motion
  • 337026 10/27/10 16:01 adaq Beam Time Accounting
  • 337025 10/27/10 15:59 Jixie Moller has been refilled
  • 337024 10/27/10 15:57 Saha Shift Summary
  • 337023 10/27/10 15:55 O.Glamazdin Moller results
  • 337022 10/27/10 15:45 Lamiaa Target's Monitoring tools
  • 337021 10/27/10 15:42 Missing proton mass
  • 337020 10/27/10 14:15 a-onl Start_Run_7968,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ
  • 337019 10/27/10 13:33 A. Camsonne Run7967 Junk run
  • 337018 10/27/10 13:24 a-onl Start_Run_7967,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Firmware E11 - With buffering ~100 Hz Interrupt mode
  • 337017 10/27/10 13:19 a-onl End_of_Run_7966
  • 337016 10/27/10 13:18 a-onl Start_Run_7966,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Firmware E11 - With buffering ~100 Hz Interrupt mode
  • 337015 10/27/10 13:09 a-onl End_of_Run_7965
  • 337014 10/27/10 13:06 A. Camsonne Loaded tsUtil.o on TS11 crate
  • 337013 10/27/10 12:48 a-onl Start_Run_7965,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Firmware E11 - With buffering ~100 Hz Interrupt mode
  • 337012 10/27/10 12:45 a-onl End_of_Run_7964
  • 337011 10/27/10 12:39 a-onl Start_Run_7964,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Firmware E11 - With buffering ~100 Hz Interrupt mode
  • 337010 10/27/10 12:39 A. Camsonne Run 7964 run in interrupt mode
  • 337009 10/27/10 12:33 a-onl End_of_Run_7963
  • 337008 10/27/10 12:31 a-onl Start_Run_7963,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Firmware E11 - With buffering ~100 Hz
  • 337007 10/27/10 12:24 a-onl End_of_Run_7962
  • 337006 10/27/10 12:22 A. Camsonne Empty FIFO messages Run 7692 with buffering
  • 337005 10/27/10 12:19 a-onl Start_Run_7962,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Firmware E11 - With buffering ~200 Hz
  • 337004 10/27/10 12:15 A. Camsonne ARS Firmware E11 loaded for buffering tests - HV on cosmics paddles on
  • 337003 10/27/10 12:06 a-onl End_of_Run_7961
  • 337002 10/27/10 11:54 a-onl Start_Run_7961,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Firmware E10
  • 337001 10/27/10 11:51 A. Camsonne DVCS calorimeter HV off for Moller measurement
  • 337000 10/27/10 11:50 auto_snap Radiation and Ozone snapshot
  • 336999 10/27/10 11:50 auto_snap Miscellaneous snapshot
  • 336998 10/27/10 11:50 auto_snap Hall A Tools snapshot
  • 336997 10/27/10 11:50 auto_snap Compton Polarimeter snapshot
  • 336996 10/27/10 11:50 auto_snap Beam Position and Current snapshot
  • 336995 10/27/10 10:55 auto_snap Radiation and Ozone snapshot
  • 336994 10/27/10 10:55 auto_snap Miscellaneous snapshot
  • 336993 10/27/10 10:55 auto_snap Hall A Tools snapshot
  • 336992 10/27/10 10:55 auto_snap Compton Polarimeter snapshot
  • 336991 10/27/10 10:55 auto_snap Beam Position and Current snapshot
  • 336990 10/27/10 10:39 O.Glamazdin Moller magnet is rumping up
  • 336989 10/27/10 10:14 efuchey shift check list modified
  • 336988 10/27/10 09:58 gaskell Run plan for DAY and SWING shift
  • 336987 10/27/10 09:50 efuchey QWP scan for 10-27-10
  • 336986 10/27/10 09:26 O.Glamazdin Controlled acces for the Moller solenoid reboot
  • 336985 10/27/10 09:17 Lamiaa Strange behavior of Hal2_Cernox_2_R
  • 336984 10/27/10 08:44 Lamiaa Screenshot of target's charts
  • 336983 10/27/10 08:34 Lamiaa Move target to Empty position
  • 336982 10/27/10 08:27 a-onl End_of_Run_7960
  • 336981 10/27/10 08:11 jroche Owl shift summary
  • 336980 10/27/10 08:09 adaq Beam Time Accounting
  • 336979 10/27/10 08:02 a-onl Start_Run_7960,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production kin3 low - ARS firmware E10 for 12 boards
  • 336978 10/27/10 08:00 a-onl End_of_Run_7959
  • 336977 10/27/10 07:54 flay Run 7959 Trigger/ROC Snapshot
  • 336976 10/27/10 07:17 a-onl Start_Run_7959,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production kin3 low - ARS firmware E10 for 12 boards
  • 336975 10/27/10 07:11 flay Run 7958 is a Junk Run
  • 336974 10/27/10 07:11 a-onl End_of_Run_7958
  • 336973 10/27/10 07:11 a-onl Start_Run_7958,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production kin3 low - ARS firmware E10 for 12 boards
  • 336972 10/27/10 07:09 flay Run 7957 is a Junk Run
  • 336971 10/27/10 07:09 a-onl End_of_Run_7957
  • 336970 10/27/10 07:04 a-onl Start_Run_7957,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production kin3 low - ARS firmware E10 for 12 boards
  • 336969 10/27/10 07:03 a-onl End_of_Run_7956
  • 336968 10/27/10 06:18 flay Run 7956 Trigger/ROC Snapshot
  • 336967 10/27/10 06:09 a-onl Start_Run_7956,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production kin3 low - ARS firmware E10 for 12 boards
  • 336966 10/27/10 06:08 a-onl End_of_Run_7955
  • 336965 10/27/10 05:51 Carlos Pedestal issues with new ARS firmware (E11)
  • 336964 10/27/10 05:47 jroche Increased compton HV to 1675V
  • 336963 10/27/10 05:46 jroche It's freezing in here !
  • 336962 10/27/10 05:10 a-onl Start_Run_7955,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production kin3 low - ARS firmware E10 for 12 boards
  • 336961 10/27/10 05:06 a-onl End_of_Run_7954
  • 336960 10/27/10 05:06 a-onl Start_Run_7954,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production kin3 low - ARS firmware E10 for 12 boards
  • 336959 10/27/10 05:04 A. Camsonne ARS firmware version E10 downloaded in 12 ARS boards
  • 336958 10/27/10 05:01 a-onl End_of_Run_7953
  • 336957 10/27/10 04:59 flay Some CALO Blocks Not Working
  • 336956 10/27/10 04:48 A. Camsonne DVCS sync check work
  • 336955 10/27/10 04:39 a-onl Start_Run_7953,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production kin3 low
  • 336954 10/27/10 04:34 a-onl Start_Run_7952,Run_type=Pedestal,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Pedestals
  • 336953 10/27/10 04:33 a-onl End_of_Run_7951
  • 336952 10/27/10 04:23 a-onl Start_Run_7951,Run_type=Pedestal,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production kin3 low
  • 336951 10/27/10 04:16 a-onl End_of_Run_7950
  • 336950 10/27/10 04:15 a-onl Start_Run_7950,Run_type=Pedestal,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Pedestal run
  • 336949 10/27/10 04:14 a-onl End_of_Run_7949
  • 336948 10/27/10 04:00 a-onl Start_Run_7949,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kinematic 3 -- Low Energy [production]
  • 336947 10/27/10 03:50 auto_snap Radiation and Ozone snapshot
  • 336946 10/27/10 03:50 auto_snap Miscellaneous snapshot
  • 336945 10/27/10 03:49 auto_snap Hall A Tools snapshot
  • 336944 10/27/10 03:49 auto_snap Compton Polarimeter snapshot
  • 336943 10/27/10 03:49 auto_snap Beam Position and Current snapshot
  • 336942 10/27/10 03:46 jroche Compton rates look reasonnable
  • 336941 10/27/10 03:46 itard target graph
  • 336940 10/27/10 03:46 a-onl Start_Run_7947,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kinematic 3 -- Low Energy [production]
  • 336939 10/27/10 03:28 jroche Calo anode currents
  • 336938 10/27/10 03:25 a-onl Start_Run_7946,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kinematic 3 -- Low Energy [production]
  • 336937 10/27/10 03:23 jroche general Tools at the beginning of the data taking
  • 336936 10/27/10 03:20 jroche harps looks good
  • 336935 10/27/10 03:02 a-onl End_of_Run_7945
  • 336934 10/27/10 03:02 jroche data taking to resume shortly
  • 336933 10/27/10 02:55 auto_snap Radiation and Ozone snapshot
  • 336932 10/27/10 02:55 auto_snap Miscellaneous snapshot
  • 336931 10/27/10 02:55 auto_snap Hall A Tools snapshot
  • 336930 10/27/10 02:55 auto_snap Compton Polarimeter snapshot
  • 336929 10/27/10 02:55 auto_snap Beam Position and Current snapshot
  • 336928 10/27/10 02:35 a-onl Start_Run_7945,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check - sync signal on ROC4 slot 22 chan 48
  • 336927 10/27/10 02:27 a-onl Start_Run_7944,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check - sync signal on ROC4 slot 22 chan 48
  • 336926 10/27/10 02:21 a-onl Start_Run_7943,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check - sync signal on ROC4 slot 22 chan 48
  • 336925 10/27/10 02:20 a-onl End_of_Run_7942
  • 336924 10/27/10 02:15 a-onl Start_Run_7942,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check - sync signal on ROC4 slot 22 chan 48
  • 336923 10/27/10 02:13 a-onl End_of_Run_7939
  • 336922 10/27/10 01:59 a-onl Start_Run_7941,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check - sync signal on ROC4 slot 22 chan 48
  • 336921 10/27/10 01:58 a-onl End_of_Run_7940
  • 336920 10/27/10 01:55 a-onl Start_Run_7940,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check - sync signal on ROC4 slot 22 chan 48
  • 336919 10/27/10 01:54 a-onl End_of_Run_7938
  • 336918 10/27/10 01:50 a-onl Start_Run_7939,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check - sync signal on ROC4 slot 22 chan 48
  • 336917 10/27/10 01:48 a-onl End_of_Run_7936
  • 336916 10/27/10 01:39 a-onl Start_Run_7938,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check - sync signal on ROC4 slot 22 chan 48
  • 336915 10/27/10 01:37 a-onl End_of_Run_7937
  • 336914 10/27/10 01:33 a-onl Start_Run_7937,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check - sync signal on ROC4 slot 22 chan 48
  • 336913 10/27/10 01:26 a-onl Start_Run_7936,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check - sync signal on ROC4 slot 22 chan 48
  • 336912 10/27/10 01:25 a-onl End_of_Run_7935
  • 336911 10/27/10 01:24 a-onl End_of_Run_7933
  • 336910 10/27/10 01:04 a-onl Start_Run_7935,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336909 10/27/10 01:03 a-onl End_of_Run_7934
  • 336908 10/27/10 00:49 a-onl Start_Run_7934,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336907 10/27/10 00:48 a-onl End_of_Run_7930
  • 336906 10/27/10 00:34 a-onl Start_Run_7933,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336905 10/27/10 00:34 a-onl End_of_Run_7932
  • 336904 10/27/10 00:25 a-onl Start_Run_7932,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336903 10/27/10 00:24 a-onl End_of_Run_7931
  • 336902 10/27/10 00:15 a-onl Start_Run_7931,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336901 10/27/10 00:14 a-onl End_of_Run_7928
  • 336900 10/27/10 00:02 a-onl Start_Run_7930,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336899 10/27/10 00:01 a-onl End_of_Run_7929
  • 336897 10/27/10 00:01 GOLGE Swing Shift Summary
  • 336898 10/27/10 00:00 adaq Beam Time Accounting
  • Tuesday

  • 336896 10/26/10 23:44 a-onl Start_Run_7929,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336895 10/26/10 23:41 a-onl Start_Run_7928,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336894 10/26/10 23:40 a-onl End_of_Run_7925
  • 336892 10/26/10 23:36 a-onl Start_Run_7927,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336891 10/26/10 23:35 a-onl End_of_Run_7926
  • 336890 10/26/10 23:33 a-onl Start_Run_7926,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336889 10/26/10 23:32 a-onl End_of_Run_7924
  • 336888 10/26/10 23:27 a-onl Start_Run_7925,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336887 10/26/10 23:19 a-onl Start_Run_7924,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336886 10/26/10 22:56 a-onl End_of_Run_7923
  • 336885 10/26/10 22:55 a-onl Start_Run_7923,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS sync ADC check
  • 336884 10/26/10 22:54 a-onl End_of_Run_7922
  • 336883 10/26/10 22:35 a-onl Start_Run_7922,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=no beam
  • 336882 10/26/10 22:34 a-onl End_of_Run_7921
  • 336881 10/26/10 22:13 a-onl Start_Run_7921,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=no beam
  • 336880 10/26/10 22:07 Golge Controlled access
  • 336879 10/26/10 22:03 a-onl Start_Run_7920,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=no beam
  • 336878 10/26/10 21:57 a-onl End_of_Run_7918
  • 336877 10/26/10 21:54 a-onl Start_Run_7918,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=no beam
  • 336876 10/26/10 21:30 a-onl End_of_Run_7917
  • 336875 10/26/10 20:33 A. Camsonne Networked FPGA programmer replugged
  • 336874 10/26/10 20:22 a-onl Start_Run_7917,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=no beam
  • 336873 10/26/10 20:09 a-onl Start_Run_7916,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=no beam
  • 336872 10/26/10 19:49 auto_snap Radiation and Ozone snapshot
  • 336871 10/26/10 19:49 auto_snap Miscellaneous snapshot
  • 336870 10/26/10 19:49 auto_snap Hall A Tools snapshot
  • 336869 10/26/10 19:49 auto_snap Compton Polarimeter snapshot
  • 336868 10/26/10 19:49 auto_snap Beam Position and Current snapshot
  • 336867 10/26/10 19:25 a-onl End_of_Run_7915
  • 336866 10/26/10 19:23 a-onl Start_Run_7915,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin III 4 pass - 2 uA
  • 336865 10/26/10 19:20 gaskell Run plan for Tuesday Swing --> Wednesday Day
  • 336864 10/26/10 19:19 a-onl Start_Run_7914,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin III 4 pass
  • 336863 10/26/10 19:19 a-onl Start_Run_7913,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DAQ check
  • 336862 10/26/10 19:14 a-onl End_of_Run_7913
  • 336861 10/26/10 19:11 a-onl End_of_Run_7912
  • 336860 10/26/10 19:07 a-onl Start_Run_7912,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3 only run
  • 336859 10/26/10 19:06 a-onl End_of_Run_7911
  • 336858 10/26/10 19:05 a-onl End_of_Run_7908
  • 336857 10/26/10 18:54 auto_snap Radiation and Ozone snapshot
  • 336856 10/26/10 18:54 auto_snap Miscellaneous snapshot
  • 336855 10/26/10 18:54 auto_snap Hall A Tools snapshot
  • 336854 10/26/10 18:54 auto_snap Compton Polarimeter snapshot
  • 336853 10/26/10 18:54 auto_snap Beam Position and Current snapshot
  • 336852 10/26/10 18:53 Magali news trigger fpga ctrl firmware
  • 336851 10/26/10 18:47 a-onl Start_Run_7911,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production run
  • 336850 10/26/10 18:46 A. Camsonne ROC18 power cycled
  • 336849 10/26/10 18:43 a-onl End_of_Run_7910
  • 336848 10/26/10 18:39 a-onl Start_Run_7910,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production run
  • 336847 10/26/10 18:38 a-onl End_of_Run_7909
  • 336846 10/26/10 18:36 efuchey Compton rates for 10-26-10
  • 336845 10/26/10 18:31 a-onl Start_Run_7909,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Buffering test
  • 336844 10/26/10 18:31 a-onl End_of_Run_7906
  • 336843 10/26/10 18:29 a-onl Start_Run_7908,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Buffering test
  • 336842 10/26/10 18:28 a-onl End_of_Run_7907
  • 336841 10/26/10 18:25 a-onl Start_Run_7907,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Buffering test
  • 336840 10/26/10 18:19 a-onl Start_Run_7906,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Buffering test
  • 336839 10/26/10 18:16 a-onl End_of_Run_7905
  • 336838 10/26/10 18:14 jroche BCM calibration: some result.
  • 336837 10/26/10 18:06 a-onl Start_Run_7905,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production run
  • 336836 10/26/10 18:01 mcanan Target Motion
  • 336835 10/26/10 17:44 mcanan Target Motion
  • 336834 10/26/10 17:42 A. Camsonne DVCS calorimeter cosmics paddle off
  • 336833 10/26/10 17:15 A. Camsonne Target operator refresher by JP on DVCS wiki
  • 336832 10/26/10 16:24 A. Camsonne APC src2 channel assignement
  • 336831 10/26/10 16:19 A. Camsonne iochacs serial port on dvcstlab1 2008
  • 336829 10/26/10 16:03 Holmstrom Day Shift Summary
  • 336830 10/26/10 16:03 adaq Beam Time Accounting
  • 336828 10/26/10 16:02 A. Camsonne Slow control APC PDU switched to src2 hostname
  • 336827 10/26/10 15:48 A. Camsonne Calorimeter cosmics paddles HV on for buffering test
  • 336826 10/26/10 15:45 a-onl Start_Run_7904,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=test ROC 17 24 with buffering version E11
  • 336825 10/26/10 15:42 A. Camsonne ARS Firware version E11
  • 336824 10/26/10 15:40 a-onl Start_Run_7903,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=test ROC 17 24 version E11
  • 336823 10/26/10 15:24 a-onl Start_Run_7901,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=test ROC 17 E11
  • 336822 10/26/10 15:06 a-onl End_of_Run_7899
  • 336821 10/26/10 15:01 a-onl Start_Run_7900,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=test ROC 17 E11
  • 336820 10/26/10 14:58 a-onl Start_Run_7899,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=test ROC 17 E11
  • 336819 10/26/10 14:22 Jixie Moller magnet has been filled
  • 336818 10/26/10 13:19 Carlos HRS moved to new angle (31.82 deg: kin3 low)
  • 336817 10/26/10 12:27 hyde S0 Efficiency
  • 336816 10/26/10 11:49 auto_snap Radiation and Ozone snapshot
  • 336815 10/26/10 11:49 auto_snap Miscellaneous snapshot
  • 336814 10/26/10 11:49 auto_snap Hall A Tools snapshot
  • 336813 10/26/10 11:49 auto_snap Compton Polarimeter snapshot
  • 336812 10/26/10 11:49 auto_snap Beam Position and Current snapshot
  • 336811 10/26/10 11:46 a-onl Start_Run_7898,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=test
  • 336810 10/26/10 11:27 efuchey Compton beamline
  • 336809 10/26/10 11:12 Fansler/Ewin Left Dipole
  • 336808 10/26/10 11:01 a-onl Start_Run_7897,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=test
  • 336807 10/26/10 10:54 auto_snap Radiation and Ozone snapshot
  • 336806 10/26/10 10:54 auto_snap Miscellaneous snapshot
  • 336805 10/26/10 10:54 auto_snap Hall A Tools snapshot
  • 336804 10/26/10 10:54 auto_snap Compton Polarimeter snapshot
  • 336803 10/26/10 10:54 auto_snap Beam Position and Current snapshot
  • 336802 10/26/10 10:36 a-onl Start_Run_7896,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=test
  • 336801 10/26/10 10:32 gaskell Status and plan for the day
  • 336800 10/26/10 10:29 gaskell HRS work in progress
  • 336799 10/26/10 10:07 Holmstrom Spectrometer Off
  • 336798 10/26/10 09:14 gomez Re: Problem cycling quadrupole current
  • 336797 10/26/10 08:28 a-onl End_of_Run_7895
  • 336796 10/26/10 08:19 Hisham Run 7895: ROC 18 and 14
  • 336795 10/26/10 08:17 Hisham Run 7895 scalers, RunControl gui
  • 336794 10/26/10 08:11 Hisham Hall in controlled access
  • 336793 10/26/10 08:09 Hisham Owl shift summary
  • 336792 10/26/10 08:00 adaq Beam Time Accounting
  • 336791 10/26/10 07:52 a-onl Start_Run_7895,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2c conditions
  • 336790 10/26/10 07:46 cshloka target status
  • 336789 10/26/10 07:34 flay Run 7894 is Junk
  • 336788 10/26/10 07:31 a-onl End_of_Run_7893
  • 336787 10/26/10 07:24 a-onl Start_Run_7894,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2c conditions
  • 336786 10/26/10 07:06 a-onl End_of_Run_7892
  • 336785 10/26/10 07:02 flay Run 7893 Trigger/ROC Snapshots
  • 336784 10/26/10 06:17 a-onl Start_Run_7893,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2c conditions
  • 336783 10/26/10 06:15 flay Run 7892 Physics Quantities
  • 336782 10/26/10 06:12 flay Run 7892 Calo TDCs
  • 336781 10/26/10 06:09 flay Run 7892 LHRS Gas Cherenkov
  • 336780 10/26/10 06:07 flay Run 7892 S1 TDCs
  • 336779 10/26/10 05:57 flay Run 7892 Trigger/ROC Snapshots
  • 336778 10/26/10 05:13 cshloka target status and L2 cernox
  • 336777 10/26/10 05:12 a-onl Start_Run_7892,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2c conditions
  • 336776 10/26/10 04:50 flay Run 7891 is Junk
  • 336775 10/26/10 04:43 Hisham Hall A General Tools snapshot.
  • 336774 10/26/10 04:37 flay Run 7890 LHRS GC
  • 336773 10/26/10 04:32 a-onl End_of_Run_7890
  • 336772 10/26/10 04:13 flay Run 7890 Trigger/Snapshots
  • 336771 10/26/10 03:49 auto_snap Radiation and Ozone snapshot
  • 336770 10/26/10 03:49 auto_snap Miscellaneous snapshot
  • 336769 10/26/10 03:49 auto_snap Hall A Tools snapshot
  • 336768 10/26/10 03:49 auto_snap Compton Polarimeter snapshot
  • 336767 10/26/10 03:48 auto_snap Beam Position and Current snapshot
  • 336766 10/26/10 03:35 flay Run 7889 Exhibits Same Behavior as Run 7888
  • 336765 10/26/10 03:28 a-onl Start_Run_7890,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.131,leftarm_theta=37.86,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2c comdotions
  • 336764 10/26/10 03:20 Hisham Hall A General Tools snapshot
  • 336763 10/26/10 03:14 flay Run 7889: GC Sum
  • 336762 10/26/10 03:10 a-onl End_of_Run_7889
  • 336761 10/26/10 02:59 flay Run 7889 Trigger/ROC Snapshots
  • 336760 10/26/10 02:54 auto_snap Radiation and Ozone snapshot
  • 336759 10/26/10 02:54 auto_snap Miscellaneous snapshot
  • 336758 10/26/10 02:54 auto_snap Hall A Tools snapshot
  • 336757 10/26/10 02:54 auto_snap Compton Polarimeter snapshot
  • 336756 10/26/10 02:54 auto_snap Beam Position and Current snapshot
  • 336755 10/26/10 02:34 Carlos Calorimeter HV update
  • 336754 10/26/10 02:29 Hisham Problems in ycling quadrupole currents
  • 336753 10/26/10 02:29 a-onl Start_Run_7889,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.131,leftarm_theta=37.86,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2c comdotions
  • 336752 10/26/10 02:09 cshloka target status
  • 336751 10/26/10 01:53 flay Important: Database Updated
  • 336750 10/26/10 01:17 Hisham LHRS moved to 37.86 deg
  • 336749 10/26/10 01:11 flay Run 7888 Physics Quantities
  • 336748 10/26/10 01:08 flay Run 7888 Calorimeter TDCs
  • 336747 10/26/10 01:01 flay Run 7888
  • 336746 10/26/10 00:40 a-onl End_of_Run_7888
  • 336745 10/26/10 00:14 flay Run 7888 Trigger/ROC Snapshot
  • 336744 10/26/10 00:05 Min Huang target charts screen
  • 336742 10/26/10 00:03 BoZhao Swing shift summary
  • 336743 10/26/10 00:03 adaq Beam Time Accounting
  • Monday

  • 336741 10/25/10 23:49 a-onl Start_Run_7888,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.00,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2c comdotions
  • 336740 10/25/10 23:25 Margaziotis Run 7887 ROC18 screen
  • 336739 10/25/10 23:24 Margaziotis Run 7887 ROC14 and ROC17 screens
  • 336738 10/25/10 23:23 Margaziotis Run 7887 CODA and scaler screens
  • 336737 10/25/10 23:08 a-onl Start_Run_7887,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.00,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2c comdotions Config trig DVCS
  • 336736 10/25/10 22:51 Min Huang target moved to loop2, 15cm
  • 336735 10/25/10 22:39 BoZhao 4A and 4B Harp scan
  • 336734 10/25/10 21:59 a-onl End_of_Run_7886
  • 336733 10/25/10 21:41 a-onl End_of_Run_7885
  • 336732 10/25/10 21:10 a-onl Start_Run_7886,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.00,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Config DVCS to bot - no beam - test run
  • 336731 10/25/10 21:07 Magali new firmware version
  • 336730 10/25/10 21:06 efuchey Compton tuning
  • 336729 10/25/10 20:50 a-onl Start_Run_7885,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.00,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2c condition
  • 336728 10/25/10 20:48 a-onl End_of_Run_7883
  • 336727 10/25/10 19:55 a-onl End_of_Run_7884
  • 336726 10/25/10 19:52 BoZhao Stop taking data for compton tuning
  • 336725 10/25/10 19:48 auto_snap Radiation and Ozone snapshot
  • 336724 10/25/10 19:48 auto_snap Miscellaneous snapshot
  • 336723 10/25/10 19:48 auto_snap Hall A Tools snapshot
  • 336722 10/25/10 19:48 auto_snap Compton Polarimeter snapshot
  • 336721 10/25/10 19:48 auto_snap Beam Position and Current snapshot
  • 336720 10/25/10 19:47 Min Huang move to empty target
  • 336719 10/25/10 19:46 efuchey Compton rates at another volage
  • 336718 10/25/10 19:40 Margaziotis Run 7884 ROC18 screen
  • 336717 10/25/10 19:39 Margaziotis Run 7884 ROC14 and ROC17
  • 336716 10/25/10 19:38 Margaziotis Run 7884 CODA and scaler screens
  • 336715 10/25/10 19:31 a-onl Start_Run_7884,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.00,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2c condition
  • 336714 10/25/10 19:25 efuchey Compton rates 10-25-10
  • 336713 10/25/10 19:11 a-onl End_of_Run_7881
  • 336712 10/25/10 19:00 Margaziotis Run 7883 ROC18 screen
  • 336711 10/25/10 19:00 Margaziotis Run 7883 ROC14 and ROC17 screens
  • 336710 10/25/10 18:59 Margaziotis Run 7883 CODA and scalers screen
  • 336709 10/25/10 18:43 a-onl Start_Run_7883,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.00,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2c condition
  • 336708 10/25/10 18:22 a-onl End_of_Run_7882
  • 336707 10/25/10 18:16 Margaziotis Run 7882 ROC 18 screen
  • 336706 10/25/10 18:15 Margaziotis Run 7882 ROC14 and ROC17 screens
  • 336705 10/25/10 18:14 Margaziotis Run 7882 CODA and scaler screens
  • 336704 10/25/10 17:52 a-onl Start_Run_7882,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.00,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2c condition
  • 336703 10/25/10 17:31 BoZhao Hall A general tools screen
  • 336702 10/25/10 17:27 Jack (Techs) DVCS grounding strap is too short
  • 336701 10/25/10 16:38 a-onl End_of_Run_7880
  • 336700 10/25/10 16:24 A. Camsonne Calorimeter ADCs channels in end of run time outs
  • 336699 10/25/10 16:21 gaskell Run plan for Monday Day and Swing
  • 336697 10/25/10 16:08 JPChen Day shit summary
  • 336698 10/25/10 16:08 adaq Beam Time Accounting
  • 336696 10/25/10 16:04 a-onl Start_Run_7881,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.131,leftarm_theta=37.86,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2b condition
  • 336695 10/25/10 15:51 a-onl End_of_Run_7879
  • 336694 10/25/10 15:30 O.Glamazdin Run 7880: ROC1
  • 336693 10/25/10 15:30 O.Glamazdin Run 7880: ROC14 and ROC17
  • 336692 10/25/10 15:29 O.Glamazdin Run 7880: CODA and Scalers
  • 336691 10/25/10 15:16 R. Michaels added roc14,17 scalers to datastream
  • 336690 10/25/10 15:14 a-onl Start_Run_7880,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.131,leftarm_theta=37.86,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2b condition
  • 336689 10/25/10 15:07 R. Michaels xscaler updated, can display ROC14, 17
  • 336688 10/25/10 14:59 jpchen HRSL to elastic 2b
  • 336687 10/25/10 14:46 a-onl End_of_Run_7878
  • 336686 10/25/10 14:34 O.Glamazdin Run 7879: ROC18
  • 336685 10/25/10 14:33 O.Glamazdin Run 7879: ROC14 and ROC17
  • 336684 10/25/10 14:32 O.Glamazdin Run 7879: CODA and Scalers
  • 336683 10/25/10 14:11 O.Glamazdin Calorimeter HV
  • 336682 10/25/10 13:45 a-onl Start_Run_7879,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2a condition
  • 336681 10/25/10 13:45 a-onl End_of_Run_7876
  • 336680 10/25/10 13:38 Hisham s1 and s2 efficiency
  • 336679 10/25/10 13:32 O.Glamazdin Run 7878: ROC18
  • 336678 10/25/10 13:32 O.Glamazdin Run 7878: ROC14, ROC17
  • 336677 10/25/10 13:30 O.Glamazdin Run 7878 Coda and Scaler ratesd
  • 336676 10/25/10 13:06 Hisham s2 efficiency with the right figure
  • 336675 10/25/10 13:01 R. Michaels datamon updated again (use T10)
  • 336674 10/25/10 12:51 Hisham s2 efficiency
  • 336673 10/25/10 12:48 a-onl Start_Run_7878,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=24.3,calo_distance=5.5,comment_text=Elastic 2a condition
  • 336672 10/25/10 12:44 jpchen Harp 1H04A
  • 336671 10/25/10 12:43 jpchen Harp 1H04B
  • 336670 10/25/10 12:31 Carlos HV update
  • 336669 10/25/10 12:20 gaskell Unser monitor temp control bad?
  • 336668 10/25/10 12:16 jpchen request beam permit
  • 336667 10/25/10 12:14 A. Camsonne Scope traces block checks
  • 336666 10/25/10 12:13 Magali trigger firmware
  • 336665 10/25/10 11:53 A. Camsonne Scope traces block checks
  • 336664 10/25/10 11:52 jpchen Back to controlled access
  • 336663 10/25/10 11:24 hyde S1 HV adjustment
  • 336662 10/25/10 11:18 Jack Pion Rejector cabling
  • 336661 10/25/10 11:06 jpchen controlled access over, back to beam permit
  • 336660 10/25/10 11:05 jpchen calorimeter and HRSL moved and set for ealstic 2-a
  • 336659 10/25/10 11:03 a-onl End_of_Run_7877
  • 336658 10/25/10 10:53 a-onl Start_Run_7877,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter cosmics with buffering ROC14 and ROC17 - V792 added
  • 336657 10/25/10 10:16 A. Camsonne Left HRS positive polarity and set to 2.245 GeV/c
  • 336656 10/25/10 09:51 A. Camsonne Calorimeter cosmics HV from 600 V to 700 V
  • 336655 10/25/10 09:42 A. Camsonne Run 7876 cosmics run with buffering
  • 336654 10/25/10 09:42 a-onl Start_Run_7876,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter cosmics with buffering ROC14 and ROC17
  • 336653 10/25/10 09:40 O.Glamazdin Moller magnet has been filled
  • 336652 10/25/10 09:38 a-onl End_of_Run_7875
  • 336651 10/25/10 09:35 a-onl Start_Run_7875,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter cosmics
  • 336650 10/25/10 09:13 jpchen In controlled access
  • 336649 10/25/10 08:45 jpchen Controlled access to change to Elastic 2 (a) setting
  • 336648 10/25/10 08:34 R. Michaels datamon updated for DVCS
  • 336647 10/25/10 08:29 jroche turning on and off various pion rejector HV
  • 336646 10/25/10 08:07 adaq Beam Time Accounting
  • 336645 10/25/10 08:06 jroche owl shift summary
  • 336644 10/25/10 07:50 Markowitz moving to empty target, LHRS ramping down
  • 336643 10/25/10 07:41 jroche pion rejector mystery solved?
  • 336642 10/25/10 07:33 lamiaa Rocs[14,17&18] screenshot of a run 7874
  • 336641 10/25/10 07:22 a-onl End_of_Run_7874
  • 336640 10/25/10 07:15 Lamiaa Recapitulation of CODA Screenshot of a run 7873
  • 336639 10/25/10 07:13 Lamiaa Recapitulation of Screenshot of Rocs[14,17&18] of a run 7873
  • 336638 10/25/10 07:04 a-onl Start_Run_7874,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration TrigDVCS -
  • 336637 10/25/10 06:49 Lamiaa CODA screenshot of a run 7873
  • 336636 10/25/10 06:44 Lamiaa Screenshot of Rocs[14,17&18] of a run 7873
  • 336635 10/25/10 06:33 a-onl End_of_Run_7873
  • 336634 10/25/10 06:22 Lamiaa Late halog entry of the end of run #7867
  • 336633 10/25/10 06:07 a-onl Start_Run_7873,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration TrigDVCS -
  • 336632 10/25/10 06:07 jroche all HV back on for the pion rejector
  • 336631 10/25/10 06:06 a-onl End_of_Run_7872
  • 336630 10/25/10 06:06 a-onl End_of_Run_7867
  • 336629 10/25/10 06:02 Lamiaa Screenshot of CODA & Scalers of a run 7872
  • 336628 10/25/10 06:01 Lamiaa Screenshot of Rocs[14,17&18] of a run 7872
  • 336627 10/25/10 05:22 a-onl Start_Run_7872,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration TrigDVCS -
  • 336626 10/25/10 05:21 jroche only prl2 ch 12 and 13 off
  • 336625 10/25/10 05:20 a-onl End_of_Run_7871
  • 336624 10/25/10 05:16 Lamiaa Screenshot of CODA & Scalers of a run 7871
  • 336623 10/25/10 05:15 Lamiaa Screenshot of Rocs[14,17&18] of a run 7871
  • 336622 10/25/10 04:35 a-onl Start_Run_7871,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration TrigDVCS -
  • 336621 10/25/10 04:33 a-onl End_of_Run_7870
  • 336620 10/25/10 04:29 Lamiaa Screenshot of CODA & Scalers of a run 7870
  • 336619 10/25/10 04:29 Lamiaa Screenshot of Rocs[14,17&18] of a run 7870
  • 336618 10/25/10 03:53 Lamiaa Screenshot of CODA & Scalers of a run 7869
  • 336617 10/25/10 03:51 lamiaa Screenshot of Rocs[14&17] of a run 7869
  • 336616 10/25/10 03:47 a-onl Start_Run_7870,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration TrigDVCS -
  • 336615 10/25/10 03:45 a-onl End_of_Run_7869
  • 336614 10/25/10 03:41 jroche more pion rejector off
  • 336613 10/25/10 03:40 a-onl Start_Run_7869,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration DVCSTopBot -
  • 336612 10/25/10 03:39 a-onl End_of_Run_7868
  • 336611 10/25/10 03:36 Lamiaa Screenshot of CODA & Scalers of a run 7868
  • 336610 10/25/10 03:35 Lamiaa Screenshot of Rocs[14&17] of a run 7868
  • 336609 10/25/10 03:20 jroche prl signals during run 7867
  • 336608 10/25/10 03:05 jroche some prl turned off (the screen)
  • 336607 10/25/10 03:04 jroche other Pion rejector turned off.
  • 336606 10/25/10 02:54 a-onl Start_Run_7868,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration DVCSTopBot -
  • 336605 10/25/10 02:50 Lamiaa Screenshot of DOCA & Scalers of a run 7867
  • 336604 10/25/10 02:50 Lamiaa Screenshot of Rocs[14&17] of a run 7867
  • 336603 10/25/10 02:16 jroche pion rejector: some pion rejector off during run 7867
  • 336602 10/25/10 02:09 a-onl Start_Run_7867,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration DVCSTopBot -
  • 336601 10/25/10 02:07 jroche change pedestal for S2R15
  • 336600 10/25/10 02:07 a-onl End_of_Run_7866
  • 336599 10/25/10 02:06 a-onl End_of_Run_7865
  • 336598 10/25/10 02:02 Lamiaa Screenshot of CODA & Scalers of a run 7866
  • 336597 10/25/10 02:01 Lamiaa Screenshot of Rocs[14&17] of a run 7866
  • 336596 10/25/10 01:52 Markowitz target screens
  • 336595 10/25/10 01:35 Lamiaa S0 and S2 distributions of a run 7865
  • 336594 10/25/10 01:22 a-onl Start_Run_7866,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration TricgDVCS -
  • 336593 10/25/10 01:14 Carlos Calorimeter work
  • 336592 10/25/10 00:56 a-onl Start_Run_7865,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration TricgDVCS -
  • 336591 10/25/10 00:50 Markowitz target to empty and back to LH2
  • 336590 10/25/10 00:00 adaq Beam Time Accounting
  • Sunday

  • 336589 10/24/10 23:56 Margaziotis Run 7864 ROC18 screenshot
  • 336588 10/24/10 23:56 Margaziotis Run 7864 ROC14 and ROC17 screenshots
  • 336587 10/24/10 23:54 Margaziotis Run 7854 CODA and Scaler screens
  • 336586 10/24/10 23:54 a-onl End_of_Run_7864
  • 336585 10/24/10 23:54 Ramesh Subed Shift Summary
  • 336584 10/24/10 23:53 Margaziotis Run 7864 HRS spectra need attention
  • 336583 10/24/10 23:53 a-onl Start_Run_7862,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration TricgDVCS -
  • 336582 10/24/10 23:52 Ramesh Asked for short escorted access
  • 336581 10/24/10 23:44 Margaziotis Run 7864 Calorimeter ADC and TDC spectra empty
  • 336580 10/24/10 23:38 Margaziotis Run 7864 Calorimeter ADC 84 spectrum empty
  • 336579 10/24/10 23:22 Carlos T10 fixed
  • 336578 10/24/10 23:13 a-onl Start_Run_7864,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration TricgDVCS -
  • 336577 10/24/10 23:12 a-onl End_of_Run_7863
  • 336576 10/24/10 23:11 a-onl Start_Run_7863,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Configuration TricgDVCS -
  • 336575 10/24/10 23:10 a-onl End_of_Run_7862
  • 336574 10/24/10 23:08 BoZhao Target charts screen
  • 336573 10/24/10 23:03 BoZhao Target to LH2
  • 336572 10/24/10 22:44 BoZhao Target to Empty
  • 336571 10/24/10 22:39 Ramesh Hall in beam permit
  • 336570 10/24/10 22:23 a-onl End_of_Run_7861
  • 336569 10/24/10 22:23 a-onl End_of_Run_7860
  • 336568 10/24/10 22:06 a-onl End_of_Run_7854
  • 336567 10/24/10 21:51 a-onl Start_Run_7860,Run_type=LED,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter LED scan
  • 336566 10/24/10 21:49 a-onl End_of_Run_7859
  • 336565 10/24/10 21:45 a-onl Start_Run_7859,Run_type=LED,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter LED block 1
  • 336564 10/24/10 21:44 a-onl End_of_Run_7858
  • 336563 10/24/10 21:31 a-onl Start_Run_7858,Run_type=LED,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter LED scan
  • 336562 10/24/10 21:30 a-onl End_of_Run_7857
  • 336561 10/24/10 21:27 a-onl Start_Run_7857,Run_type=LED,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter LED block 0,0
  • 336560 10/24/10 21:25 a-onl End_of_Run_7856
  • 336559 10/24/10 21:08 a-onl Start_Run_7856,Run_type=LED,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter LED scan
  • 336558 10/24/10 21:06 a-onl End_of_Run_7855
  • 336557 10/24/10 21:02 a-onl Start_Run_7855,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test
  • 336556 10/24/10 20:50 a-onl Start_Run_7854,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test
  • 336555 10/24/10 20:46 a-onl Start_Run_7853,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test led block 207
  • 336554 10/24/10 20:08 Jixie Moller magnet filling was unusual slow today
  • 336553 10/24/10 20:05 A. Camsonne vxserver loaded on iochacs and dvcstlab6
  • 336552 10/24/10 20:05 jroche, sega pion rejector cabling check
  • 336549 10/24/10 20:00 Hisham BNC synch cable
  • 336548 10/24/10 19:54 J. Zhang Moller magnet has been filled
  • 336547 10/24/10 19:50 a-onl Start_Run_7851,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test led block 207
  • 336546 10/24/10 19:49 a-onl End_of_Run_7850
  • 336545 10/24/10 19:32 a-onl Start_Run_7850,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test led block 207
  • 336544 10/24/10 19:31 a-onl End_of_Run_7849
  • 336543 10/24/10 19:29 Carlos Calo HV ON
  • 336542 10/24/10 19:24 a-onl Start_Run_7849,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test led block 207
  • 336541 10/24/10 19:24 a-onl End_of_Run_7848
  • 336540 10/24/10 19:21 a-onl Start_Run_7848,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test led block 207
  • 336539 10/24/10 19:11 gaskell Run plan for Sunday swing to Monday Day
  • 336538 10/24/10 19:08 a-onl Start_Run_7847,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test led block 207
  • 336537 10/24/10 18:08 Carlos Calorimeter HV off
  • 336536 10/24/10 18:01 jroche more on the pion rejector
  • 336535 10/24/10 17:41 a-onl End_of_Run_7846
  • 336534 10/24/10 17:40 Ramesh Asked for controlled access
  • 336533 10/24/10 17:37 Margaziotis Run 7846 ROC18
  • 336532 10/24/10 17:37 Margaziotis Run 7846 ROC14 and ROC17 screen
  • 336531 10/24/10 17:35 Margaziotis Run 7846 ROC14 and ROC17
  • 336530 10/24/10 17:34 Margaziotis Run 7846 CODA and scalers screenshot
  • 336529 10/24/10 17:26 BoZhao target charts screen
  • 336528 10/24/10 17:13 Carlos 2-photon invariant mass
  • 336527 10/24/10 17:11 jroche one more fix to the end of run epics script (better formatting)
  • 336526 10/24/10 17:11 jroche one more fix to the end of run epics script
  • 336525 10/24/10 17:09 a-onl Start_Run_7846,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout - Changed to TrigDVCS
  • 336524 10/24/10 17:05 a-onl End_of_Run_7845
  • 336523 10/24/10 17:02 Margaziotis Run 7845 ROC14 and ROC17 screenshots
  • 336522 10/24/10 17:02 Margaziotis Run 7845 CODA and scalers screenshot
  • 336521 10/24/10 16:21 a-onl Start_Run_7845,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336520 10/24/10 16:19 a-onl End_of_Run_7844
  • 336519 10/24/10 16:17 Carlos Calo hit map
  • 336518 10/24/10 16:12 Margaziotis run 7844 ROC14 and ROC17
  • 336516 10/24/10 16:11 Margaziotis Run 7844 CODA and scaler screenshots
  • 336517 10/24/10 16:11 adaq Beam Time Accounting
  • 336515 10/24/10 16:05 Lindgren Shift summary
  • 336514 10/24/10 16:03 BoZhao Target charts screen
  • 336513 10/24/10 15:36 Lamiaa LHRS Cerenkov signal on PMT2 and sum back to normal on previous run 7843!
  • 336512 10/24/10 15:36 Carlos Run 7844 with DVCSTopBot configuration
  • 336511 10/24/10 15:30 Lindgren current strip chart
  • 336510 10/24/10 15:29 Lindgren HV DVCS
  • 336509 10/24/10 15:29 a-onl Start_Run_7844,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336508 10/24/10 15:28 Lindgren DVCS DC Levels
  • 336507 10/24/10 15:27 a-onl End_of_Run_7843
  • 336506 10/24/10 15:26 Lindgren Hall A Tools
  • 336505 10/24/10 15:26 Lamiaa Screenshot of CODA & Scalers of a run 7843
  • 336504 10/24/10 15:25 Lamiaa Screenshot of Rocs[14,17&18] of a run 7843
  • 336503 10/24/10 15:11 jroche Modified epics run start and end run to make them faster
  • 336502 10/24/10 15:01 a-onl End_of_Run_7843
  • 336501 10/24/10 15:00 Lamiaa Cerenkov data deviation on PMT2 and sum noticed on runs 7841 & 7842
  • 336500 10/24/10 14:57 gaskell Controlled access at 6:15 pm
  • 336499 10/24/10 14:54 a-onl Start_Run_7843,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336498 10/24/10 14:50 a-onl Start_Run_7843,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336497 10/24/10 14:49 a-onl Start_Run_7843,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336496 10/24/10 14:48 a-onl Start_Run_7843,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336495 10/24/10 14:47 a-onl End_of_Run_7842
  • 336494 10/24/10 14:40 Lamiaa Screenshot of CODA & Scalers of a run 7842
  • 336493 10/24/10 14:37 Lamiaa Screenshot of Rocs[14,17&18] of a run 7842
  • 336492 10/24/10 14:29 M trigger informations
  • 336491 10/24/10 14:24 Magali firmware of the calorimeter trigger
  • 336490 10/24/10 14:12 a-onl Start_Run_7842,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336489 10/24/10 14:11 a-onl End_of_Run_7841
  • 336488 10/24/10 13:51 Lamiaa Screenshot of CODA & Scalers of a run 7841
  • 336487 10/24/10 13:50 Lamiaa Screenshot of Rocs[14,17&18] of a run 7841
  • 336486 10/24/10 13:22 a-onl Start_Run_7841,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336485 10/24/10 13:20 a-onl End_of_Run_7840
  • 336484 10/24/10 13:07 Lamiaa Screenshot of CODA & Scalers of a run 7840
  • 336483 10/24/10 13:06 Lamiaa Screenshot of Rocs[14,17&18] of a run 7840
  • 336482 10/24/10 12:51 Carlos Dead time reading
  • 336481 10/24/10 12:33 Lindgren Lower beam current
  • 336480 10/24/10 12:27 a-onl Start_Run_7840,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336479 10/24/10 12:26 a-onl End_of_Run_7839
  • 336478 10/24/10 12:26 Lindgren Half shift summary
  • 336477 10/24/10 12:20 Lamiaa Screenshot of CODA & Scalers of a run 7839
  • 336476 10/24/10 12:19 Lamiaa Screenshot of Rocs[14,17&18] of a run 7839
  • 336475 10/24/10 12:16 Lamiaa Missing entry of the "End of run 7838"!
  • 336474 10/24/10 12:09 Lamiaa Real Start time is 11:34AM instead of the shown posting time 11:51AM
  • 336473 10/24/10 12:05 Carlos Start/End of run scripts
  • 336472 10/24/10 11:52 a-onl Start_Run_7839,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336470 10/24/10 11:33 Lamiaa Screenshot of CODA and Scalers of a run 7838
  • 336469 10/24/10 11:32 Lamiaa Screenshot of Rocs[14,17&18] of a run 7838
  • 336468 10/24/10 11:27 Lamiaa Halog timing issue!
  • 336467 10/24/10 11:12 Carlos ROC screenshots
  • 336466 10/24/10 11:02 a-onl Start_Run_7838,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336465 10/24/10 11:01 a-onl End_of_Run_7837
  • 336464 10/24/10 10:36 Lamiaa Screenshot of CODA and Scalers of a run 7837
  • 336463 10/24/10 10:35 Lamiaa Screenshot of Rocs[14,17&18] of a run 7837
  • 336462 10/24/10 10:11 gaskell Hall C energy modulation
  • 336461 10/24/10 10:04 a-onl Start_Run_7837,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336460 10/24/10 10:03 a-onl End_of_Run_7836
  • 336459 10/24/10 09:51 lamiaa Screenshot of Rocs[14,17&18] by the end of a run 7836
  • 336458 10/24/10 09:49 lamiaa Screenshot of CODA and Scalers of a run 7836
  • 336457 10/24/10 09:23 Lamiaa Screenshot of Roc17, Roc18 & Roc 14 of run 7836
  • 336456 10/24/10 09:17 a-onl Start_Run_7836,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336455 10/24/10 09:16 a-onl End_of_Run_7835
  • 336454 10/24/10 08:57 a-onl Start_Run_7835,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336453 10/24/10 08:39 Lamiaa Screenshot of CODA and Scaler of a run 7833
  • 336452 10/24/10 08:35 Lamiaa Screenshot of Roc17, Roc18 & Roc14 of a run 7833
  • 336451 10/24/10 08:20 a-onl Start_Run_7833,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336450 10/24/10 08:18 a-onl End_of_Run_7832
  • 336449 10/24/10 08:16 Lindgren Starting Day shift
  • 336448 10/24/10 08:01 efuchey 10/24/10 Owl shift summary
  • 336447 10/24/10 07:59 adaq Beam Time Accounting
  • 336446 10/24/10 07:23 a-onl Start_Run_7832,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336445 10/24/10 07:20 a-onl End_of_Run_7831
  • 336444 10/24/10 07:08 D.Wang rocs
  • 336443 10/24/10 07:07 D.Wang Coda and scalers
  • 336442 10/24/10 06:35 a-onl Start_Run_7831,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336441 10/24/10 05:35 a-onl Start_Run_7830,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336440 10/24/10 05:33 a-onl End_of_Run_7829
  • 336439 10/24/10 05:23 D.Wang rocs
  • 336438 10/24/10 05:22 D.Wang coda and scalers
  • 336437 10/24/10 04:54 efuchey DVCS HV alarm
  • 336436 10/24/10 04:46 a-onl Start_Run_7829,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336435 10/24/10 04:46 a-onl End_of_Run_7825
  • 336434 10/24/10 04:45 a-onl End_of_Run_7828
  • 336433 10/24/10 04:32 D.Wang rocs
  • 336432 10/24/10 04:31 D.Wang Coda and scalers
  • 336431 10/24/10 03:49 a-onl Start_Run_7828,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336430 10/24/10 03:48 a-onl End_of_Run_7827
  • 336429 10/24/10 03:42 D.Wang rocs
  • 336428 10/24/10 03:42 D.Wang coda and scalers
  • 336427 10/24/10 03:15 a-onl Start_Run_7827,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336426 10/24/10 03:13 a-onl End_of_Run_7826
  • 336425 10/24/10 02:56 D.Wang rocs
  • 336424 10/24/10 02:56 D.Wang coda and scalers
  • 336423 10/24/10 02:28 a-onl Start_Run_7826,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336422 10/24/10 02:20 efuchey HV trip: complementary information.
  • 336421 10/24/10 02:18 a-onl Start_Run_7825,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336420 10/24/10 02:16 M. Magne Current logic module version
  • 336419 10/24/10 02:16 a-onl Start_Run_7824,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336418 10/24/10 02:16 a-onl End_of_Run_7824
  • 336417 10/24/10 02:15 a-onl End_of_Run_7823
  • 336416 10/24/10 02:10 a-onl Start_Run_7822,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336415 10/24/10 02:03 a-onl Start_Run_7823,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336414 10/24/10 02:02 a-onl End_of_Run_7821
  • 336413 10/24/10 02:00 a-onl End_of_Run_7822
  • 336412 10/24/10 01:54 efuchey HV trip
  • 336411 10/24/10 01:39 a-onl Start_Run_7820,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336410 10/24/10 01:39 a-onl Start_Run_7821,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Trigger in readout
  • 336409 10/24/10 01:32 a-onl End_of_Run_7820
  • 336408 10/24/10 01:30 A. Camsonne Running config from now on TrigDVCS
  • 336407 10/24/10 01:27 a-onl End_of_Run_7819
  • 336406 10/24/10 01:20 A. Camsonne trigROC18.crl modified to read 127 words TrigDVCS modified - Run 7821
  • 336405 10/24/10 01:07 a-onl Start_Run_7819,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 3 crates - Triiger in readout
  • 336404 10/24/10 01:07 A. Camsonne Added dvcs3 flag to DVCS3 config - starting run 7819
  • 336403 10/24/10 01:01 A. Camsonne Run 7818 Run with ROC18 trigger and Top and Bottom crate
  • 336402 10/24/10 01:00 a-onl Start_Run_7817,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336401 10/24/10 00:59 a-onl End_of_Run_7816
  • 336400 10/24/10 00:52 D.Wang rocs
  • 336399 10/24/10 00:52 D.Wang coda and scalers
  • 336398 10/24/10 00:32 efuchey PbF2 base currents
  • 336397 10/24/10 00:31 efuchey Hall A tools screenshot
  • 336396 10/24/10 00:10 a-onl Start_Run_7816,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336395 10/24/10 00:10 a-onl End_of_Run_7815
  • 336394 10/24/10 00:01 Margaziotis Run 7815 ROC14 and ROC17 screenshots
  • 336392 10/24/10 00:00 Margaziotis Run 7815 DAQ screenshot
  • 336393 10/24/10 00:00 adaq Beam Time Accounting
  • 336391 10/24/10 00:00 GOLGE shift summary for swing
  • Saturday

  • 336390 10/23/10 23:59 Margaziotis Run 7815 scalers screenshot
  • 336389 10/23/10 23:45 Carlos T10 scaler
  • 336388 10/23/10 23:43 Margaziotis HRS spectra need checking
  • 336387 10/23/10 23:21 a-onl Start_Run_7815,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336386 10/23/10 23:16 Maargaziotis Run 7814 ROC14 and ROC17 screenshots
  • 336385 10/23/10 23:15 Margaziotis Run 7814 DAQ screenshot
  • 336384 10/23/10 23:14 Margaziotis Run 7814 scalers screenshot
  • 336383 10/23/10 22:49 a-onl Start_Run_7814,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336382 10/23/10 22:46 a-onl End_of_Run_7813
  • 336381 10/23/10 22:02 a-onl Start_Run_7813,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336380 10/23/10 22:01 a-onl End_of_Run_7812
  • 336379 10/23/10 21:55 Margaziotis Run 7812 ROC14 and ROC17 screenshots
  • 336378 10/23/10 21:54 Margaziotis Run 7812 runcontrol screenshot
  • 336377 10/23/10 21:53 Margaziotis Run 7812 scalers screenshot
  • 336376 10/23/10 21:15 a-onl Start_Run_7812,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336375 10/23/10 21:14 a-onl End_of_Run_7811
  • 336374 10/23/10 21:13 Carlos HV change: block 119
  • 336373 10/23/10 21:03 a-onl Start_Run_7811,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336372 10/23/10 21:01 Margaziotis Restarted Coda to fix bug
  • 336371 10/23/10 20:51 Margaziotis Run 7809 Rock14 and Rock17 screenshots
  • 336370 10/23/10 20:49 Margaziotis Run 7809 runcontrol screenshot
  • 336369 10/23/10 20:47 Margaziotis Run 7809 scalers screenshot
  • 336368 10/23/10 20:37 Margaziotis Calo Amp/TDC 119 spectra run 7809
  • 336367 10/23/10 20:07 a-onl Start_Run_7809,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336366 10/23/10 20:02 ZAFAR AHMED Target is LH2 Loop 2 15 cm
  • 336365 10/23/10 19:56 R. Michaels re: work disk cleanup
  • 336364 10/23/10 19:37 a-onl End_of_Run_7808
  • 336363 10/23/10 19:31 a-onl Start_Run_7808,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336362 10/23/10 19:29 Carlos Block 119 bloken, fixed now
  • 336361 10/23/10 19:23 A. Camsonne HV cosmics off - reloaded HV from Oct 19
  • 336360 10/23/10 19:06 a-onl End_of_Run_7807
  • 336359 10/23/10 19:04 a-onl Start_Run_7807,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - test DVCSTopBot - Still working !
  • 336358 10/23/10 19:02 a-onl Start_Run_7806,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - test 3 crates
  • 336357 10/23/10 19:01 a-onl End_of_Run_7802
  • 336356 10/23/10 19:01 a-onl Start_Run_7805,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 2 crates with buffering - sync events set to 1000
  • 336355 10/23/10 18:59 a-onl Start_Run_7804,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 2 crates with buffering - sync events set to 10
  • 336354 10/23/10 18:59 a-onl End_of_Run_7804
  • 336353 10/23/10 18:55 a-onl End_of_Run_7803
  • 336352 10/23/10 18:47 A. Camsonne Changed sync event frequency to 10
  • 336351 10/23/10 18:24 a-onl Start_Run_7803,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 2 crates with buffering
  • 336350 10/23/10 18:24 A. Camsonne Nothing on BNC sync signal
  • 336349 10/23/10 18:23 a-onl Start_Run_7800,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 2 crates with buffering
  • 336348 10/23/10 18:23 A. Camsonne Buffering making empty FIFO at 60 Hz of cosmics
  • 336347 10/23/10 18:16 A. Camsonne crl for ROC17 and ROC14 modified
  • 336346 10/23/10 18:11 a-onl Start_Run_7802,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 2 crates with buffering
  • 336345 10/23/10 18:10 a-onl End_of_Run_7801
  • 336344 10/23/10 18:04 a-onl Start_Run_7801,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 2 crates with buffering
  • 336343 10/23/10 18:04 a-onl End_of_Run_7800
  • 336342 10/23/10 18:02 a-onl End_of_Run_7799
  • 336341 10/23/10 17:40 a-onl Start_Run_7799,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - 2 crates with buffering
  • 336340 10/23/10 17:09 a-onl Start_Run_7798,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp -test 3 crates
  • 336339 10/23/10 17:07 a-onl Start_Run_7797,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp -test
  • 336338 10/23/10 17:06 a-onl End_of_Run_7796
  • 336337 10/23/10 17:04 a-onl Start_Run_7796,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp -test
  • 336336 10/23/10 17:03 a-onl End_of_Run_7795
  • 336335 10/23/10 16:56 a-onl Start_Run_7795,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp -test
  • 336334 10/23/10 16:54 A. Camsonne ROC18 hardware ROC ID switched from 5 to 1
  • 336333 10/23/10 16:51 a-onl End_of_Run_7794
  • 336332 10/23/10 16:32 a-onl Start_Run_7794,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp -test
  • 336331 10/23/10 16:19 J. Zhang Moller magnet has been filled
  • 336330 10/23/10 16:04 Lindgren Shift summary
  • 336329 10/23/10 16:00 adaq Beam Time Accounting
  • 336328 10/23/10 15:45 a-onl End_of_Run_7790
  • 336327 10/23/10 15:39 a-onl Start_Run_7788,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - test buffering -with validation
  • 336326 10/23/10 15:38 a-onl End_of_Run_7789
  • 336325 10/23/10 15:38 A. Camsonne Test with validation - need to check timing
  • 336324 10/23/10 15:38 a-onl Start_Run_7789,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - test buffering -with validation
  • 336323 10/23/10 15:30 a-onl End_of_Run_7788
  • 336322 10/23/10 15:29 a-onl End_of_Run_7787
  • 336321 10/23/10 15:26 a-onl Start_Run_7787,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp - test buffering
  • 336320 10/23/10 15:21 A. Camsonne Run 7787 test with buffering
  • 336319 10/23/10 14:58 A. Camsonne ROC1417Buff modified for buffering
  • 336318 10/23/10 14:56 a-onl End_of_Run_7786
  • 336317 10/23/10 14:56 a-onl End_of_Run_7785
  • 336316 10/23/10 14:52 chench scalers and DAQ of run7786
  • 336315 10/23/10 14:51 chench Screenshot of ROC14 and ROC17 for run7786
  • 336314 10/23/10 14:24 a-onl Start_Run_7786,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336313 10/23/10 14:23 a-onl End_of_Run_7781
  • 336312 10/23/10 14:03 chench Screenshot of ROC14 and ROC17 for run7785
  • 336311 10/23/10 14:02 chench Scalers of run7785
  • 336310 10/23/10 14:02 chench Screenshot of DAQ for run#7785
  • 336309 10/23/10 14:00 Lindgren DVCS Calo
  • 336308 10/23/10 13:59 lindgren Lecroy HV
  • 336307 10/23/10 13:57 Lindgren beam chart
  • 336306 10/23/10 13:54 Lindgren Hall A Tools
  • 336305 10/23/10 13:53 Lindgren DVCS ADC Readback
  • 336304 10/23/10 13:52 Ole Hansen Several work1 disks are full/nearly full
  • 336303 10/23/10 13:40 chench Missing channels of pioin rejectors1 ADC
  • 336302 10/23/10 13:36 markowit target screens
  • 336301 10/23/10 13:36 Ole Hansen Don't fill adaq home directory quota!
  • 336300 10/23/10 13:31 a-onl Start_Run_7785,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336299 10/23/10 13:29 a-onl End_of_Run_7784
  • 336298 10/23/10 13:10 Margaziotis Screenshot DAQ Run 7784
  • 336297 10/23/10 13:09 Margaziotis ROCK14 &ROCK17 run 7784
  • 336296 10/23/10 13:08 Margaziotis Screenshot for scalers run 7784
  • 336295 10/23/10 12:24 a-onl Start_Run_7784,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336294 10/23/10 12:24 a-onl End_of_Run_7783
  • 336293 10/23/10 12:13 chench Screenshot for DAQ of run7783
  • 336292 10/23/10 12:12 chench Screenshot for ROC14 and ROC17 of run7783
  • 336291 10/23/10 12:11 chench Scalers for run7783
  • 336290 10/23/10 11:37 a-onl Start_Run_7783,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336289 10/23/10 11:35 a-onl End_of_Run_7782
  • 336288 10/23/10 11:18 chench screen shot of ROC14 and ROC17 for run7782
  • 336287 10/23/10 11:17 chench Screen shot of DAQ for run7782
  • 336286 10/23/10 11:16 chench Scalers of s1, s2 for run7782
  • 336285 10/23/10 11:06 A. Camsonne Can you also take the scalers norm tab ?
  • 336284 10/23/10 10:43 a-onl Start_Run_7782,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336283 10/23/10 10:36 gaskell Access scheduled for 3 pm
  • 336282 10/23/10 10:25 chench Screenshot of DAQ for run7781
  • 336281 10/23/10 10:24 chench Scalers for run7781
  • 336280 10/23/10 10:23 chench Roc14 and Roc17 screenshot for run7781
  • 336279 10/23/10 10:01 A. Camsonne Issues with power cycling the crates
  • 336278 10/23/10 10:01 A. Camsonne Issues with power cycling the crates
  • 336277 10/23/10 09:43 Lindgren Coda failure
  • 336276 10/23/10 09:41 a-onl Start_Run_7781,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336275 10/23/10 09:12 chench Error from DAQ during run7780
  • 336274 10/23/10 08:54 a-onl Start_Run_7780,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336273 10/23/10 08:52 a-onl End_of_Run_7779
  • 336272 10/23/10 08:50 chench Roc14 and Roc17 screen shot at the end of run7779
  • 336271 10/23/10 08:14 LIndgren Statring Day Shift
  • 336270 10/23/10 08:11 a-onl Start_Run_7779,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336269 10/23/10 08:10 a-onl End_of_Run_7778
  • 336268 10/23/10 08:00 adaq Beam Time Accounting
  • 336267 10/23/10 07:59 efuchey Owl 10-23-10 shift summary
  • 336266 10/23/10 07:49 D.Wang rocs
  • 336265 10/23/10 07:49 D.Wang coda and scalers
  • 336264 10/23/10 07:09 a-onl Start_Run_7778,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336263 10/23/10 07:08 a-onl End_of_Run_7777
  • 336262 10/23/10 06:56 D.Wang rocs
  • 336261 10/23/10 06:56 D.Wang coda and scalers
  • 336260 10/23/10 06:12 a-onl Start_Run_7777,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336259 10/23/10 06:12 a-onl End_of_Run_7776
  • 336258 10/23/10 06:06 D.Wang rocs
  • 336257 10/23/10 06:05 D.Wang coda and scalers
  • 336256 10/23/10 05:44 efuchey PbF2 base currents
  • 336255 10/23/10 05:24 a-onl Start_Run_7776,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336254 10/23/10 05:21 efuchey Target screens
  • 336253 10/23/10 05:21 a-onl End_of_Run_7775
  • 336252 10/23/10 05:20 efuchey Beam back
  • 336251 10/23/10 05:10 D.Wang rocs
  • 336250 10/23/10 05:10 D.Wang Coda and scalers
  • 336249 10/23/10 04:38 a-onl Start_Run_7775,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336248 10/23/10 04:37 a-onl End_of_Run_7774
  • 336247 10/23/10 03:43 efuchey Beqam down again
  • 336246 10/23/10 03:26 efuchey Hall C feedback test done
  • 336245 10/23/10 03:14 a-onl Start_Run_7774,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336244 10/23/10 03:13 a-onl End_of_Run_7773
  • 336243 10/23/10 03:09 efuchey Beam is back
  • 336242 10/23/10 02:50 a-onl Start_Run_7773,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336241 10/23/10 02:50 a-onl End_of_Run_7771
  • 336240 10/23/10 02:49 a-onl End_of_Run_7772
  • 336239 10/23/10 02:47 efuchey Beam down for more than 15 minutes
  • 336238 10/23/10 02:23 D.Wang rocs
  • 336237 10/23/10 02:23 D.Wang coda and scalers
  • 336236 10/23/10 01:54 efuchey Hall A tools screenshot
  • 336235 10/23/10 01:53 a-onl Start_Run_7772,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336234 10/23/10 01:52 a-onl End_of_Run_7763
  • 336233 10/23/10 01:32 D.Wang rocs
  • 336232 10/23/10 01:32 D.Wang coda and scaler
  • 336231 10/23/10 01:16 efuchey Compton rates for tonight
  • 336230 10/23/10 00:53 a-onl Start_Run_7771,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336229 10/23/10 00:52 a-onl End_of_Run_7770
  • 336228 10/23/10 00:46 D.Wang rocs
  • 336227 10/23/10 00:45 D.Wang runcontrol and scalers
  • 336225 10/23/10 00:03 ZAFAR AHMED Friday 10-22-2010 Swing Shift Summary
  • 336226 10/23/10 00:02 adaq Beam Time Accounting
  • Friday

  • 336224 10/22/10 23:44 a-onl Start_Run_7770,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336223 10/22/10 23:44 a-onl End_of_Run_7769
  • 336222 10/22/10 23:42 Nuruzzaman Target GUI and strip chart screenshot.
  • 336221 10/22/10 23:38 Punjabi ROC screens
  • 336220 10/22/10 23:38 Punjabi RC and scaler screens
  • 336219 10/22/10 23:21 ZAFAR AHMED Hall C is testing feedback
  • 336218 10/22/10 23:03 a-onl Start_Run_7769,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336217 10/22/10 22:55 a-onl End_of_Run_7768
  • 336216 10/22/10 22:50 Punjabi ROC screens
  • 336215 10/22/10 22:49 Punjabi RC and scaler screens
  • 336214 10/22/10 22:08 a-onl Start_Run_7768,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336213 10/22/10 22:07 a-onl End_of_Run_7767
  • 336212 10/22/10 22:01 Punjabi roc sreens
  • 336211 10/22/10 22:00 Punjabi RC and scalers
  • 336210 10/22/10 21:41 ZAFAR AHMED PbF Calorimeter ADC readback
  • 336209 10/22/10 21:19 a-onl Start_Run_7767,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 3 4 pass beam, 2 microamp
  • 336208 10/22/10 21:17 a-onl End_of_Run_7766
  • 336207 10/22/10 21:16 a-onl Start_Run_7766,Run_type=Pedestal,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Pedastal run
  • 336206 10/22/10 21:10 a-onl End_of_Run_7765
  • 336205 10/22/10 21:08 Punjabi run control, scalers screens
  • 336204 10/22/10 21:06 Punjabi Roc 17 and Roc 14 screens
  • 336203 10/22/10 20:40 gaskell Status and Runplan for remainder of swing and owl shift
  • 336202 10/22/10 20:29 a-onl Start_Run_7765,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin3 - 4pass PRODUCTION (2uA)
  • 336201 10/22/10 20:29 Carlos jmysql is back
  • 336200 10/22/10 20:27 a-onl End_of_Run_7764
  • 336199 10/22/10 20:11 a-onl Start_Run_7764,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin3 - 4pass PRODUCTION (1uA)
  • 336198 10/22/10 20:02 Nuruzzaman Target moved to LH2, LOOP 2 15 cm.
  • 336197 10/22/10 19:56 Carlos Pedestal update
  • 336196 10/22/10 19:55 ZAFAR AHMED HARP scan
  • 336195 10/22/10 19:28 Carlos jmysql server down
  • 336194 10/22/10 19:25 efuchey weird cavity behavior
  • 336193 10/22/10 18:02 hyde Deadtime study vs rate
  • 336192 10/22/10 17:52 A. Camsonne DVCS DAQ test ROC18
  • 336191 10/22/10 17:43 a-onl Start_Run_7763,Run_type=Other,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Check ROC14 ROC17
  • 336190 10/22/10 17:15 a-onl End_of_Run_7761
  • 336189 10/22/10 17:15 a-onl Start_Run_7761,Run_type=Pedestal,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter pedestal run (lower pulse rate)
  • 336188 10/22/10 17:14 a-onl End_of_Run_7760
  • 336187 10/22/10 17:14 a-onl End_of_Run_7759
  • 336186 10/22/10 17:12 a-onl Start_Run_7759,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test ROC14 17 test crate
  • 336185 10/22/10 17:07 gaskell Beam permit
  • 336184 10/22/10 17:04 gaskell New kinematics: Kin 3, low energy
  • 336183 10/22/10 16:55 a-onl End_of_Run_7757
  • 336182 10/22/10 16:46 a-onl Start_Run_7757,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test ROC14 17 test crate
  • 336181 10/22/10 16:46 a-onl Start_Run_7753,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test ROC14 17 test crate
  • 336180 10/22/10 16:45 a-onl Start_Run_7756,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test ROC14 17 test crate
  • 336179 10/22/10 16:44 a-onl End_of_Run_7755
  • 336178 10/22/10 16:42 a-onl End_of_Run_7745
  • 336177 10/22/10 16:42 a-onl Start_Run_7755,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test ROC14 17 test crate
  • 336176 10/22/10 16:41 a-onl End_of_Run_7754
  • 336175 10/22/10 16:39 a-onl Start_Run_7754,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test ROC14 17 test crate
  • 336174 10/22/10 16:37 a-onl End_of_Run_7753
  • 336173 10/22/10 16:37 jroche pion rejector signal when HV labelled PRL1 are off
  • 336172 10/22/10 16:33 a-onl End_of_Run_7751
  • 336171 10/22/10 16:32 a-onl Start_Run_7751,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test ROC18 test crate
  • 336170 10/22/10 16:31 a-onl End_of_Run_7750
  • 336169 10/22/10 16:25 hyde clock scalers in replay
  • 336168 10/22/10 16:16 jroche, cshl Pion rejector: database detector map match reality
  • 336167 10/22/10 16:06 A. Camsonne VME ADC
  • 336166 10/22/10 16:05 a-onl End_of_Run_7749
  • 336165 10/22/10 16:04 Saha Shift Summary
  • 336164 10/22/10 16:03 adaq Beam Time Accounting
  • 336163 10/22/10 15:57 a-onl Start_Run_7749,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test ROC18 test crate
  • 336162 10/22/10 15:56 a-onl End_of_Run_7748
  • 336161 10/22/10 15:55 a-onl Start_Run_7748,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test ROC18 test crate
  • 336160 10/22/10 15:54 a-onl End_of_Run_7747
  • 336159 10/22/10 15:54 a-onl Start_Run_7747,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text= test ROC18 test crate
  • 336158 10/22/10 15:23 a-onl End_of_Run_7743
  • 336157 10/22/10 15:22 a-onl Start_Run_7743,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test prl
  • 336156 10/22/10 15:21 a-onl End_of_Run_7742
  • 336155 10/22/10 15:17 a-onl Start_Run_7742,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test prl
  • 336154 10/22/10 15:16 a-onl End_of_Run_7741
  • 336153 10/22/10 15:00 a-onl Start_Run_7741,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test prl
  • 336152 10/22/10 14:59 a-onl End_of_Run_7740
  • 336151 10/22/10 14:48 a-onl Start_Run_7740,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Cosmics - test buffering
  • 336150 10/22/10 14:20 ellie ADCs look worse after pedestal update (run 7737)
  • 336149 10/22/10 14:16 J. Zhang Moller magnet is full
  • 336148 10/22/10 13:55 Carlos Pedestals update
  • 336147 10/22/10 10:58 a-onl Start_Run_7739,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Cosmics - test buffering
  • 336146 10/22/10 10:37 a-onl Start_Run_7738,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Cosmics - test buffering
  • 336145 10/22/10 10:34 ellie (Run 7737) CaloAmp17 ADC looks off and has noisy TDCs
  • 336144 10/22/10 10:29 Lindgren Target charts
  • 336143 10/22/10 10:17 lindgren target photo
  • 336142 10/22/10 08:38 gaskell Status update and plan for the day
  • 336141 10/22/10 08:36 a-onl End_of_Run_7737
  • 336140 10/22/10 08:27 a-onl Start_Run_7737,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, 1 uA, DVCSTopBot, ps5=1
  • 336139 10/22/10 08:26 a-onl End_of_Run_7736
  • 336138 10/22/10 08:26 a-onl End_of_Run_7732
  • 336137 10/22/10 08:00 horn Owl shift summary
  • 336136 10/22/10 08:00 adaq Beam Time Accounting
  • 336135 10/22/10 07:48 a-onl Start_Run_7736,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, 1 muA, DVCSTopBot, ps5=1
  • 336134 10/22/10 07:45 a-onl End_of_Run_7735
  • 336133 10/22/10 06:58 a-onl Start_Run_7735,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, 1 muA, DVCSTopBot, ps5=1
  • 336132 10/22/10 06:41 horn Beam back
  • 336131 10/22/10 06:39 a-onl End_of_Run_7734
  • 336130 10/22/10 06:31 horn Beam down for beam studies preparation
  • 336129 10/22/10 06:27 a-onl Start_Run_7734,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, 1 muA, DVCSTopBot, ps5=1
  • 336128 10/22/10 06:25 a-onl End_of_Run_7733
  • 336127 10/22/10 05:38 a-onl Start_Run_7733,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, 1 muA, DVCSTopBot, ps5=1
  • 336126 10/22/10 05:07 ent calo TDCs back
  • 336125 10/22/10 04:58 Carlos online gui
  • 336124 10/22/10 04:52 a-onl Start_Run_7732,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, 1 muA, DVCSTopBot, ps5=1
  • 336123 10/22/10 04:49 horn Rebooted calo HV gui
  • 336122 10/22/10 04:35 ent many calo TDC channels missing
  • 336121 10/22/10 04:35 a-onl Start_Run_7731,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, 1 muA, DVCSTopBot, ps5=1
  • 336120 10/22/10 04:23 a-onl End_of_Run_7730
  • 336119 10/22/10 04:18 ent problems analyzing full file
  • 336118 10/22/10 04:12 ent dead time study - runs
  • 336117 10/22/10 03:57 a-onl Start_Run_7730,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=0.5,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, dead time study, DVCSTopBot, ps5=1
  • 336116 10/22/10 03:56 a-onl End_of_Run_7729
  • 336115 10/22/10 03:41 a-onl Start_Run_7729,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, dead time study, DVCSTopBot, ps5=1
  • 336114 10/22/10 03:38 a-onl End_of_Run_7728
  • 336113 10/22/10 03:23 a-onl Start_Run_7728,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.2,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, dead time study, DVCSTopBot, ps5=1
  • 336112 10/22/10 03:22 a-onl End_of_Run_7727
  • 336111 10/22/10 03:05 a-onl Start_Run_7727,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3.5,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, dead time study, DVCSTopBot, ps5=1
  • 336110 10/22/10 02:54 a-onl End_of_Run_7726
  • 336109 10/22/10 02:41 a-onl Start_Run_7726,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, dead time study, DVCSTopBot, ps5=1
  • 336108 10/22/10 02:40 a-onl End_of_Run_7725
  • 336107 10/22/10 02:24 a-onl Start_Run_7725,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3.5,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, dead time study, DVCSTopBot, ps5=1
  • 336106 10/22/10 02:15 a-onl End_of_Run_7724
  • 336105 10/22/10 01:59 a-onl Start_Run_7724,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.6,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, dead time study, DVCSTopBot, ps5=1
  • 336104 10/22/10 01:53 a-onl End_of_Run_7723
  • 336103 10/22/10 01:38 a-onl Start_Run_7723,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.1,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, dead time study, DVCSTopBot, ps5=1
  • 336102 10/22/10 01:36 a-onl End_of_Run_7722
  • 336101 10/22/10 01:24 a-onl Start_Run_7722,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.6,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, dead time study, DVCSTopBot, ps5=1
  • 336100 10/22/10 01:20 a-onl End_of_Run_7721
  • 336099 10/22/10 01:20 A. Camsonne DAQ status
  • 336098 10/22/10 01:06 A. Camsonne Run 7720 with calorimeter read out
  • 336097 10/22/10 01:05 Markowitz target status
  • 336096 10/22/10 01:01 a-onl Start_Run_7721,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS production run, dead time study, DVCSTopBot, ps5=1
  • 336095 10/22/10 00:58 a-onl End_of_Run_7720
  • 336094 10/22/10 00:57 horn Starting dead time study
  • 336093 10/22/10 00:50 a-onl Start_Run_7720,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.591,leftarm_theta=21.489,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS test run, calo on, DVCSTopBot, ps5=1
  • 336092 10/22/10 00:48 horn Calorimeter HV on for test run
  • 336091 10/22/10 00:47 horn Change to nominal configuration
  • 336090 10/22/10 00:41 Markowitz more cernox L2 oscillations
  • 336089 10/22/10 00:13 a-onl Start_Run_7718,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.539,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps3=ps4=1
  • 336088 10/22/10 00:03 a-onl End_of_Run_7717
  • 336086 10/22/10 00:03 Markowitz move target to LH2. 15 cm
  • 336085 10/22/10 00:02 Solvignon Swing shift summary
  • 336087 10/22/10 00:02 adaq Beam Time Accounting
  • 336084 10/22/10 00:02 a-onl Start_Run_7717,Run_type=Production,target_type=15cm dummy,beam_e0=5.5,beam_raster=on,beam_current=4.5,leftarm_p=2.539,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps3=ps4=1
  • Thursday

  • 336083 10/21/10 23:52 a-onl End_of_Run_7716
  • 336082 10/21/10 23:50 Solvignon Spot++
  • 336081 10/21/10 23:43 a-onl Start_Run_7716,Run_type=Production,target_type=15cm dummy,beam_e0=5.5,beam_raster=on,beam_current=1,leftarm_p=2.539,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps3=ps4=1
  • 336080 10/21/10 23:36 cshloka target moved to Dummy, 15cm
  • 336079 10/21/10 23:30 a-onl End_of_Run_7715
  • 336078 10/21/10 23:08 cshloka fluctuation in cernox 2 R
  • 336077 10/21/10 23:08 a-onl Start_Run_7715,Run_type=Production,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=off,beam_current=1,leftarm_p=2.539,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps3=ps4=1
  • 336076 10/21/10 23:05 a-onl End_of_Run_7714
  • 336075 10/21/10 22:55 Solvignon LHRS at 2.539GeV/c
  • 336074 10/21/10 22:41 a-onl Start_Run_7714,Run_type=Production,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=off,beam_current=1,leftarm_p=2.591,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps5=1
  • 336073 10/21/10 22:41 A. Camsonne Pion rejector focal plane y amplitude distribution
  • 336072 10/21/10 22:36 A. Camsonne TDCs look ok now in T5
  • 336071 10/21/10 22:34 a-onl End_of_Run_7713
  • 336070 10/21/10 22:28 a-onl Start_Run_7713,Run_type=Production,target_type=15cm dummy,beam_e0=5.5,beam_raster=on,beam_current=1,leftarm_p=2.591,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps5=1
  • 336069 10/21/10 22:24 cshloka target moved to carbon
  • 336068 10/21/10 22:13 a-onl End_of_Run_7710
  • 336067 10/21/10 22:13 a-onl Start_Run_7712,Run_type=Production,target_type=15cm dummy,beam_e0=5.5,beam_raster=off,beam_current=1,leftarm_p=2.591,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps3=ps4=1
  • 336066 10/21/10 22:09 Solvignon Spot++
  • 336065 10/21/10 21:56 cshloka target moved to Dummy, 15cm
  • 336064 10/21/10 21:55 a-onl End_of_Run_7711
  • 336063 10/21/10 21:38 a-onl Start_Run_7711,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1,leftarm_p=2.591,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps3=ps4=1
  • 336062 10/21/10 21:30 Solvignon LHRS at 2.591GeV/c
  • 336061 10/21/10 21:16 a-onl Start_Run_7710,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1,leftarm_p=2.645,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps3=ps4=1
  • 336060 10/21/10 21:10 cshloka target moved to Loop 2, 15cm
  • 336059 10/21/10 21:09 a-onl End_of_Run_7709
  • 336058 10/21/10 20:58 Solvignon spot++: now raster pattern looks fine at 2x2
  • 336057 10/21/10 20:56 a-onl Start_Run_7709,Run_type=Production,target_type=15cm dummy,beam_e0=5.5,beam_raster=on,beam_current=4.5,leftarm_p=2.645,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps3=ps4=1
  • 336056 10/21/10 20:52 Nuruzzaman Coda crashed during begining of the run 7707.
  • 336055 10/21/10 20:49 a-onl End_of_Run_7707
  • 336054 10/21/10 20:45 cshloka fluctuation in cernox_2_R still exists,
  • 336053 10/21/10 20:41 a-onl Start_Run_7707,Run_type=Production,target_type=15cm dummy,beam_e0=5.5,beam_raster=on,beam_current=4.5,leftarm_p=2.645,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps3=ps4=1
  • 336052 10/21/10 20:36 a-onl End_of_Run_7706
  • 336051 10/21/10 20:36 Solvignon Weird raster pattern
  • 336050 10/21/10 20:34 a-onl Start_Run_7706,Run_type=Production,target_type=15cm dummy,beam_e0=5.5,beam_raster=on,beam_current=4.5,leftarm_p=2.645,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps3=ps4=1
  • 336049 10/21/10 20:31 a-onl End_of_Run_7705
  • 336048 10/21/10 20:26 cshloka fluctuation in cernox_2_R temperature
  • 336047 10/21/10 20:24 cshloka Target moved to Dummy, 15cm
  • 336046 10/21/10 20:18 Solvignon Cerenkov 5,6 and 7 a bit weak
  • 336045 10/21/10 20:05 a-onl Start_Run_7705,Run_type=Production,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps3=ps4=1
  • 336044 10/21/10 19:51 cshloka target moved to Carbon
  • 336043 10/21/10 19:50 Solvignon Start LHRS delta scan: P0=2.643 GeV/c
  • 336042 10/21/10 19:49 a-onl Start_Run_7704,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production DIS delta Scan, Config LHRs, ps3=ps4=1
  • 336041 10/21/10 19:44 gaskell Setting up for DIS delta scan
  • 336040 10/21/10 19:39 A. Camsonne DVCS Logic module
  • 336039 10/21/10 19:35 a-onl End_of_Run_7703
  • 336038 10/21/10 19:26 Carlos Calorimeter HV off
  • 336037 10/21/10 19:25 A. Camsonne Timing spectra for T3 only run
  • 336036 10/21/10 19:17 a-onl Start_Run_7703,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production
  • 336035 10/21/10 19:13 cshloka Fluctuation in cernox_2_R value
  • 336034 10/21/10 19:10 a-onl End_of_Run_7699
  • 336033 10/21/10 19:10 a-onl End_of_Run_7702
  • 336032 10/21/10 19:02 A. Camsonne ARS firmware version E10 downloaded
  • 336031 10/21/10 19:01 cshloka target status
  • 336030 10/21/10 18:57 a-onl Start_Run_7702,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - LeftHRS, ps3=1
  • 336029 10/21/10 18:54 a-onl End_of_Run_7701
  • 336028 10/21/10 18:21 a-onl Start_Run_7701,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom
  • 336027 10/21/10 18:20 a-onl End_of_Run_7700
  • 336026 10/21/10 17:55 efuchey Compton rates recheck
  • 336025 10/21/10 17:54 a-onl Start_Run_7700,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom
  • 336024 10/21/10 17:52 a-onl End_of_Run_7696
  • 336023 10/21/10 17:51 a-onl Start_Run_7699,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom
  • 336022 10/21/10 17:51 a-onl End_of_Run_7698
  • 336021 10/21/10 17:33 a-onl Start_Run_7698,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom
  • 336020 10/21/10 17:32 a-onl End_of_Run_7697
  • 336019 10/21/10 17:29 A. Camsonne Run 7694 Cerenkov
  • 336018 10/21/10 17:29 hyde DVCS Trig scalers
  • 336017 10/21/10 17:28 A. Camsonne Timing plots 7694 s2m TDCs
  • 336016 10/21/10 17:27 A. Camsonne Weird timing distribution in data - Run 7694
  • 336015 10/21/10 17:20 a-onl Start_Run_7697,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom
  • 336014 10/21/10 17:18 a-onl Start_Run_7689,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom - ARS firmware version E09
  • 336013 10/21/10 17:12 hyde DVCS Trig scalers
  • 336012 10/21/10 17:03 a-onl Start_Run_7696,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom
  • 336011 10/21/10 17:01 a-onl End_of_Run_7695
  • 336010 10/21/10 17:00 hyde DVCS Trig scalers
  • 336009 10/21/10 16:51 gaskell Pockels cell OFF last night
  • 336008 10/21/10 16:26 a-onl Start_Run_7695,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom - ARS firmware version E09 - Soft reset - No validation
  • 336007 10/21/10 16:23 a-onl End_of_Run_7694
  • 336006 10/21/10 16:19 cshloka target snapshot at the start of shift
  • 336005 10/21/10 16:13 Solvignon Gas shed airflow alarm
  • 336004 10/21/10 16:02 adaq Beam Time Accounting
  • 336003 10/21/10 16:01 efuchey Day shift summary 10/21/2010
  • 336002 10/21/10 15:41 a-onl Start_Run_7694,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom - ARS firmware version E09 - Soft reset - No validation
  • 336001 10/21/10 15:40 a-onl End_of_Run_7693
  • 336000 10/21/10 15:36 A. Camsonne Run 7694 no validation run
  • 335999 10/21/10 15:33 a-onl Start_Run_7693,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom - ARS firmware version E09 - Soft reset
  • 335998 10/21/10 15:32 efuchey master IOC reboot from MCC
  • 335997 10/21/10 15:31 a-onl End_of_Run_7692
  • 335996 10/21/10 15:28 a-onl Start_Run_7692,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom - ARS firmware version E09 - Soft reset
  • 335995 10/21/10 15:26 a-onl Start_Run_7691,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom - ARS firmware version E09 - Soft reset
  • 335994 10/21/10 15:26 a-onl End_of_Run_7691
  • 335993 10/21/10 15:25 a-onl End_of_Run_7690
  • 335992 10/21/10 15:21 a-onl Start_Run_7690,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom - ARS firmware version E09 - Soft reset
  • 335991 10/21/10 15:21 A. Camsonne Potential good run with E09 Run 7692
  • 335990 10/21/10 15:16 A. Camsonne Readout file for ROC14 and ROC17 modified to use soft reset
  • 335989 10/21/10 15:05 A. Camsonne Oops it was run 7689
  • 335988 10/21/10 15:05 A. Camsonne Time out in read out increased Run 7688
  • 335987 10/21/10 14:53 a-onl End_of_Run_7688
  • 335986 10/21/10 14:51 a-onl Start_Run_7688,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom
  • 335985 10/21/10 14:51 A. Camsonne Loading ARS firmware E09 for testing
  • 335984 10/21/10 14:50 a-onl End_of_Run_7687
  • 335983 10/21/10 14:08 Hisham VME4: warning
  • 335982 10/21/10 14:05 a-onl Start_Run_7687,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom
  • 335981 10/21/10 13:58 Hisham Run# 7686
  • 335980 10/21/10 13:39 efuchey Compton rates for today
  • 335979 10/21/10 13:34 Hisham vme4--hatsv4-port-7 warning
  • 335978 10/21/10 13:27 a-onl Start_Run_7686,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - top bottom
  • 335977 10/21/10 13:18 Hisham Target moved to LH2, 15cm
  • 335976 10/21/10 13:12 efuchey Beam recovery procedure
  • 335975 10/21/10 12:55 Hisham Target moved to empty
  • 335974 10/21/10 12:51 hyde dvcs scalar problems
  • 335973 10/21/10 12:38 efuchey Back to beam permit
  • 335972 10/21/10 12:37 A. Camsonne DVCS DAQ work
  • 335971 10/21/10 12:33 a-onl Start_Run_7685,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - top bottom
  • 335970 10/21/10 12:32 a-onl Start_Run_7684,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - 3 rocs
  • 335969 10/21/10 12:32 a-onl End_of_Run_7684
  • 335968 10/21/10 12:31 a-onl End_of_Run_7683
  • 335967 10/21/10 12:30 a-onl Start_Run_7683,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - roc 14
  • 335966 10/21/10 12:26 a-onl End_of_Run_7682
  • 335965 10/21/10 12:20 a-onl Start_Run_7682,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - roc 14 17
  • 335964 10/21/10 12:19 a-onl End_of_Run_7681
  • 335963 10/21/10 12:17 a-onl Start_Run_7681,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - roc 14 17
  • 335962 10/21/10 12:15 a-onl End_of_Run_7680
  • 335961 10/21/10 12:14 a-onl Start_Run_7680,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - roc 18
  • 335960 10/21/10 12:11 Magali Lemo INPUT and OUTPUT of the claorimeter trigger
  • 335959 10/21/10 12:08 a-onl End_of_Run_7679
  • 335958 10/21/10 12:08 a-onl Start_Run_7671,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - top crate with valid
  • 335957 10/21/10 12:08 a-onl Start_Run_7679,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - roc 18
  • 335956 10/21/10 12:08 hyde Trigger modifications
  • 335955 10/21/10 12:02 a-onl Start_Run_7678,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - roc 17 and 14
  • 335954 10/21/10 12:01 a-onl End_of_Run_7677
  • 335953 10/21/10 12:01 a-onl Start_Run_7677,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - 3 crates
  • 335952 10/21/10 12:00 a-onl End_of_Run_7676
  • 335951 10/21/10 11:56 a-onl Start_Run_7676,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - top crate with valid
  • 335950 10/21/10 11:54 a-onl End_of_Run_7675
  • 335949 10/21/10 11:51 a-onl Start_Run_7675,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - top crate with valid
  • 335948 10/21/10 11:48 a-onl Start_Run_7674,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - top crate with valid
  • 335947 10/21/10 11:47 a-onl End_of_Run_7673
  • 335946 10/21/10 11:35 a-onl Start_Run_7673,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - top crate with valid
  • 335945 10/21/10 11:34 a-onl End_of_Run_7672
  • 335944 10/21/10 11:31 efuchey Hall A tools screenshot
  • 335943 10/21/10 11:30 a-onl Start_Run_7672,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - top crate with valid
  • 335942 10/21/10 11:28 gaskell Runplan for Thursday Day --> Swing
  • 335941 10/21/10 11:22 a-onl End_of_Run_7671
  • 335940 10/21/10 11:21 a-onl Start_Run_7669,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ
  • 335939 10/21/10 11:18 a-onl Start_Run_7670,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DAQ - top crate with valid
  • 335938 10/21/10 11:00 R. Michaels attempt to get scaler server on DVCS rocs
  • 335937 10/21/10 10:43 efuchey Beam available again
  • 335936 10/21/10 10:41 efuchey Alarm from DVCS HV
  • 335935 10/21/10 10:24 R. Michaels DVCSTopBot put at top of list
  • 335934 10/21/10 10:22 J. Zhang Moller magnet has been filled
  • 335933 10/21/10 09:23 hisham Hall in controlled access
  • 335932 10/21/10 09:14 Hisham Target status snapshot
  • 335931 10/21/10 09:04 Hisham ROC 14, ROC 17
  • 335930 10/21/10 09:02 A. Camsonne Calorimeter missing channels recovered
  • 335929 10/21/10 08:35 a-onl Start_Run_7667,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1.0 uA on kin3
  • 335928 10/21/10 08:25 A. Camsonne Power cycled DVCS crates
  • 335927 10/21/10 08:19 Carlos Pedestal shift in boards
  • 335926 10/21/10 08:09 a-onl Start_Run_7665,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1.0 uA on kin3
  • 335925 10/21/10 08:06 a-onl End_of_Run_7664
  • 335924 10/21/10 08:03 adaq Beam Time Accounting
  • 335923 10/21/10 08:03 horn Owl shift summary
  • 335922 10/21/10 07:51 ent end-of-run-7664 scaler screen
  • 335921 10/21/10 07:50 ent end-of-run-7664 ROC screens
  • 335920 10/21/10 07:27 A. Camsonne Waveforms block 35
  • 335919 10/21/10 07:14 a-onl Start_Run_7664,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1.0 uA on kin3
  • 335918 10/21/10 07:12 A. Camsonne Missing calorimeter channels 35 and 38
  • 335917 10/21/10 07:10 a-onl End_of_Run_7663
  • 335916 10/21/10 07:09 a-onl Start_Run_7663,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1.0 uA on kin3 - Test TrigDVCS config
  • 335915 10/21/10 07:07 A. Camsonne Run 7663 with TrigDVCS configuration
  • 335914 10/21/10 07:05 a-onl End_of_Run_7662
  • 335913 10/21/10 07:00 ent scaler screen run 7662
  • 335912 10/21/10 07:00 ent ROC screens
  • 335911 10/21/10 06:51 ent missing calo channels, snapshots
  • 335910 10/21/10 06:35 ent calo channels missing, runs 7660 etc.
  • 335909 10/21/10 06:22 a-onl Start_Run_7662,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1.0 uA on kin3
  • 335908 10/21/10 06:20 a-onl End_of_Run_7661
  • 335907 10/21/10 06:16 ent end-of-run-7661 scaler screen
  • 335906 10/21/10 06:15 ent end-of-run-7661 ROC screens
  • 335905 10/21/10 06:10 a-onl Start_Run_7661,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1.0 uA on kin3
  • 335904 10/21/10 05:34 a-onl End_of_Run_7660
  • 335903 10/21/10 04:49 a-onl Start_Run_7660,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1.0 uA on kin3
  • 335902 10/21/10 04:45 horn Configuration change to DVCS production
  • 335901 10/21/10 04:29 ent boiling test runs
  • 335900 10/21/10 04:21 a-onl Start_Run_7659,Run_type=Other,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=3.2,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 3.2 uA on kin3 target boiling raster 4 x 4
  • 335899 10/21/10 04:20 a-onl End_of_Run_7658
  • 335898 10/21/10 04:03 a-onl Start_Run_7658,Run_type=Other,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=3.2,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 3.2 uA on kin3 target boiling raster 4 x 4
  • 335897 10/21/10 03:36 a-onl End_of_Run_7657
  • 335896 10/21/10 03:34 a-onl Start_Run_7657,Run_type=Other,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=6.7,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 6.7 uA on kin3 target boiling raster 4 x 4
  • 335895 10/21/10 03:33 a-onl End_of_Run_7656
  • 335894 10/21/10 03:28 a-onl Start_Run_7656,Run_type=Other,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=6.7,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 6.7 uA on kin3 target boiling raster 4 x 4
  • 335893 10/21/10 03:27 a-onl End_of_Run_7655
  • 335892 10/21/10 03:20 a-onl Start_Run_7655,Run_type=Other,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=10.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 10 uA on kin3 target boiling raster 4 x 4
  • 335891 10/21/10 03:20 a-onl End_of_Run_7654
  • 335890 10/21/10 03:14 a-onl Start_Run_7654,Run_type=Other,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=10.0,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 10 uA on kin3 target boiling raster 4 x 4
  • 335889 10/21/10 03:04 a-onl End_of_Run_7653
  • 335888 10/21/10 03:00 a-onl Start_Run_7653,Run_type=Other,target_type=15cm dummy,beam_e0=5.5,beam_raster=on,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 4.5uA on kin3 target boiling raster 4 x 4
  • 335887 10/21/10 02:58 a-onl End_of_Run_7652
  • 335886 10/21/10 02:52 a-onl Start_Run_7652,Run_type=Other,target_type=15cm dummy,beam_e0=5.5,beam_raster=on,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 4.5uA on kin3 target boiling raster 4 x 4
  • 335885 10/21/10 02:39 a-onl End_of_Run_7651
  • 335884 10/21/10 02:37 a-onl Start_Run_7651,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 1.6uA on kin3 target boiling raster 4 x 4
  • 335883 10/21/10 02:36 a-onl End_of_Run_7650
  • 335882 10/21/10 02:33 a-onl Start_Run_7650,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 1.6uA on kin3 target boiling raster 4 x 4
  • 335881 10/21/10 02:31 a-onl End_of_Run_7649
  • 335880 10/21/10 02:27 a-onl Start_Run_7649,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 1.6uA on kin3 target boiling raster 2 x 2
  • 335879 10/21/10 02:26 a-onl End_of_Run_7648
  • 335878 10/21/10 02:20 a-onl Start_Run_7648,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 1.6uA on kin3 target boiling raster 2 x 2
  • 335877 10/21/10 02:16 a-onl End_of_Run_7647
  • 335876 10/21/10 02:11 a-onl Start_Run_7647,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 1.6uA on kin3 target boiling no raster
  • 335875 10/21/10 02:10 a-onl End_of_Run_7646
  • 335874 10/21/10 02:03 a-onl Start_Run_7646,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 1.6uA on kin3 target boiling no raster
  • 335873 10/21/10 01:57 a-onl End_of_Run_7645
  • 335872 10/21/10 01:52 a-onl Start_Run_7645,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 3.1uA on kin3 target boiling no raster
  • 335871 10/21/10 01:47 a-onl End_of_Run_7644
  • 335870 10/21/10 01:43 a-onl End_of_Run_7642
  • 335869 10/21/10 01:40 a-onl End_of_Run_7643
  • 335868 10/21/10 01:35 a-onl Start_Run_7643,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 3.1uA on kin3 target boiling raster 2 x 2
  • 335867 10/21/10 01:20 a-onl Start_Run_7641,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 4.5uA on kin3 target boiling raster 4 x 4
  • 335866 10/21/10 01:19 a-onl Start_Run_7642,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 4.5uA on kin3 target boiling raster 2 x 2
  • 335865 10/21/10 01:12 a-onl End_of_Run_7641
  • 335864 10/21/10 01:02 a-onl End_of_Run_7640
  • 335863 10/21/10 00:58 a-onl Start_Run_7640,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 4.5uA on kin3 target boiling no raster
  • 335862 10/21/10 00:57 a-onl End_of_Run_7639
  • 335861 10/21/10 00:51 a-onl Start_Run_7639,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 4.5uA on kin3 target boiling no raster
  • 335860 10/21/10 00:11 a-onl End_of_Run_7638
  • 335859 10/21/10 00:10 a-onl Start_Run_7638,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 T5=0 leftHRS 4.5uA on kin3 target boiling
  • 335858 10/21/10 00:08 a-onl End_of_Run_7637
  • 335857 10/21/10 00:00 adaq Beam Time Accounting
  • Wednesday

  • 335856 10/20/10 23:58 ZAFAR AHMED Wednesday 10-20-2010 Swing Shift Summary
  • 335855 10/20/10 23:50 a-onl Start_Run_7637,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 TrigDVCS 4.5uA on kin3
  • 335854 10/20/10 23:48 a-onl End_of_Run_7636
  • 335853 10/20/10 23:38 a-onl Start_Run_7636,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 TrigDVCS 4.5uA on kin3
  • 335852 10/20/10 23:37 a-onl End_of_Run_7635
  • 335851 10/20/10 23:22 Carlos ARS firmware version E09
  • 335850 10/20/10 22:57 a-onl Start_Run_7635,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 TrigDVCS 1uA on kin3
  • 335849 10/20/10 22:55 a-onl End_of_Run_7634
  • 335848 10/20/10 22:27 Nuruzzaman ROC 14 and 17 errors during run 7634.
  • 335847 10/20/10 22:12 A. Camsonne Added dvcs3 flag to TrigDVCS configuration
  • 335846 10/20/10 22:08 a-onl Start_Run_7634,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 TrigDVCS 1uA on kin3
  • 335845 10/20/10 22:00 A. Camsonne ROC18 crashed
  • 335844 10/20/10 21:53 a-onl Start_Run_7633,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 TrigDVCS 1uA on kin3
  • 335843 10/20/10 21:51 A. Camsonne Run with Calorimeter Trigger crate run 7633
  • 335842 10/20/10 21:51 a-onl End_of_Run_7632
  • 335841 10/20/10 21:36 A. Camsonne DVCS Calorimeter data event dump
  • 335840 10/20/10 21:17 a-onl Start_Run_7632,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3
  • 335839 10/20/10 21:16 a-onl End_of_Run_7631
  • 335838 10/20/10 20:14 a-onl Start_Run_7631,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3
  • 335837 10/20/10 20:13 a-onl End_of_Run_7630
  • 335836 10/20/10 20:09 a-onl Start_Run_7630,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3
  • 335835 10/20/10 19:59 a-onl Start_Run_7629,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=0.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 0.5uA on kin3
  • 335834 10/20/10 19:49 a-onl End_of_Run_7628
  • 335833 10/20/10 19:46 A. Camsonne Reverted to ARS version E08
  • 335832 10/20/10 19:41 a-onl Start_Run_7628,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=0.5,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 0.5uA on kin3
  • 335831 10/20/10 19:31 Nuruzzaman Rebooted ROCs and Restarted CODA
  • 335830 10/20/10 19:19 gaskell Runplan for Wednesday Swing --> Thursday Day
  • 335829 10/20/10 19:08 a-onl Start_Run_7625,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 2uA on kin3
  • 335828 10/20/10 19:07 a-onl End_of_Run_7624
  • 335827 10/20/10 19:04 Nuruzzaman Beam Current changed to 2 uA for run # 7625.
  • 335826 10/20/10 19:01 Nuruzzaman ROC 14 and 17 Screens during end of the run 7624.
  • 335825 10/20/10 19:01 Nuruzzaman Scaler and event rate during end of the run 7624.
  • 335824 10/20/10 18:34 a-onl Start_Run_7624,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3
  • 335823 10/20/10 18:27 Nuruzzaman ROC 14 and 17 Screens during end of the run 7623.
  • 335822 10/20/10 18:26 Nuruzzaman Scaler and event rate during end of the run 7623.
  • 335821 10/20/10 17:55 a-onl Start_Run_7623,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3 - No more FIFO clear
  • 335820 10/20/10 17:53 A. Camsonne Removed partial clear starting run 7623
  • 335819 10/20/10 17:50 a-onl End_of_Run_7622
  • 335818 10/20/10 17:49 a-onl Start_Run_7622,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3
  • 335817 10/20/10 17:16 a-onl Start_Run_7621,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.59,leftarm_theta=21.4,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3
  • 335816 10/20/10 17:15 a-onl End_of_Run_7620
  • 335815 10/20/10 17:04 efuchey rates variation
  • 335814 10/20/10 16:56 A. Camsonne DVCS T5 trigger with and without Cerenkov
  • 335813 10/20/10 16:49 a-onl Start_Run_7620,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3 -Test DAQ firmware version E09
  • 335812 10/20/10 16:42 efuchey Compton rates for today
  • 335811 10/20/10 16:42 a-onl End_of_Run_7619
  • 335810 10/20/10 16:40 a-onl Start_Run_7619,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3 -Test DAQ firmware version E09
  • 335809 10/20/10 16:39 A. Camsonne Run 7620 with version E09
  • 335808 10/20/10 16:38 A. Camsonne FPGA programmers on ARS
  • 335807 10/20/10 16:34 a-onl End_of_Run_7618
  • 335806 10/20/10 16:17 hyde DVCS trig outputs
  • 335805 10/20/10 16:15 A. Camsonne ARS firmware version E09 downloaded
  • 335804 10/20/10 16:01 Saha Shift Summary
  • 335803 10/20/10 16:00 adaq Beam Time Accounting
  • 335802 10/20/10 15:41 a-onl End_of_Run_7616
  • 335801 10/20/10 14:36 a-onl Start_Run_7616,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3
  • 335800 10/20/10 14:31 a-onl End_of_Run_7615
  • 335799 10/20/10 14:31 a-onl End_of_Run_7610
  • 335798 10/20/10 14:27 a-onl Start_Run_7615,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3
  • 335797 10/20/10 14:24 a-onl Start_Run_7614,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3
  • 335796 10/20/10 14:15 a-onl End_of_Run_7613
  • 335795 10/20/10 14:11 a-onl Start_Run_7613,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 1uA on kin3
  • 335794 10/20/10 14:11 a-onl End_of_Run_7612
  • 335793 10/20/10 14:11 hyde S1 S2 retiming
  • 335792 10/20/10 14:11 a-onl Start_Run_7612,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 DVCSTopBot 3uA on kin3
  • 335791 10/20/10 14:10 a-onl End_of_Run_7611
  • 335790 10/20/10 14:10 a-onl Start_Run_7611,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T5=1 LeftHRS 3uA on kin3
  • 335789 10/20/10 14:09 a-onl End_of_Run_7609
  • 335788 10/20/10 14:09 a-onl End_of_Run_7610
  • 335787 10/20/10 14:08 a-onl Start_Run_7610,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=T3=1 LeftHRS 3uA on kin3
  • 335786 10/20/10 14:06 a-onl Start_Run_7609,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=T5=3 LeftHRS 1uA on kin3
  • 335785 10/20/10 14:05 a-onl End_of_Run_7608
  • 335784 10/20/10 14:00 a-onl Start_Run_7608,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=T5=1 LeftHRS 1uA on kin3
  • 335783 10/20/10 13:57 a-onl End_of_Run_7607
  • 335782 10/20/10 13:53 a-onl Start_Run_7607,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=second run with beam after changing magnet polarity to detect electrons and HV on Calo is on now
  • 335781 10/20/10 13:48 a-onl End_of_Run_7606
  • 335780 10/20/10 13:26 a-onl Start_Run_7606,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=first run with beam after changing magnet polarity to detect electrons
  • 335779 10/20/10 13:21 a-onl End_of_Run_7605
  • 335778 10/20/10 13:09 Kiadtisak Target moved to LH2
  • 335777 10/20/10 13:04 Saha Harp Scans after beam recovery
  • 335776 10/20/10 13:00 R. Michaels disk quota problems on adaq
  • 335775 10/20/10 12:51 a-onl Start_Run_7605,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test daq
  • 335774 10/20/10 12:50 a-onl End_of_Run_7604
  • 335773 10/20/10 11:38 a-onl Start_Run_7604,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test daq
  • 335772 10/20/10 11:36 gaskell Change of kinematics
  • 335771 10/20/10 11:34 a-onl End_of_Run_7603
  • 335770 10/20/10 11:32 a-onl Start_Run_7603,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #1
  • 335769 10/20/10 11:08 hyde T1 input
  • 335768 10/20/10 10:53 auto_snap Radiation and Ozone snapshot
  • 335767 10/20/10 10:53 auto_snap Miscellaneous snapshot
  • 335765 10/20/10 10:53 auto_snap Compton Polarimeter snapshot
  • 335764 10/20/10 10:53 auto_snap Beam Position and Current snapshot
  • 335763 10/20/10 10:50 Fansler Ice on Moeller lines
  • 335762 10/20/10 10:38 A. Camsonne Changed scaler header in datastream
  • 335761 10/20/10 10:37 hyde S0/T2 input
  • 335760 10/20/10 10:16 hyde T2 double pulsing
  • 335759 10/20/10 10:06 J. Zhang Moller magnet has been filled
  • 335758 10/20/10 09:54 jpchen Summary of cryotarget events overnight and what TO should know
  • 335757 10/20/10 08:04 adaq Beam Time Accounting
  • 335756 10/20/10 07:36 gaskell Status and Controlled Access
  • 335755 10/20/10 07:33 hisham Hall in controlled access
  • 335754 10/20/10 07:31 Hisham Owl shift summary
  • 335753 10/20/10 07:11 Hisham Controlled access
  • 335752 10/20/10 07:03 Hisham Harp scan & BPM
  • 335751 10/20/10 06:26 florian itar target issue
  • 335750 10/20/10 05:49 Hisham Harp scan
  • 335749 10/20/10 05:05 Hisham Target update
  • 335748 10/20/10 04:44 a-onl End_of_Run_7600
  • 335747 10/20/10 04:38 Pawel run 7600 rates
  • 335746 10/20/10 04:37 Pawel run 7600 rocs
  • 335745 10/20/10 04:35 Hisham Target alarm
  • 335744 10/20/10 03:36 a-onl Start_Run_7600,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #1
  • 335743 10/20/10 03:33 a-onl End_of_Run_7599
  • 335742 10/20/10 03:29 Pawel run 7599 rocs
  • 335741 10/20/10 03:28 Pawel run 7599 rates (end)
  • 335740 10/20/10 03:26 florian itar target problem report
  • 335739 10/20/10 03:12 florian itar alarm report
  • 335738 10/20/10 02:53 auto_snap Radiation and Ozone snapshot
  • 335737 10/20/10 02:53 auto_snap Miscellaneous snapshot
  • 335736 10/20/10 02:53 auto_snap Hall A Tools snapshot
  • 335735 10/20/10 02:53 auto_snap Compton Polarimeter snapshot
  • 335734 10/20/10 02:53 auto_snap Beam Position and Current snapshot
  • 335733 10/20/10 02:27 a-onl Start_Run_7599,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #1
  • 335732 10/20/10 02:26 Pawel run 7599 rates
  • 335731 10/20/10 02:25 a-onl End_of_Run_7598
  • 335730 10/20/10 02:21 Pawel run 7598 rocs
  • 335729 10/20/10 01:31 a-onl Start_Run_7598,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #1
  • 335728 10/20/10 01:24 Pawel run 7598 rates
  • 335727 10/20/10 01:16 Pawel run 7597 rocs
  • 335726 10/20/10 00:26 a-onl Start_Run_7597,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #1
  • 335725 10/20/10 00:13 adaq Beam Time Accounting
  • 335724 10/20/10 00:12 Solvignon Swing shift summary
  • 335723 10/20/10 00:00 a-onl End_of_Run_7596
  • Tuesday

  • 335722 10/19/10 23:55 Nuruzzaman ROC 14 and 17 Screens during end of the run 7596.
  • 335721 10/19/10 23:55 Nuruzzaman Scaler and event rate during end of the run 7596.
  • 335720 10/19/10 23:54 mcanan swing target screens
  • 335719 10/19/10 23:18 a-onl Start_Run_7596,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #2
  • 335718 10/19/10 23:13 Nuruzzaman CODA showed error when I tried to end the run 7595.
  • 335717 10/19/10 23:07 Nuruzzaman ROC 14 and 17 Screens during end of the run 7595.
  • 335716 10/19/10 23:06 Nuruzzaman Scaler and event rate during end of the run 7595.
  • 335715 10/19/10 22:58 mcanan L2 Power is zero
  • 335714 10/19/10 22:37 a-onl Start_Run_7595,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #2
  • 335713 10/19/10 22:36 a-onl End_of_Run_7594
  • 335712 10/19/10 22:28 mcanan L2 Power fluctuates
  • 335711 10/19/10 22:24 mcanan Cryo Temps (green and red) dropped
  • 335710 10/19/10 22:21 Nuruzzaman ROC 14 and 17 Screens during end of the run 7594.
  • 335709 10/19/10 22:21 Nuruzzaman Scaler and event rate during end of the run 7594.
  • 335708 10/19/10 21:41 a-onl Start_Run_7594,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #2
  • 335707 10/19/10 21:35 Solvignon Kinematics changed to P0=2.406GeV/c, Theta=37.0
  • 335706 10/19/10 21:27 a-onl End_of_Run_7593
  • 335705 10/19/10 21:22 Nuruzzaman ROC 14 and 17 Screens during end of the run 7593.
  • 335704 10/19/10 21:21 Nuruzzaman Scaler and event rate during end of the run 7593.
  • 335703 10/19/10 21:11 a-onl Start_Run_7593,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #2
  • 335702 10/19/10 21:06 a-onl End_of_Run_7592
  • 335701 10/19/10 21:01 Nuruzzaman ROC 14 and 17 Screens during end of the run 7592.
  • 335700 10/19/10 21:00 Nuruzzaman Scaler and event rate during end of the run 7592.
  • 335699 10/19/10 20:52 efuchey Table scan attempt
  • 335698 10/19/10 20:30 a-onl Start_Run_7592,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #2
  • 335697 10/19/10 20:29 a-onl End_of_Run_7591
  • 335696 10/19/10 20:25 Nuruzzaman ROC 14 and 17 Screens.
  • 335695 10/19/10 20:24 Nuruzzaman Scaler and event rate during end of the run 7591.
  • 335694 10/19/10 19:45 a-onl Start_Run_7591,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #2
  • 335693 10/19/10 19:44 a-onl End_of_Run_7590
  • 335692 10/19/10 19:44 Nuruzzaman Scaler Rate after Run # 7590.
  • 335691 10/19/10 19:43 Nuruzzaman ROC 14 and 17 Error.
  • 335690 10/19/10 19:34 gaskell Runplan update for Tuesday Swing --> Wednesday Owl
  • 335689 10/19/10 19:27 Nuruzzaman Run # 7590: BPM A and B.
  • 335688 10/19/10 18:53 auto_snap Radiation and Ozone snapshot
  • 335687 10/19/10 18:53 auto_snap Miscellaneous snapshot
  • 335686 10/19/10 18:53 auto_snap Hall A Tools snapshot
  • 335685 10/19/10 18:53 auto_snap Compton Polarimeter snapshot
  • 335684 10/19/10 18:53 auto_snap Beam Position and Current snapshot
  • 335683 10/19/10 18:50 a-onl Start_Run_7590,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #2
  • 335682 10/19/10 18:45 Nuruzzaman Scaler Rate after Run # 7589.
  • 335681 10/19/10 18:44 Nuruzzaman ROC 14 and 17 error after run 7589
  • 335680 10/19/10 18:20 efuchey Rates for tonight
  • 335679 10/19/10 18:17 Nuruzzaman BPM A and B.
  • 335678 10/19/10 18:09 Nuruzzaman Analizer showing the following error.
  • 335677 10/19/10 18:03 a-onl Start_Run_7589,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #2
  • 335676 10/19/10 18:02 a-onl End_of_Run_7588
  • 335675 10/19/10 17:39 Nuruzzaman Problem in Calorimeter ADC channel 163 and 167 ?
  • 335674 10/19/10 17:21 Nuruzzaman Error message in ROC 14 and 17
  • 335673 10/19/10 17:04 R. Michaels scalers in ROC14 and 17
  • 335672 10/19/10 17:02 a-onl Start_Run_7588,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration #2
  • 335671 10/19/10 16:54 Solvignon Beam back
  • 335670 10/19/10 16:22 a-onl End_of_Run_7586
  • 335669 10/19/10 16:19 A. Camsonne ROC14 triggers
  • 335668 10/19/10 16:05 a-onl End_of_Run_7585
  • 335667 10/19/10 16:04 A. Camsonne Scaler in ROC4
  • 335665 10/19/10 16:03 a-onl End_of_Run_7584
  • 335666 10/19/10 16:03 adaq Beam Time Accounting
  • 335664 10/19/10 16:03 a-onl Start_Run_7584,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test DAQ
  • 335663 10/19/10 15:52 Hisham Day shift summary
  • 335662 10/19/10 15:37 a-onl Start_Run_7582,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test DAQ
  • 335661 10/19/10 15:37 a-onl End_of_Run_7582
  • 335660 10/19/10 15:11 A. Camsonne ARS in ROC14 not accepting triggers
  • 335659 10/19/10 15:06 a-onl End_of_Run_7580
  • 335658 10/19/10 15:06 a-onl Start_Run_7580,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test DAQ
  • 335657 10/19/10 15:06 a-onl Start_Run_7579,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test DAQ
  • 335656 10/19/10 15:04 hisham Controlled access
  • 335655 10/19/10 15:00 a-onl End_of_Run_7579
  • 335654 10/19/10 15:00 a-onl End_of_Run_7577
  • 335653 10/19/10 14:53 a-onl End_of_Run_7578
  • 335652 10/19/10 14:52 a-onl Start_Run_7578,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting 2
  • 335651 10/19/10 14:50 a-onl Start_Run_7577,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting 2
  • 335650 10/19/10 14:33 a-onl Start_Run_7576,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting 2
  • 335649 10/19/10 14:09 a-onl Start_Run_7575,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting 2
  • 335648 10/19/10 14:05 a-onl End_of_Run_7574
  • 335647 10/19/10 14:02 Hisham Empty FIFO, ROC14,Run7574
  • 335646 10/19/10 13:53 Kiadtisak Target moved to LH2
  • 335645 10/19/10 13:44 Hisham Beam Recovery
  • 335644 10/19/10 13:13 hyde DVCS DAQ deadtime study
  • 335643 10/19/10 12:58 HIsham Beam update
  • 335642 10/19/10 12:51 a-onl Start_Run_7572,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test DAQ
  • 335641 10/19/10 12:50 a-onl End_of_Run_7571
  • 335640 10/19/10 12:41 a-onl Start_Run_7571,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test DAQ
  • 335639 10/19/10 12:33 A. Camsonne Scalers added to VME readout
  • 335638 10/19/10 11:57 gaskell Update and plan for Day and Swing shift
  • 335637 10/19/10 11:50 hyde S0 HV
  • 335636 10/19/10 11:46 Kiadtisak Target moved to empty
  • 335635 10/19/10 11:46 Hisham Beam recovery procedure
  • 335634 10/19/10 11:46 Kiadtisak Target moved to empty
  • 335633 10/19/10 11:45 hyde S0 S2m scalers in top ARS VME crate
  • 335632 10/19/10 11:42 A. Camsonne ARS readout to reset the FIFO at every event again
  • 335631 10/19/10 11:28 A. Camsonne Clocks on scalers
  • 335630 10/19/10 11:27 a-onl Start_Run_7570,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test DAQ Deadtime study S0A HV 950 S0 at 80 Hz
  • 335629 10/19/10 11:25 a-onl End_of_Run_7569
  • 335628 10/19/10 11:20 a-onl Start_Run_7569,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test DAQ
  • 335627 10/19/10 11:06 a-onl Start_Run_7568,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Deadtime study, S0 copied to S2m in DVCS trigger S0A HV 950V, rate=80Hz
  • 335626 10/19/10 11:06 a-onl End_of_Run_7567
  • 335625 10/19/10 10:59 J. Zhang Moller magnet has been filled
  • 335624 10/19/10 10:53 auto_snap Radiation and Ozone snapshot
  • 335623 10/19/10 10:52 auto_snap Miscellaneous snapshot
  • 335622 10/19/10 10:52 auto_snap Hall A Tools snapshot
  • 335621 10/19/10 10:52 auto_snap Compton Polarimeter snapshot
  • 335620 10/19/10 10:52 auto_snap Beam Position and Current snapshot
  • 335619 10/19/10 10:51 a-onl Start_Run_7567,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Deadtime study, S0 copied to S2m in DVCS trigger S0A HV 950V, rate=80Hz, prescale removed by Magali?
  • 335618 10/19/10 10:46 A. Camsonne Upper level Left HRS hut scope on hallavme12
  • 335617 10/19/10 10:38 A. Camsonne DVCS crates scalers assignment
  • 335616 10/19/10 10:23 a-onl End_of_Run_7563
  • 335615 10/19/10 10:21 a-onl End_of_Run_7562
  • 335614 10/19/10 09:57 a-onl Start_Run_7562,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Deadtime study, S0 copied to S2m in DVCS trigger S0A HV 1050V, rate=200Hz
  • 335613 10/19/10 09:55 a-onl End_of_Run_7561
  • 335612 10/19/10 09:50 a-onl Start_Run_7561,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Deadtime study, S0 copied to S2m in DVCS trigger S0A HV 1000V, rate=120Hz
  • 335611 10/19/10 09:47 a-onl End_of_Run_7560
  • 335610 10/19/10 09:46 a-onl Start_Run_7560,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Deadtime study, S0 copied to S2m in DVCS trigger S0A HV 1100V, rate=300Hz
  • 335609 10/19/10 09:44 a-onl End_of_Run_7559
  • 335608 10/19/10 09:43 a-onl Start_Run_7559,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Deadtime study, S0 copied to S2m in DVCS trigger S0A HV 1100V, rate=300Hz
  • 335607 10/19/10 08:55 Carlos ROCs at the end of run 7557
  • 335606 10/19/10 08:31 jroche Owl shift summary
  • 335605 10/19/10 08:29 a-onl End_of_Run_7557
  • 335604 10/19/10 08:25 jroche Hall in controlled access
  • 335603 10/19/10 08:11 jroche N2 level in Q3 really low
  • 335602 10/19/10 08:03 adaq Beam Time Accounting
  • 335601 10/19/10 07:59 jroche Turn on pion rejector 1 back on
  • 335600 10/19/10 07:40 Pawel run 7557 roc17
  • 335599 10/19/10 07:32 a-onl Start_Run_7557,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic Calibration-I
  • 335598 10/19/10 07:26 yez a plot this shows the problem of pion rejectors
  • 335597 10/19/10 07:24 Pawel run 7557 roc14
  • 335596 10/19/10 07:17 Pawel run 7556 errors
  • 335595 10/19/10 07:13 jroche pion rejector (HV off for a test)
  • 335594 10/19/10 07:11 Pawel run 7555 roc17
  • 335593 10/19/10 07:07 a-onl Start_Run_7555,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic Calibration-I
  • 335592 10/19/10 06:58 Pawel run 7554 errors
  • 335591 10/19/10 06:47 a-onl Start_Run_7554,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic Calibration-I
  • 335590 10/19/10 06:45 a-onl End_of_Run_7553
  • 335589 10/19/10 06:41 jroche roc 14 and 17 at the end of run 7553
  • 335588 10/19/10 06:25 Pawel run 7553 roc 14 error
  • 335587 10/19/10 06:00 Pawel run 7553 roc17 error
  • 335586 10/19/10 05:56 Pawel coda and raw rates at 2 uA
  • 335585 10/19/10 05:48 a-onl Start_Run_7553,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic Calibration-I
  • 335584 10/19/10 05:48 Pawel coda and raw rates @ 3 uA
  • 335583 10/19/10 05:35 Pawel run 7551 errors
  • 335582 10/19/10 05:08 a-onl Start_Run_7551,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic Calibration-I
  • 335581 10/19/10 05:07 a-onl Start_Run_7544,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic Calibration-I
  • 335580 10/19/10 05:06 a-onl End_of_Run_7544
  • 335579 10/19/10 05:00 Pawel run 7551 roc17
  • 335578 10/19/10 04:55 a-onl Start_Run_7550,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335577 10/19/10 04:50 yez target screen 2
  • 335576 10/19/10 04:49 yez target screen 1
  • 335575 10/19/10 04:39 a-onl End_of_Run_7549
  • 335574 10/19/10 04:38 a-onl Start_Run_7548,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335573 10/19/10 04:38 a-onl Start_Run_7549,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335572 10/19/10 04:35 a-onl End_of_Run_7548
  • 335571 10/19/10 04:29 a-onl End_of_Run_7547
  • 335570 10/19/10 04:27 a-onl Start_Run_7547,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335569 10/19/10 04:21 jroche that was 4 uA
  • 335568 10/19/10 04:19 a-onl End_of_Run_7546
  • 335567 10/19/10 04:18 a-onl Start_Run_7546,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335566 10/19/10 04:16 a-onl End_of_Run_7545
  • 335565 10/19/10 04:12 a-onl Start_Run_7545,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic Calibration-I
  • 335564 10/19/10 04:01 a-onl End_of_Run_7543
  • 335563 10/19/10 03:54 Carlos Coda and raw rates for 3uA
  • 335562 10/19/10 03:45 a-onl Start_Run_7543,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic Calibration-I
  • 335561 10/19/10 03:28 yez target moved to LH2 15cm in Loop2
  • 335560 10/19/10 03:22 yez target moved to BeO
  • 335559 10/19/10 03:17 jroche it's freezing in here !
  • 335558 10/19/10 03:06 Jack NMR scope signal
  • 335557 10/19/10 02:54 jroche Compton dipole problem: decided to bypass the chicane all together
  • 335556 10/19/10 02:52 auto_snap Radiation and Ozone snapshot
  • 335555 10/19/10 02:52 auto_snap Miscellaneous snapshot
  • 335554 10/19/10 02:52 auto_snap Hall A Tools snapshot
  • 335553 10/19/10 02:52 auto_snap Compton Polarimeter snapshot
  • 335552 10/19/10 02:52 auto_snap Beam Position and Current snapshot
  • 335551 10/19/10 02:35 jroche Status of Hall A quite misleading
  • 335550 10/19/10 02:03 efuchey QWP scan for today
  • 335549 10/19/10 01:52 jroche difficult beam recovery.
  • 335548 10/19/10 01:17 a-onl End_of_Run_7542
  • 335547 10/19/10 00:30 a-onl Start_Run_7542,Run_type=Calibration,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test
  • 335546 10/19/10 00:13 a-onl End_of_Run_7539
  • 335545 10/19/10 00:10 jroche NMR gui screen shot
  • 335543 10/19/10 00:02 a-onl Start_Run_7539,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335544 10/19/10 00:02 adaq Beam Time Accounting
  • 335542 10/19/10 00:00 Nuruzzaman Target GUI and stripchart screenshot at end of swing shift.
  • Monday

  • 335541 10/18/10 23:59 efuchey redid QWP scan analysis from 10-16
  • 335540 10/18/10 23:49 efuchey Results of QWP scan from 10/16/2010
  • 335539 10/18/10 23:42 BoZhao Swing shift summary
  • 335538 10/18/10 23:42 a-onl Start_Run_7534,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS - with clear ROC14 and ROC17 - Cleared when data shifts or FIFO are empty
  • 335537 10/18/10 23:35 BoZhao dipole NMR signal
  • 335536 10/18/10 23:23 efuchey Results of QWP scan from 10/15/2010
  • 335535 10/18/10 23:08 a-onl Start_Run_7532,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS - with clear ROC14 and ROC17
  • 335534 10/18/10 23:07 Jack (Techs) LCW pressure regulator for Arc/BSY feed from Hall A is shot
  • 335533 10/18/10 23:06 A. Camsonne ARS board resetting FIFO after each event
  • 335532 10/18/10 22:56 Jack (Techs) Bogie controls issue
  • 335531 10/18/10 22:48 gaskell Runplan for Monday swing to Tuesday day
  • 335530 10/18/10 22:44 Jack Magnets status
  • 335529 10/18/10 22:44 A. Camsonne reset fan board moved to ROC14
  • 335528 10/18/10 22:30 a-onl End_of_Run_7528
  • 335527 10/18/10 22:28 a-onl Start_Run_7528,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS - with clear ROC14 and ROC17
  • 335526 10/18/10 22:27 a-onl End_of_Run_7527
  • 335525 10/18/10 22:06 a-onl Start_Run_7527,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS - with clear ROC14 and ROC17
  • 335524 10/18/10 22:03 a-onl End_of_Run_7526
  • 335523 10/18/10 22:02 a-onl Start_Run_7526,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS - with clear ROC14 and ROC17
  • 335522 10/18/10 22:00 a-onl End_of_Run_7525
  • 335521 10/18/10 22:00 a-onl End_of_Run_7523
  • 335520 10/18/10 22:00 a-onl Start_Run_7524,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS - with clear ROC14 and ROC17
  • 335519 10/18/10 22:00 a-onl End_of_Run_7524
  • 335518 10/18/10 21:57 a-onl Start_Run_7523,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS - with clear ROC14 and ROC17
  • 335517 10/18/10 21:55 a-onl End_of_Run_7522
  • 335516 10/18/10 21:50 a-onl Start_Run_7522,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS - with clear ROC14 and ROC17
  • 335515 10/18/10 21:49 a-onl End_of_Run_7521
  • 335514 10/18/10 21:46 a-onl End_of_Run_7519
  • 335513 10/18/10 21:46 a-onl Start_Run_7521,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS - with clear ROC14 and ROC17
  • 335512 10/18/10 21:42 BoZhao LHRS magnets down
  • 335511 10/18/10 21:40 a-onl End_of_Run_7520
  • 335510 10/18/10 21:37 a-onl Start_Run_7520,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS - with clear ROC14 only
  • 335509 10/18/10 21:37 a-onl End_of_Run_7518
  • 335508 10/18/10 21:33 a-onl Start_Run_7519,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS - with clear
  • 335507 10/18/10 21:26 a-onl Start_Run_7518,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS - with clear
  • 335506 10/18/10 21:25 a-onl End_of_Run_7517
  • 335505 10/18/10 21:23 a-onl Start_Run_7517,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS
  • 335504 10/18/10 21:14 a-onl End_of_Run_7514
  • 335503 10/18/10 20:15 a-onl Start_Run_7513,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test magali
  • 335502 10/18/10 19:47 a-onl End_of_Run_7511
  • 335501 10/18/10 19:47 a-onl End_of_Run_7509
  • 335500 10/18/10 19:39 gaskell LCW update
  • 335499 10/18/10 18:52 auto_snap Radiation and Ozone snapshot
  • 335498 10/18/10 18:52 auto_snap Miscellaneous snapshot
  • 335497 10/18/10 18:52 auto_snap Hall A Tools snapshot
  • 335496 10/18/10 18:52 auto_snap Compton Polarimeter snapshot
  • 335495 10/18/10 18:52 auto_snap Beam Position and Current snapshot
  • 335494 10/18/10 18:02 a-onl Start_Run_7511,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test magali
  • 335493 10/18/10 17:24 a-onl End_of_Run_7510
  • 335492 10/18/10 16:51 a-onl Start_Run_7510,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test magali
  • 335491 10/18/10 16:34 a-onl Start_Run_7509,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test magali
  • 335490 10/18/10 16:08 Hisham Correction: BCM calibration
  • 335489 10/18/10 16:04 adaq Beam Time Accounting
  • 335488 10/18/10 16:04 Nuruzzaman Target GUI and strip chart screenshot at the begining of the swing shift.
  • 335487 10/18/10 16:02 ellie Day Shift Summary
  • 335486 10/18/10 15:59 A. Camsonne 2 ARS board surveyed and take to the CODA DAQ lab
  • 335485 10/18/10 15:28 O.Glamazdin Moller magnet has been filled
  • 335484 10/18/10 15:07 jpchen cryotarget: restoring to normal running condition
  • 335483 10/18/10 14:56 ellie LHRS Q1 flow giving yellow, occasionally red, alerts
  • 335482 10/18/10 14:31 Hisham BCM Calibration
  • 335481 10/18/10 14:04 Jack (Techs) Resets for the Bogies controls
  • 335480 10/18/10 13:33 a-onl End_of_Run_7508
  • 335479 10/18/10 13:24 a-onl Start_Run_7508,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=S0A HV 1200 V, S0 rate 600 hz S0 copied to S2 in DVCS daq DVCS cosmic HV off Deadtime study
  • 335478 10/18/10 13:22 a-onl End_of_Run_7507
  • 335477 10/18/10 13:17 a-onl End_of_Run_7506
  • 335476 10/18/10 13:05 a-onl Start_Run_7506,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=S0A HV 1000 V, S0 rate 110 hz S0 copied to S2 in DVCS daq DVCS cosmic HV off Deadtime study
  • 335475 10/18/10 13:03 a-onl End_of_Run_7505
  • 335474 10/18/10 12:58 a-onl Start_Run_7505,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=S0A HV 900 V, S0 rate 45 hz S0 copied to S2 in DVCS daq DVCS cosmic HV off Deadtime study
  • 335473 10/18/10 12:41 hyde deadtime study
  • 335472 10/18/10 12:41 a-onl Start_Run_7504,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=S0A HV 950 V, S0 rate 75 hz S0 copied to S2 in DVCS daq DVCS cosmic HV off Deadtime study
  • 335471 10/18/10 12:40 a-onl End_of_Run_7503
  • 335470 10/18/10 12:35 hyde DVCS Cosmic HV Off
  • 335469 10/18/10 12:28 Carlos Owl shift summary
  • 335468 10/18/10 12:22 a-onl Start_Run_7503,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=S0A HV 950 V, S0 rate 75 hz S0 copied to S2 in DVCS daq Deadtime study
  • 335467 10/18/10 12:19 a-onl End_of_Run_7502
  • 335466 10/18/10 12:08 a-onl Start_Run_7502,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=S0A HV 1000 V, S0 copied to S2 in DVCS daq Deadtime study
  • 335465 10/18/10 12:07 a-onl End_of_Run_7501
  • 335464 10/18/10 11:58 a-onl Start_Run_7501,Run_type=Check out,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=S0A HV 1000 V, S0 plugged to S2 in DVCS daq
  • 335463 10/18/10 11:55 jroche DVCS2 Target informations
  • 335462 10/18/10 11:39 A. Camsonne output modified
  • 335461 10/18/10 11:33 ellie HV Change for S2m L/R 12-15, Channel 8
  • 335460 10/18/10 11:33 a-onl End_of_Run_7500
  • 335459 10/18/10 11:31 hyde S0 HV
  • 335458 10/18/10 10:52 auto_snap Radiation and Ozone snapshot
  • 335457 10/18/10 10:52 auto_snap Miscellaneous snapshot
  • 335456 10/18/10 10:51 auto_snap Hall A Tools snapshot
  • 335455 10/18/10 10:51 auto_snap Compton Polarimeter snapshot
  • 335454 10/18/10 10:51 auto_snap Beam Position and Current snapshot
  • 335453 10/18/10 10:49 gaskell Status update
  • 335452 10/18/10 10:23 a-onl Start_Run_7500,Run_type=Cosmics,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Cosmics HV at 600 V
  • 335451 10/18/10 10:21 A. Camsonne DVCS calorimeter paddle on for cosmics data taking - Run 7500
  • 335450 10/18/10 10:02 hyde S1.and.S2 timing into T3: Correction
  • 335449 10/18/10 09:56 adev End_of_Run_20021
  • 335448 10/18/10 08:52 A. Camsonne DVCS DAQ meeting summary
  • 335447 10/18/10 08:41 adaq Beam Time Accounting
  • 335446 10/18/10 07:55 jmammei target status at start of shift
  • 335445 10/18/10 06:47 a-onl End_of_Run_7499
  • 335444 10/18/10 06:44 yez target moved to Empty
  • 335443 10/18/10 06:04 yez target screen 2
  • 335442 10/18/10 06:03 yez target screen 1
  • 335441 10/18/10 05:45 a-onl Start_Run_7499,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-III
  • 335440 10/18/10 05:37 a-onl End_of_Run_7498
  • 335439 10/18/10 05:04 a-onl Start_Run_7498,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-II
  • 335438 10/18/10 03:32 a-onl Start_Run_7497,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-II
  • 335437 10/18/10 03:31 Carlos Bogies troubles (TECHS)
  • 335436 10/18/10 03:28 Carlos Elastic setting 2
  • 335435 10/18/10 02:51 auto_snap Radiation and Ozone snapshot
  • 335434 10/18/10 02:51 auto_snap Miscellaneous snapshot
  • 335433 10/18/10 02:51 auto_snap Hall A Tools snapshot
  • 335432 10/18/10 02:51 auto_snap Compton Polarimeter snapshot
  • 335431 10/18/10 02:51 auto_snap Beam Position and Current snapshot
  • 335430 10/18/10 02:23 a-onl End_of_Run_7496
  • 335429 10/18/10 02:17 Pawel Nadel- run 7496
  • 335428 10/18/10 02:04 a-onl Start_Run_7496,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-I
  • 335427 10/18/10 02:03 a-onl End_of_Run_7495
  • 335426 10/18/10 01:57 Pawel Nadel- run 7495
  • 335425 10/18/10 01:46 a-onl Start_Run_7495,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-I
  • 335424 10/18/10 01:35 Pawel Nadel- run 7494
  • 335423 10/18/10 01:11 Carlos ARS sync problem
  • 335422 10/18/10 00:57 Pawel Nadel- run 7493
  • 335421 10/18/10 00:38 a-onl Start_Run_7493,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-I
  • 335420 10/18/10 00:29 Carlos Run 7491
  • 335419 10/18/10 00:11 a-onl Start_Run_7491,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-I
  • 335418 10/18/10 00:10 a-onl End_of_Run_7490
  • 335417 10/18/10 00:09 adaq Beam Time Accounting
  • 335416 10/18/10 00:05 Carlos fifo 7490
  • 335415 10/18/10 00:03 LeRose Shift Summary (Swing)
  • Sunday

  • 335414 10/17/10 23:43 a-onl Start_Run_7490,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-I
  • 335413 10/17/10 23:43 Abdurahim Loop 2 temperature sensor readout fluctuation
  • 335412 10/17/10 23:37 Abdurahim temperature sensor readout fluctuation
  • 335411 10/17/10 23:30 Ramesh fifo from 7489
  • 335410 10/17/10 23:28 a-onl Start_Run_7489,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-I
  • 335409 10/17/10 23:10 a-onl Start_Run_7488,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-I
  • 335408 10/17/10 23:08 a-onl End_of_Run_7487
  • 335407 10/17/10 23:04 Ramesh Run 7487 stopped due to fifo appearance
  • 335406 10/17/10 23:03 Abdurahim Compton
  • 335405 10/17/10 22:51 a-onl Start_Run_7487,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-I
  • 335404 10/17/10 22:42 Ramesh ROC17 stopped run 7486
  • 335403 10/17/10 22:18 a-onl Start_Run_7486,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-I
  • 335402 10/17/10 22:14 a-onl End_of_Run_7485
  • 335401 10/17/10 22:14 Ramesh Scalar rate during run 7486 @ 3 uA
  • 335400 10/17/10 22:08 LeRose Beam is back
  • 335399 10/17/10 22:07 LeRose Quad Cycling
  • 335398 10/17/10 22:06 Abdurahim Target Snapshot
  • 335397 10/17/10 22:05 Ramesh Run 7485 mostly a cosmic run
  • 335396 10/17/10 21:30 LeRose No Beam!
  • 335395 10/17/10 20:47 Carlos Calorimeter HV updated
  • 335394 10/17/10 20:42 a-onl Start_Run_7485,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration setting-I
  • 335393 10/17/10 20:27 a-onl Start_Run_7484,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.748,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335392 10/17/10 20:25 Ramesh anti-correlation of momentum and phi angle
  • 335391 10/17/10 20:20 Abdurahim Target -> LH2 (Loop2, 15 cm)
  • 335390 10/17/10 20:18 LeRose Beam is back
  • 335389 10/17/10 20:07 Abdurahim Target -> Empty
  • 335388 10/17/10 20:04 Abdurahim Target screenshot
  • 335387 10/17/10 19:57 LeRose Spectrometer Settings
  • 335386 10/17/10 19:48 LeRose Controlled Access
  • 335385 10/17/10 19:35 a-onl End_of_Run_7480
  • 335384 10/17/10 18:51 auto_snap Radiation and Ozone snapshot
  • 335383 10/17/10 18:51 auto_snap Miscellaneous snapshot
  • 335382 10/17/10 18:51 auto_snap Hall A Tools snapshot
  • 335381 10/17/10 18:51 auto_snap Compton Polarimeter snapshot
  • 335380 10/17/10 18:51 auto_snap Beam Position and Current snapshot
  • 335379 10/17/10 18:20 Carlos Sync test run
  • 335378 10/17/10 18:12 a-onl Start_Run_7480,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=0,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=ARS test - validation on
  • 335377 10/17/10 18:08 Abdurahim Cryogen Flow Snapshots
  • 335376 10/17/10 18:07 Abdurahim Target Loops Snapshots
  • 335375 10/17/10 18:05 Abdurahim Target Screenshot
  • 335374 10/17/10 17:43 Abdurahim Target -> H2 (Loop 2, 15 cm)
  • 335373 10/17/10 17:42 A. Camsonne Clock switched
  • 335372 10/17/10 17:33 A. Camsonne ARS setup to use validation signal
  • 335371 10/17/10 17:30 Abdurahim Target -> BeO
  • 335370 10/17/10 17:19 LeRose NMR Lock
  • 335369 10/17/10 17:12 gaskell Run plan for Sunday swing and Monday owl
  • 335368 10/17/10 17:04 A. Camsonne Compton photon rates background at 4 KHz/uA
  • 335367 10/17/10 16:40 A. Camsonne Timeout in ROC14 and ROC17 set to 300
  • 335366 10/17/10 16:39 LeRose NMR Lock
  • 335365 10/17/10 16:31 Abdurahim Target -> Empty
  • 335364 10/17/10 16:01 a-onl Start_Run_7477,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=0,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=cosmic trigger
  • 335363 10/17/10 16:01 adaq Beam Time Accounting
  • 335362 10/17/10 15:59 jmammei Day Shift summary
  • 335361 10/17/10 15:52 cshloka loop 2 change in temp in cernox_2_R
  • 335360 10/17/10 15:44 cshloka empty histos in some calo TDCs
  • 335359 10/17/10 15:41 cshloka empty histos for calo adcs
  • 335358 10/17/10 15:35 cshloka no counts in some calo ADCs and TDCs
  • 335357 10/17/10 15:30 a-onl Start_Run_7476,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=0,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=cosmic trigger
  • 335356 10/17/10 15:29 a-onl End_of_Run_7475
  • 335355 10/17/10 15:10 a-onl Start_Run_7475,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=0,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test clock in tdc T5 only - LED pulser on
  • 335354 10/17/10 15:10 a-onl End_of_Run_7474
  • 335353 10/17/10 15:02 A. Camsonne Run test pulser 7474
  • 335352 10/17/10 14:59 a-onl Start_Run_7474,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=0,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test clock in tdc T5 only - LED pulser on
  • 335351 10/17/10 14:40 a-onl Start_Run_7473,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=0,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test clock in tdc T5 only
  • 335350 10/17/10 14:39 a-onl End_of_Run_7472
  • 335349 10/17/10 14:38 a-onl Start_Run_7472,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=0,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test clock in tdc T3 only
  • 335348 10/17/10 14:37 cshloka target status
  • 335347 10/17/10 14:33 a-onl Start_Run_7471,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=0,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test clock in tdc T3 only
  • 335346 10/17/10 14:31 a-onl End_of_Run_7470
  • 335345 10/17/10 14:25 a-onl Start_Run_7470,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=0,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test clock in tdc
  • 335344 10/17/10 14:01 cshloka empty fifo
  • 335343 10/17/10 13:49 jmammei MOLLER refill complete
  • 335342 10/17/10 13:41 J. Zhang Moller magnet has been filled
  • 335341 10/17/10 13:26 a-onl Start_Run_7469,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=0,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test run
  • 335340 10/17/10 13:03 a-onl Start_Run_7466,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=0,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=test
  • 335339 10/17/10 13:00 jmammei survey complete; Hall in Controlled access
  • 335338 10/17/10 12:38 cshloka no counts in some calo TDCs
  • 335337 10/17/10 12:34 cshloka no counts in some calo ADCs
  • 335336 10/17/10 12:28 gaskell Called for Controlled Access
  • 335335 10/17/10 11:45 a-onl Start_Run_7465,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=elastic calibration - test trigger config
  • 335334 10/17/10 11:30 a-onl Start_Run_7463,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=elastic calibration - test trigger config
  • 335333 10/17/10 11:30 cshloka target status
  • 335332 10/17/10 11:26 A. Camsonne Run 7463 - test calorimeter trigger read out
  • 335331 10/17/10 11:22 A. Camsonne removing DVCS root files on disk
  • 335330 10/17/10 11:17 a-onl End_of_Run_7462
  • 335329 10/17/10 11:17 A. Camsonne /adaql4/work1 full
  • 335328 10/17/10 11:11 A. Camsonne adaql7 NFS issues
  • 335327 10/17/10 11:05 a-onl Start_Run_7462,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=elastic calibration
  • 335326 10/17/10 11:02 a-onl End_of_Run_7461
  • 335325 10/17/10 10:58 cshloka empty fifo for roc 14 and 17 at the end of run# 7461
  • 335324 10/17/10 10:51 auto_snap Radiation and Ozone snapshot
  • 335323 10/17/10 10:51 auto_snap Miscellaneous snapshot
  • 335322 10/17/10 10:51 auto_snap Hall A Tools snapshot
  • 335321 10/17/10 10:50 auto_snap Compton Polarimeter snapshot
  • 335320 10/17/10 10:50 auto_snap Beam Position and Current snapshot
  • 335319 10/17/10 10:39 gaskell Status update
  • 335318 10/17/10 10:00 cshloka empty fifo in roc 14 and 17
  • 335317 10/17/10 09:57 cshloka Empty Fifo error in ROC 17
  • 335316 10/17/10 09:55 cshloka ion rejector sum plots are blank for prl1.
  • 335315 10/17/10 09:39 jmammei changed the energy in the db_run.dat
  • 335314 10/17/10 09:32 jmammei p0=2.607; LHRS @ 34.760 degrees
  • 335313 10/17/10 09:16 a-onl Start_Run_7461,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335312 10/17/10 09:12 a-onl End_of_Run_7460
  • 335311 10/17/10 09:09 A. Camsonne DVCS clock in TDC seen now
  • 335310 10/17/10 08:57 brads Beam energy a bit unstable
  • 335309 10/17/10 08:52 a-onl Start_Run_7460,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335308 10/17/10 08:33 adaq Beam Time Accounting
  • 335307 10/17/10 08:32 cshloka target status
  • 335306 10/17/10 08:17 efuchey Q1 + D alarms
  • 335305 10/17/10 08:09 efuchey 10/17/10 Owl Shift summary
  • 335304 10/17/10 07:43 a-onl End_of_Run_7459
  • 335303 10/17/10 07:41 A. Camsonne Requested access for HRS move back to elastic kinematics
  • 335302 10/17/10 07:40 A. Camsonne No signal seen on D.dvcsClk
  • 335301 10/17/10 07:36 a-onl Start_Run_7459,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=ARS tests clock check
  • 335300 10/17/10 07:33 a-onl End_of_Run_7458
  • 335299 10/17/10 07:32 a-onl Start_Run_7450,Run_type=Calibration,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=4.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test, raster 2.75x2 (MCC)
  • 335298 10/17/10 07:26 A. Camsonne DVCS DAQ tests
  • 335297 10/17/10 06:24 a-onl End_of_Run_7457
  • 335296 10/17/10 06:23 D.Wang target status
  • 335295 10/17/10 06:16 a-onl Start_Run_7457,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test, raster 2.75x2
  • 335294 10/17/10 06:11 a-onl End_of_Run_7456
  • 335293 10/17/10 06:05 a-onl Start_Run_7456,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test, raster 5.5x4
  • 335292 10/17/10 05:58 a-onl End_of_Run_7455
  • 335291 10/17/10 05:51 a-onl Start_Run_7455,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test, raster off
  • 335290 10/17/10 05:48 a-onl End_of_Run_7454
  • 335289 10/17/10 05:41 a-onl Start_Run_7454,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=3,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test, raster 2.75x2
  • 335288 10/17/10 05:37 a-onl End_of_Run_7453
  • 335287 10/17/10 05:31 a-onl Start_Run_7453,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=3,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test, raster 5.5x4
  • 335286 10/17/10 05:26 a-onl End_of_Run_7452
  • 335285 10/17/10 05:16 a-onl Start_Run_7452,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test, raster off
  • 335284 10/17/10 05:14 a-onl End_of_Run_7451
  • 335283 10/17/10 05:07 a-onl Start_Run_7451,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=4.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test, raster 2.75x2 (MCC)
  • 335282 10/17/10 05:06 a-onl End_of_Run_7450
  • 335281 10/17/10 05:01 D.Wang target moved to Loop2, 15cm
  • 335280 10/17/10 04:55 a-onl End_of_Run_7449
  • 335279 10/17/10 04:46 a-onl Start_Run_7449,Run_type=Calibration,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=1.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test, raster 2.75x2 (MCC)
  • 335278 10/17/10 04:44 a-onl End_of_Run_7448
  • 335277 10/17/10 04:33 a-onl Start_Run_7448,Run_type=Calibration,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=1.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test
  • 335276 10/17/10 04:24 a-onl End_of_Run_7446
  • 335275 10/17/10 04:09 a-onl Start_Run_7446,Run_type=Calibration,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=3.0,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test
  • 335274 10/17/10 04:01 a-onl End_of_Run_7436
  • 335273 10/17/10 03:57 a-onl End_of_Run_7445
  • 335272 10/17/10 03:50 a-onl Start_Run_7445,Run_type=Calibration,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=3.0,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test
  • 335271 10/17/10 03:42 a-onl End_of_Run_7444
  • 335270 10/17/10 03:31 a-onl Start_Run_7444,Run_type=Calibration,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=3.0,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test
  • 335269 10/17/10 03:27 a-onl End_of_Run_7443
  • 335268 10/17/10 03:17 a-onl Start_Run_7443,Run_type=Calibration,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=3.0,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test
  • 335267 10/17/10 03:14 a-onl End_of_Run_7442
  • 335266 10/17/10 03:06 a-onl Start_Run_7442,Run_type=Calibration,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=4.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test
  • 335265 10/17/10 02:58 a-onl End_of_Run_7441
  • 335264 10/17/10 02:50 auto_snap Radiation and Ozone snapshot
  • 335263 10/17/10 02:50 auto_snap Miscellaneous snapshot
  • 335262 10/17/10 02:50 auto_snap Hall A Tools snapshot
  • 335261 10/17/10 02:50 auto_snap Compton Polarimeter snapshot
  • 335260 10/17/10 02:50 auto_snap Beam Position and Current snapshot
  • 335259 10/17/10 02:44 a-onl Start_Run_7441,Run_type=Calibration,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=on,beam_current=4.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test
  • 335258 10/17/10 02:35 D.Wang target moved to Carbon
  • 335257 10/17/10 02:31 efuchey Hall A tools snapshot
  • 335256 10/17/10 02:31 a-onl End_of_Run_7440
  • 335255 10/17/10 02:23 a-onl Start_Run_7440,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=4.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test
  • 335254 10/17/10 02:19 D.Wang target moved to Loop2, 15cm
  • 335253 10/17/10 02:14 a-onl End_of_Run_7439
  • 335252 10/17/10 02:07 a-onl Start_Run_7439,Run_type=Calibration,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test
  • 335251 10/17/10 02:07 a-onl End_of_Run_7438
  • 335250 10/17/10 02:05 a-onl Start_Run_7438,Run_type=Calibration,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=target boiling test
  • 335249 10/17/10 01:59 D.Wang target moved to Carbon
  • 335248 10/17/10 01:54 a-onl End_of_Run_7437
  • 335247 10/17/10 01:48 D.Wang target status
  • 335246 10/17/10 01:47 a-onl Start_Run_7437,Run_type=Spot++,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=spot++
  • 335245 10/17/10 01:46 a-onl Start_Run_7436,Run_type=Spot++,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=4.5,leftarm_p=2.0,leftarm_theta=25.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=spot++
  • 335244 10/17/10 01:40 D.Wang target moved to Loop2, 15cm
  • 335243 10/17/10 01:30 efuchey Second harpscan
  • 335242 10/17/10 01:22 D.Wang target moved to empty
  • 335241 10/17/10 01:20 efuchey Dipole crash
  • 335240 10/17/10 01:18 efuchey Harp scan
  • 335239 10/17/10 01:01 D.Wang target moved to BeO
  • 335238 10/17/10 01:00 a-onl End_of_Run_7435
  • 335237 10/17/10 00:57 D.Wang spot++
  • 335236 10/17/10 00:55 a-onl Start_Run_7435,Run_type=Spot++,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=spot++
  • 335235 10/17/10 00:28 a-onl Start_Run_7434,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Check sync signal in ARS
  • 335234 10/17/10 00:15 a-onl Start_Run_7433,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Check sync signal in ARS
  • 335233 10/17/10 00:02 adaq Beam Time Accounting
  • 335232 10/17/10 00:01 a-onl End_of_Run_7431
  • 335231 10/17/10 00:01 a-onl Start_Run_7431,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Check sync signal in ARS
  • Saturday

  • 335230 10/16/10 23:58 a-onl Start_Run_7430,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Check sync signal in ARS
  • 335229 10/16/10 23:53 Ramesh Shift Summary
  • 335228 10/16/10 23:41 a-onl Start_Run_7429,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Check sync signal in ARS
  • 335227 10/16/10 23:20 a-onl Start_Run_7425,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Check sync signal in ARS
  • 335226 10/16/10 23:14 A. Camsonne Added clock in D class db_D.dat
  • 335225 10/16/10 22:56 A. Camsonne Clock on TDC for sync check
  • 335224 10/16/10 22:53 brads Logbook now has a preview option
  • 335223 10/16/10 22:25 Ramesh hall in controlled access
  • 335222 10/16/10 22:01 ZAFAR AHMEED Empty target
  • 335221 10/16/10 21:45 brads Tools and NMR screens for 2.0 GeV/c
  • 335220 10/16/10 21:18 Ramesh Beam down for a while for an IOC reboot
  • 335219 10/16/10 21:14 Ramesh Spectrometer set for boiling study
  • 335218 10/16/10 21:13 Abdurahim L-HRS elastic momentum vs phi (run7422)
  • 335217 10/16/10 21:04 Carlos Calorimeter HV off for boiling studies
  • 335216 10/16/10 21:03 Ramesh Preparing for boiling study
  • 335215 10/16/10 20:58 Abdurahim empity fifo5 for roc15 7 roc17
  • 335214 10/16/10 20:57 a-onl End_of_Run_7424
  • 335213 10/16/10 20:02 Abdurahim scalers for run 7424
  • 335212 10/16/10 19:55 a-onl Start_Run_7424,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335211 10/16/10 19:54 Abdurahim empty histos 7422
  • 335210 10/16/10 19:48 hyde T3 timing HRS S1 S2m
  • 335209 10/16/10 19:40 a-onl Start_Run_7423,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335208 10/16/10 19:39 a-onl End_of_Run_7422
  • 335207 10/16/10 19:38 Abdurahim L-HRS elastic momentum vs phi (run7422)
  • 335206 10/16/10 19:36 Abdurahim scalers window
  • 335205 10/16/10 19:35 Abdurahim xwindows for roc14 and roc17
  • 335204 10/16/10 19:24 Abdurahim L HRS elastic momentum vs phi run7421
  • 335203 10/16/10 18:50 auto_snap Radiation and Ozone snapshot
  • 335202 10/16/10 18:50 auto_snap Miscellaneous snapshot
  • 335201 10/16/10 18:50 auto_snap Hall A Tools snapshot
  • 335200 10/16/10 18:50 auto_snap Compton Polarimeter snapshot
  • 335199 10/16/10 18:50 auto_snap Beam Position and Current snapshot
  • 335198 10/16/10 18:36 a-onl Start_Run_7422,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335197 10/16/10 18:35 a-onl End_of_Run_7421
  • 335196 10/16/10 18:32 Abdurahim screen shot for roc14 & roc17 for run7421
  • 335195 10/16/10 17:48 Abdurahim Scalers
  • 335194 10/16/10 17:34 a-onl Start_Run_7421,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.406,leftarm_theta=37.00,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335193 10/16/10 17:33 A. Camsonne Run 7421 with test with buffering
  • 335192 10/16/10 17:25 Ramesh About to get beam
  • 335191 10/16/10 16:57 brads Tools and NMR screens for 2.406 GeV/c
  • 335190 10/16/10 16:53 A. Camsonne NMR locked at 2.406 GeV/c on probe B
  • 335189 10/16/10 16:53 Ramesh NMR locked
  • 335188 10/16/10 16:45 Ramesh Left NMR locking problem
  • 335187 10/16/10 16:37 adev Start_Run_20021,Run_type=Cosmics,target_type=No Target,beam_e0=0,beam_raster=on,beam_halfwaveplate=in,rightarm_p=0,rightarm_theta=-70,comment_text=cosmics
  • 335186 10/16/10 16:36 adev End_of_Run_20020
  • 335185 10/16/10 16:31 ZAFAR AHMED Target is LH2 Loop2 15cm
  • 335184 10/16/10 16:21 ZAFAR AHMED Empty target
  • 335183 10/16/10 16:14 adaq Beam Time Accounting
  • 335182 10/16/10 16:07 jmammei Day shift run summary
  • 335181 10/16/10 16:05 jmammei set LHRS to 37.000 degrees
  • 335180 10/16/10 15:59 chench HRS elastic protons
  • 335179 10/16/10 15:55 chench Run7420 has same empty Calo ADC and TDC histograms as run7419
  • 335178 10/16/10 15:50 a-onl End_of_Run_7420
  • 335177 10/16/10 15:47 chench screen shot for xwindow of Roc14 and roc17
  • 335176 10/16/10 15:10 chench Left HRS momentum vs phi run7419
  • 335175 10/16/10 15:02 chench Run7418,7419,7420 have same prescale set, same LHRS angle and momentum
  • 335174 10/16/10 15:00 chench Empty histogram for calo ADC and TDC in run7419
  • 335173 10/16/10 14:51 brads re: spurious gas sensor alarm?
  • 335172 10/16/10 14:48 a-onl Start_Run_7420,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335171 10/16/10 14:48 a-onl End_of_Run_7419
  • 335170 10/16/10 14:42 chench Empty fifos roc14 and roc17 at the end run of 7419
  • 335169 10/16/10 14:00 chench Elastic (HRS) p vs phi
  • 335168 10/16/10 13:44 a-onl Start_Run_7419,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335167 10/16/10 13:43 a-onl End_of_Run_7418
  • 335166 10/16/10 13:20 adev Start_Run_20020,Run_type=Cosmics,target_type=No Target,beam_e0=0,beam_raster=on,beam_halfwaveplate=in,rightarm_p=0,rightarm_theta=-70,comment_text=cosmics
  • 335165 10/16/10 13:18 adev End_of_Run_20019
  • 335164 10/16/10 13:10 adev Start_Run_20019,Run_type=Cosmics,target_type=No Target,beam_e0=0,beam_raster=on,beam_halfwaveplate=in,rightarm_p=0,rightarm_theta=-70,comment_text=cosmics
  • 335163 10/16/10 12:49 adev End_of_Run_20018
  • 335162 10/16/10 12:45 chench following figures for the scalers of Cer
  • 335161 10/16/10 12:40 a-onl Start_Run_7418,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335160 10/16/10 12:40 chench run7418 scalers
  • 335159 10/16/10 12:38 a-onl End_of_Run_7417
  • 335158 10/16/10 12:37 pzhu fluctuation
  • 335157 10/16/10 12:22 adev Start_Run_20018,Run_type=Cosmics,target_type=No Target,beam_e0=0,beam_raster=on,beam_halfwaveplate=in,rightarm_p=0,rightarm_theta=-70,comment_text=cosmics
  • 335156 10/16/10 12:11 chench run7417 has the same prescale and configures(leftHrs) with run7416
  • 335155 10/16/10 12:04 a-onl Start_Run_7417,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335154 10/16/10 12:00 a-onl End_of_Run_7416
  • 335153 10/16/10 11:51 chench HRS elastic protons
  • 335152 10/16/10 11:39 chench following figure for scalers
  • 335151 10/16/10 11:36 chench Empty fifos for roc17
  • 335150 10/16/10 11:33 chench scalers for run7416
  • 335149 10/16/10 11:30 a-onl Start_Run_7416,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335148 10/16/10 11:26 jmammei resuming run plan
  • 335147 10/16/10 11:22 pzhu Target moved to Loop2,15cm
  • 335146 10/16/10 11:09 jmammei MCC wants to go to 50uA; empty target
  • 335145 10/16/10 11:04 pzhu target moved to empty
  • 335144 10/16/10 11:04 jmammei MCC sending tune beam
  • 335143 10/16/10 10:50 auto_snap Radiation and Ozone snapshot
  • 335142 10/16/10 10:50 auto_snap Miscellaneous snapshot
  • 335141 10/16/10 10:50 auto_snap Hall A Tools snapshot
  • 335140 10/16/10 10:50 auto_snap Compton Polarimeter snapshot
  • 335139 10/16/10 10:50 auto_snap Beam Position and Current snapshot
  • 335138 10/16/10 10:46 pzhu D2Pressures increase
  • 335137 10/16/10 10:31 jmammei beam update
  • 335136 10/16/10 10:19 pzhu h2pressures increase
  • 335135 10/16/10 10:06 jmammei spurious gas sensor alarm?
  • 335134 10/16/10 09:42 jmammei NMR locked - thanks Brad!
  • 335133 10/16/10 09:28 brads LHRS dipole at 2.607 and NMR locked (HowTo instructions)
  • 335132 10/16/10 09:26 pzhu stable cryoPressures
  • 335131 10/16/10 09:11 pzhu moved target to Loop2,15cm
  • 335130 10/16/10 09:10 chench The number of events shown on coda keeping to flip all the time
  • 335129 10/16/10 08:54 pzhu fluctuation
  • 335128 10/16/10 08:51 jmammei beam down ~15 minutes
  • 335127 10/16/10 08:49 a-onl End_of_Run_7415
  • 335126 10/16/10 08:47 a-onl Start_Run_7415,Run_type=Spot++,target_type=BeO,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335125 10/16/10 08:46 jmammei Q2, 3 momentum increased by 2%
  • 335124 10/16/10 08:45 chench spot++ for run7415
  • 335123 10/16/10 08:37 adaq moved target to BeO
  • 335122 10/16/10 08:33 adaq Beam Time Accounting
  • 335121 10/16/10 08:25 chench The number of events for run7414 is flipping between two numbers
  • 335120 10/16/10 08:17 hyde Owl shift summary
  • 335119 10/16/10 08:07 jmammei NMR won't lock at 2.607 GeV
  • 335118 10/16/10 07:40 a-onl End_of_Run_7414
  • 335117 10/16/10 07:29 hyde HRS motion controls
  • 335116 10/16/10 06:36 hyde calo DC ADC
  • 335115 10/16/10 06:26 a-onl Start_Run_7414,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 335114 10/16/10 06:24 hyde spot++
  • 335113 10/16/10 06:16 a-onl End_of_Run_7413
  • 335112 10/16/10 05:34 D.Wang Target status
  • 335111 10/16/10 05:32 cshloka mpty histograms for calo TDC 129, 147 and 151
  • 335110 10/16/10 05:30 cshloka empty histograms in calo ADC 129, 147 and 151
  • 335109 10/16/10 05:23 a-onl Start_Run_7413,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=0,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=cosmic run
  • 335108 10/16/10 05:18 a-onl End_of_Run_7412
  • 335107 10/16/10 05:14 hyde Accel down
  • 335106 10/16/10 04:52 cshloka run 7409; calo TDC 179-193 and 194-207
  • 335105 10/16/10 04:51 cshloka run 7409; calo TDC 120-143 and 144-167
  • 335104 10/16/10 04:50 cshloka run 7409; calo TDC 72-95 and 96-119
  • 335103 10/16/10 04:49 cshloka run 7409; calo TDC 24-47 and 48-71
  • 335102 10/16/10 04:47 cshloka run 7409; calo ADC 194-207 and TDC 0-23
  • 335101 10/16/10 04:46 cshloka run 7409; calo ADC 144-167 and 179-193
  • 335100 10/16/10 04:44 cshloka run 7409; calo ADC 96-119 and 120-143
  • 335099 10/16/10 04:43 cshloka run 7409; calo adc 48-71 and 72-95
  • 335098 10/16/10 04:42 cshloka calo adc 0-23 and 24-47
  • 335097 10/16/10 04:22 correction
  • 335096 10/16/10 04:16 cshloka run number for empty fifo
  • 335095 10/16/10 04:12 cshloka Empty Fifo
  • 335094 10/16/10 04:12 a-onl Start_Run_7412,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=elastic calibration 400V
  • 335093 10/16/10 04:09 a-onl End_of_Run_7411
  • 335092 10/16/10 04:08 hyde ARS firmware
  • 335091 10/16/10 04:08 a-onl End_of_Run_7410
  • 335090 10/16/10 04:06 A. Camsonne Logic module FPGA loaded with 256 us dead time for valid and 1 us deadtime for stop
  • 335089 10/16/10 03:58 hyde HRS elastic protons
  • 335088 10/16/10 03:27 a-onl Start_Run_7410,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=elastic calibration 400V
  • 335087 10/16/10 03:26 a-onl End_of_Run_7409
  • 335086 10/16/10 03:18 A. Camsonne DVCS crls files modified
  • 335085 10/16/10 03:05 hyde spot++ run 7409
  • 335084 10/16/10 02:57 hyde empty fifo
  • 335083 10/16/10 02:49 auto_snap Radiation and Ozone snapshot
  • 335082 10/16/10 02:49 auto_snap Miscellaneous snapshot
  • 335081 10/16/10 02:49 auto_snap Hall A Tools snapshot
  • 335080 10/16/10 02:49 auto_snap Compton Polarimeter snapshot
  • 335079 10/16/10 02:49 auto_snap Beam Position and Current snapshot
  • 335078 10/16/10 02:43 a-onl Start_Run_7409,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=elastic calibration 400V
  • 335077 10/16/10 02:42 a-onl End_of_Run_7408
  • 335076 10/16/10 02:38 hyde Calo HV on
  • 335075 10/16/10 02:31 hyde cerenkov scalers
  • 335074 10/16/10 02:30 a-onl Start_Run_7408,Run_type=Spot++,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=spot++
  • 335073 10/16/10 02:30 hyde scaler s2
  • 335072 10/16/10 02:29 A. Camsonne Calorimer HV on 400 V
  • 335071 10/16/10 02:29 hyde Scaler S1
  • 335070 10/16/10 02:28 hyde scalers on LH2
  • 335069 10/16/10 02:26 a-onl End_of_Run_7407
  • 335068 10/16/10 02:25 D.Wang target moved to Loo2, 15cm
  • 335067 10/16/10 02:20 a-onl Start_Run_7407,Run_type=Spot++,target_type=BeO,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=spot++
  • 335066 10/16/10 02:16 cshloka moved target to BeO
  • 335065 10/16/10 02:08 hyde bPM
  • 335064 10/16/10 02:04 a-onl End_of_Run_7404
  • 335063 10/16/10 02:00 a-onl Start_Run_7404,Run_type=Spot++,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=spot++
  • 335062 10/16/10 01:58 a-onl End_of_Run_7403
  • 335061 10/16/10 01:55 a-onl End_of_Run_7402
  • 335060 10/16/10 01:39 D.Wang target status
  • 335059 10/16/10 01:17 a-onl Start_Run_7402,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS ROC14 and ROC17 test empty fifo signal - HV on on calo
  • 335058 10/16/10 01:05 a-onl End_of_Run_7400
  • 335057 10/16/10 01:02 A. Camsonne DVCS DAQ work
  • 335056 10/16/10 01:02 a-onl Start_Run_7400,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS ROC14 and ROC17 test empty fifo signal - HV on on calo
  • 335055 10/16/10 01:00 a-onl End_of_Run_7399
  • 335054 10/16/10 00:56 hyde bbscope/ HRS
  • 335053 10/16/10 00:54 hyde S1 S2 TDC
  • 335052 10/16/10 00:51 a-onl Start_Run_7399,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS ROC14 and ROC17 test empty fifo signal - HV on on calo
  • 335051 10/16/10 00:49 a-onl End_of_Run_7398
  • 335050 10/16/10 00:48 a-onl Start_Run_7398,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS ROC14 and ROC17 test empty fifo signal
  • 335049 10/16/10 00:46 hyde 1024 Hz Clk Double Pulsing
  • 335048 10/16/10 00:45 hyde Cerenkov TDC Threshold
  • 335047 10/16/10 00:44 a-onl End_of_Run_7397
  • 335046 10/16/10 00:40 a-onl Start_Run_7397,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS ROC14 and ROC17 test empty fifo signal
  • 335045 10/16/10 00:38 a-onl End_of_Run_7396
  • 335044 10/16/10 00:21 a-onl Start_Run_7396,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS ROC14 and ROC17
  • 335043 10/16/10 00:20 a-onl End_of_Run_7395
  • 335042 10/16/10 00:11 a-onl Start_Run_7395,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS ROC14 and ROC17
  • 335041 10/16/10 00:11 a-onl End_of_Run_7394
  • 335040 10/16/10 00:07 a-onl Start_Run_7394,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS ROC14 and ROC17
  • 335039 10/16/10 00:04 adaq Beam Time Accounting
  • 335038 10/16/10 00:02 a-onl Start_Run_7393,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS ROC14 and ROC17
  • Friday

  • 335037 10/15/10 23:56 BoZhao Swing Shift Summary
  • 335036 10/15/10 23:23 J. Zhang Moller magnet has been filled
  • 335035 10/15/10 23:18 BoZhao Raster is turned off
  • 335034 10/15/10 23:06 O.Glamazdin Moller measurements result
  • 335033 10/15/10 21:50 hyde BCM
  • 335032 10/15/10 21:31 hyde BCM calibration
  • 335031 10/15/10 20:38 O.Glamazdin Moller magnet status
  • 335030 10/15/10 20:32 O.Glamazdin Moller SC magnet quench
  • 335029 10/15/10 20:10 hyde BCM
  • 335028 10/15/10 18:49 auto_snap Radiation and Ozone snapshot
  • 335027 10/15/10 18:49 auto_snap Miscellaneous snapshot
  • 335026 10/15/10 18:49 auto_snap Hall A Tools snapshot
  • 335025 10/15/10 18:49 auto_snap Compton Polarimeter snapshot
  • 335024 10/15/10 18:49 auto_snap Beam Position and Current snapshot
  • 335023 10/15/10 18:46 Abdurahim Compton Cavity QWP Scan Procedures
  • 335022 10/15/10 18:07 Kalyan Follow-up on lumi test
  • 335021 10/15/10 17:59 a-onl End_of_Run_7392
  • 335020 10/15/10 17:54 a-onl Start_Run_7392,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=scaler check
  • 335019 10/15/10 16:48 brads How to Reset the HRS bogie IOC
  • 335018 10/15/10 16:36 efuchey QWP scan for 10/15/2010
  • 335017 10/15/10 16:04 adaq Beam Time Accounting
  • 335016 10/15/10 15:54 X. Zheng Day shift summary
  • 335015 10/15/10 15:48 brads Run Plan for Friday 15 DAY -> Saturday 16 DAY
  • 335014 10/15/10 15:19 Saha Hall A Arc Energy Measurement
  • 335013 10/15/10 15:09 X. Zheng Arc done, now Moller
  • 335012 10/15/10 14:10 efuchey change of Compton threshold
  • 335011 10/15/10 14:10 efuchey change of Compton threshold
  • 335010 10/15/10 14:09 A. Camsonne scope in LHRS plugged on network bbscope1
  • 335009 10/15/10 13:21 X. Zheng Arc energy measurement
  • 335008 10/15/10 12:42 brads DVCS Calorimeter HV off
  • 335007 10/15/10 12:07 a-onl Start_Run_7391,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Cosmics new trigger module firmware
  • 335006 10/15/10 11:50 A. Camsonne DVCS calorimeter HV ON for cosmics
  • 335005 10/15/10 11:46 A. Camsonne Logic module FPGA loaded with 256 us dead time for stop
  • 335004 10/15/10 11:14 ellie Run 7388, CaloAmp6 ADC looks wonky. Lots of TDCs have low signal:noise
  • 335003 10/15/10 11:08 ellie Pion Rejector analysis plots blank
  • 335002 10/15/10 11:01 A. Camsonne BCM log file
  • 335001 10/15/10 10:49 auto_snap Radiation and Ozone snapshot
  • 335000 10/15/10 10:49 auto_snap Miscellaneous snapshot
  • 334999 10/15/10 10:49 auto_snap Hall A Tools snapshot
  • 334998 10/15/10 10:49 auto_snap Compton Polarimeter snapshot
  • 334997 10/15/10 10:49 auto_snap Beam Position and Current snapshot
  • 334996 10/15/10 10:44 O.Glamazdin Moller magnet has been filled
  • 334995 10/15/10 10:33 kelleher Target motion -- empty
  • 334994 10/15/10 10:07 a-onl End_of_Run_7388
  • 334993 10/15/10 10:02 adaq Beam Time Accounting
  • 334992 10/15/10 09:32 X. Zheng Current plan
  • 334991 10/15/10 09:16 a-onl Start_Run_7388,Run_type=Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.784,leftarm_theta=33.29,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=
  • 334990 10/15/10 09:13 X. Zheng Left HRS moved to 33.290 degrees
  • 334989 10/15/10 09:04 R. Michaels re: helicity in scalers
  • 334988 10/15/10 08:53 jroche new Helicity board firmware installed
  • 334987 10/15/10 08:52 a-onl End_of_Run_7387
  • 334986 10/15/10 08:46 a-onl Start_Run_7387,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1 uA,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Helicity test run (ps9=1, all others at 0)
  • 334985 10/15/10 08:04 jroche Owl shift summary
  • 334984 10/15/10 07:32 a-onl End_of_Run_7386
  • 334983 10/15/10 06:57 a-onl Start_Run_7386,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1 uA,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test run 3.0
  • 334982 10/15/10 06:55 florian itar calo TDC 194-207
  • 334981 10/15/10 06:54 florian itar calo TDC 144-167 and 179-193
  • 334980 10/15/10 06:52 florian itar calo TDC 96-119 and 120-143
  • 334979 10/15/10 06:51 florian itar calo TDC 48-71 and 72-95
  • 334978 10/15/10 06:50 florian itar calo TDC 0-23 and 24-47
  • 334977 10/15/10 06:49 florian itar calo ADC 194-207
  • 334976 10/15/10 06:48 florian itar calo ADC 144-167 and 179-193
  • 334975 10/15/10 06:47 florian itar calo ADC 96-119 and 120-143
  • 334974 10/15/10 06:45 florian itar calo ADC 48-71 and 72-95
  • 334973 10/15/10 06:44 florian itar calo ADC 0-23 et 24-47
  • 334972 10/15/10 06:23 a-onl End_of_Run_7385
  • 334971 10/15/10 05:49 florian correction
  • 334970 10/15/10 05:30 a-onl Start_Run_7385,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1 uA,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test run 3.0
  • 334969 10/15/10 05:28 a-onl End_of_Run_7384
  • 334968 10/15/10 05:06 florian itar current information
  • 334967 10/15/10 05:03 florian counting rates current information
  • 334966 10/15/10 05:00 yez target screen - cryo
  • 334965 10/15/10 04:59 yez target screen -- temp & press.
  • 334964 10/15/10 04:55 florian itar counting rates elastic calibration calorimeter
  • 334963 10/15/10 04:54 florian itar counting rates elastic calibration calorimeter
  • 334962 10/15/10 04:36 jroche Left bogie being really difficult (TECHS)
  • 334961 10/15/10 04:27 a-onl Start_Run_7384,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1 uA,leftarm_p=2.607,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test run 3.0
  • 334960 10/15/10 03:41 jroche LHRS set for 2.607 GeV
  • 334959 10/15/10 03:20 jroche shirt access to check on ladders
  • 334958 10/15/10 03:06 a-onl End_of_Run_7383
  • 334957 10/15/10 02:49 auto_snap Radiation and Ozone snapshot
  • 334956 10/15/10 02:49 auto_snap Miscellaneous snapshot
  • 334955 10/15/10 02:48 auto_snap Hall A Tools snapshot
  • 334954 10/15/10 02:48 auto_snap Compton Polarimeter snapshot
  • 334953 10/15/10 02:48 auto_snap Beam Position and Current snapshot
  • 334952 10/15/10 02:48 yez Sieve Partten Analysis -- See holes for LH2 but not for Carbon foils
  • 334951 10/15/10 02:45 a-onl Start_Run_7383,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1 uA,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test run 3.0
  • 334950 10/15/10 02:24 a-onl End_of_Run_7382
  • 334949 10/15/10 02:13 florian itar counting rates elastic calibration calorimeter
  • 334948 10/15/10 02:12 florian itar counting rates elastic calibration calorimeter
  • 334947 10/15/10 01:55 a-onl Start_Run_7382,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1 uA,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test run 3.0
  • 334946 10/15/10 01:54 a-onl End_of_Run_7381
  • 334945 10/15/10 00:48 a-onl Start_Run_7381,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=1 uA,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test run 3.0
  • 334944 10/15/10 00:45 a-onl End_of_Run_7380
  • 334943 10/15/10 00:43 a-onl End_of_Run_7379
  • 334942 10/15/10 00:18 a-onl Start_Run_7379,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1 uA,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test run 4.2
  • 334941 10/15/10 00:13 a-onl End_of_Run_7378
  • 334939 10/15/10 00:00 Solvignon Swing shift summary
  • 334940 10/15/10 00:00 adaq Beam Time Accounting
  • Thursday

  • 334938 10/14/10 23:57 a-onl Start_Run_7378,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1 uA,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test run 3uA
  • 334937 10/14/10 23:56 a-onl End_of_Run_7377
  • 334936 10/14/10 23:53 A. Camsonne Calorimeter HV on - Run 7377 1 uA
  • 334935 10/14/10 23:40 a-onl Start_Run_7377,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1 uA,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test run 1uA
  • 334934 10/14/10 23:36 efuchey Compton rates became dramatic
  • 334933 10/14/10 23:31 a-onl Start_Run_7376,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1 uA,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Pulser run
  • 334932 10/14/10 23:22 a-onl Start_Run_7374,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1 uA,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Pulser run
  • 334931 10/14/10 23:21 a-onl End_of_Run_7373
  • 334930 10/14/10 23:21 a-onl Start_Run_7373,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=1 uA,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Pulser run
  • 334929 10/14/10 23:19 a-onl End_of_Run_7372
  • 334928 10/14/10 23:17 a-onl Start_Run_7372,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=3 and up,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Pulser run
  • 334927 10/14/10 22:43 a-onl Start_Run_7371,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=3 and up,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration LH2
  • 334926 10/14/10 22:42 A. Camsonne Calorimeter HV on - Run 7371
  • 334925 10/14/10 22:42 a-onl End_of_Run_7369
  • 334924 10/14/10 22:35 a-onl End_of_Run_7369
  • 334923 10/14/10 22:35 a-onl End_of_Run_7362
  • 334922 10/14/10 22:32 R. Michaels expected rates = measured rate, I think.
  • 334921 10/14/10 21:51 R. Michaels lack of holes for T3
  • 334920 10/14/10 21:38 R. Michaels LH2 rates, run 7369
  • 334919 10/14/10 21:36 R. Michaels study to analyze trigger efficiency
  • 334918 10/14/10 21:22 Solvignon Spot++
  • 334917 10/14/10 21:21 a-onl Start_Run_7369,Run_type=Check out,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=3 and up,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=trigger efficiency study
  • 334916 10/14/10 21:13 mcanan Target Motion
  • 334915 10/14/10 21:11 a-onl End_of_Run_7368
  • 334914 10/14/10 20:50 a-onl Start_Run_7368,Run_type=Check out,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=off,beam_current=20 (and others),leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=trigger efficiency study
  • 334913 10/14/10 20:49 a-onl End_of_Run_7367
  • 334912 10/14/10 20:43 a-onl Start_Run_7367,Run_type=Check out,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=off,beam_current=20 (and others),leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=trigger efficiency study
  • 334911 10/14/10 20:36 hyde s0 tdc in raw data
  • 334910 10/14/10 20:34 mcanan Target Motion
  • 334909 10/14/10 20:30 efuchey Cavity scan
  • 334908 10/14/10 19:54 mcanan Target Motion
  • 334907 10/14/10 19:49 a-onl End_of_Run_7366
  • 334906 10/14/10 19:45 Kalyan Summary of lumi test runs
  • 334905 10/14/10 19:43 a-onl Start_Run_7366,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=10,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Lumi test B with 1uA
  • 334904 10/14/10 19:42 a-onl End_of_Run_7365
  • 334903 10/14/10 19:35 a-onl Start_Run_7365,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=10,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Lumi test B with 2uA
  • 334902 10/14/10 19:34 a-onl End_of_Run_7364
  • 334901 10/14/10 19:28 a-onl Start_Run_7364,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=10,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Lumi test B with 3uA
  • 334900 10/14/10 19:26 a-onl End_of_Run_7363
  • 334899 10/14/10 19:24 hyde S1 S2 trigger timing in TDC
  • 334898 10/14/10 19:23 Kalyan Run numbers of Lumi test A
  • 334897 10/14/10 19:20 a-onl Start_Run_7363,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Lumi test B with 4uA
  • 334896 10/14/10 19:15 a-onl End_of_Run_7357
  • 334895 10/14/10 19:11 a-onl Start_Run_7362,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Lumi test B with 5uA
  • 334894 10/14/10 19:10 a-onl End_of_Run_7361
  • 334893 10/14/10 19:04 efuchey Update for Compton rates
  • 334892 10/14/10 19:04 a-onl Start_Run_7361,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Lumi test B with 10uA
  • 334891 10/14/10 19:03 Kalyan HV setting for lumi test B
  • 334890 10/14/10 18:57 a-onl End_of_Run_7360
  • 334889 10/14/10 18:55 efuchey lost readout of compton:rate_G
  • 334888 10/14/10 18:50 a-onl Start_Run_7360,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=10,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Lumi test A with 1uA
  • 334887 10/14/10 18:49 a-onl End_of_Run_7359
  • 334886 10/14/10 18:48 auto_snap Radiation and Ozone snapshot
  • 334885 10/14/10 18:48 auto_snap Miscellaneous snapshot
  • 334884 10/14/10 18:48 auto_snap Hall A Tools snapshot
  • 334883 10/14/10 18:48 auto_snap Compton Polarimeter snapshot
  • 334882 10/14/10 18:48 auto_snap Beam Position and Current snapshot
  • 334881 10/14/10 18:43 a-onl Start_Run_7359,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=10,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Lumi test A with 2uA
  • 334880 10/14/10 18:42 a-onl End_of_Run_7358
  • 334879 10/14/10 18:35 a-onl Start_Run_7358,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=10,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Lumi test A with 3uA
  • 334878 10/14/10 18:26 a-onl Start_Run_7357,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Lumi test A with 4uA
  • 334877 10/14/10 18:25 a-onl End_of_Run_7356
  • 334876 10/14/10 18:19 a-onl Start_Run_7356,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Lumi test A with 5uA
  • 334875 10/14/10 18:18 a-onl End_of_Run_7355
  • 334874 10/14/10 18:15 efuchey Compton rates
  • 334873 10/14/10 18:12 a-onl Start_Run_7355,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Lumi test A with 10uA
  • 334872 10/14/10 18:11 Kalyan HV setting for Lumi test A
  • 334871 10/14/10 18:05 brads Injector matching/tuning was completed before the BCM calibration started
  • 334870 10/14/10 18:01 brads Notes on the BCM calibration procedure
  • 334869 10/14/10 17:56 mcanan Target Motion
  • 334868 10/14/10 17:43 brads re: Left HRS NMR -- multiple probes available
  • 334867 10/14/10 17:36 hyde EDTM in S2m TDC
  • 334866 10/14/10 17:16 a-onl End_of_Run_7354
  • 334865 10/14/10 17:09 gomez Left HRS NMR -- multiple probes available
  • 334864 10/14/10 17:04 a-onl Start_Run_7354,Run_type=Cosmics,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=BCM calibration (2nd try)
  • 334863 10/14/10 17:03 a-onl End_of_Run_7353
  • 334862 10/14/10 16:46 Solvignon BCM calibration: 1st part
  • 334861 10/14/10 16:45 a-onl Start_Run_7353,Run_type=Cosmics,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=BCM calibration (2nd try)
  • 334860 10/14/10 16:23 a-onl End_of_Run_7352
  • 334859 10/14/10 16:09 a-onl Start_Run_7352,Run_type=Cosmics,target_type=Empty loop,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=BCM calibration
  • 334858 10/14/10 16:06 a-onl End_of_Run_7351
  • 334857 10/14/10 16:03 adaq Beam Time Accounting
  • 334856 10/14/10 15:49 R. Michaels network work finished, MSS copying resumed
  • 334855 10/14/10 15:36 Carlos Left sieve is OUT
  • 334854 10/14/10 15:25 Jack Magnets use LCW - They will need to be ramped down and turned off
  • 334853 10/14/10 14:38 brads LHRS NMR tricks and tips
  • 334852 10/14/10 14:31 a-onl Start_Run_7351,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=HRS cosmics
  • 334851 10/14/10 14:18 Magali Which ARS board in which crate
  • 334850 10/14/10 14:05 J. Zhang Moller magnet has been filled.
  • 334849 10/14/10 13:57 folts/tech sieve position
  • 334848 10/14/10 13:31 brads Update on LHRS sieve and NMR status
  • 334847 10/14/10 13:18 a-onl End_of_Run_7350
  • 334846 10/14/10 13:10 a-onl Start_Run_7350,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 14 17 test
  • 334845 10/14/10 13:09 a-onl End_of_Run_7349
  • 334844 10/14/10 13:01 a-onl Start_Run_7349,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 14 17 test
  • 334843 10/14/10 13:00 a-onl End_of_Run_7348
  • 334842 10/14/10 12:53 a-onl Start_Run_7348,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 14 17 test
  • 334841 10/14/10 12:48 a-onl Start_Run_7347,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 14 17 test
  • 334840 10/14/10 12:40 a-onl Start_Run_7346,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 14 17 test
  • 334839 10/14/10 12:37 brads Hall will be down on Monday, 0700 for LCW work
  • 334838 10/14/10 12:31 a-onl Start_Run_7345,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 14 17 test
  • 334837 10/14/10 12:17 a-onl Start_Run_7344,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 14 17 test
  • 334836 10/14/10 12:16 a-onl End_of_Run_7337
  • 334835 10/14/10 12:06 a-onl End_of_Run_7343
  • 334834 10/14/10 12:05 a-onl Start_Run_7343,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 17 test
  • 334833 10/14/10 11:56 a-onl End_of_Run_7342
  • 334832 10/14/10 11:56 hyde HRS-DVCS trigger timing
  • 334831 10/14/10 11:55 a-onl Start_Run_7342,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 17 test
  • 334830 10/14/10 11:51 brads Run Plan for Thurs 14 DAY -> Friday 15 DAY
  • 334829 10/14/10 11:50 a-onl End_of_Run_7341
  • 334828 10/14/10 11:45 a-onl Start_Run_7341,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 17 test
  • 334827 10/14/10 11:43 a-onl End_of_Run_7340
  • 334826 10/14/10 11:42 a-onl Start_Run_7340,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 14 test
  • 334825 10/14/10 11:37 a-onl End_of_Run_7339
  • 334824 10/14/10 11:35 a-onl Start_Run_7339,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 14 test
  • 334823 10/14/10 11:28 a-onl End_of_Run_7338
  • 334822 10/14/10 11:28 a-onl Start_Run_7338,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 14 test
  • 334821 10/14/10 11:26 a-onl Start_Run_7337,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 14 test
  • 334820 10/14/10 11:06 a-onl End_of_Run_7334
  • 334819 10/14/10 10:55 a-onl Start_Run_7334,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=roc 14 test
  • 334818 10/14/10 10:48 auto_snap Radiation and Ozone snapshot
  • 334817 10/14/10 10:48 auto_snap Miscellaneous snapshot
  • 334816 10/14/10 10:48 auto_snap Hall A Tools snapshot
  • 334815 10/14/10 10:48 auto_snap Compton Polarimeter snapshot
  • 334814 10/14/10 10:48 auto_snap Beam Position and Current snapshot
  • 334813 10/14/10 10:16 efuchey 1mm thick Lead filter put back
  • 334812 10/14/10 10:15 efuchey 1mm thick Lead filter put back
  • 334811 10/14/10 09:49 a-onl End_of_Run_7333
  • 334810 10/14/10 09:45 R. Michaels network switchover in progress
  • 334809 10/14/10 09:35 a-onl Start_Run_7333,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=no beam test
  • 334808 10/14/10 09:03 adaq Beam Time Accounting
  • 334807 10/14/10 09:02 jroche owl shift summary
  • 334806 10/14/10 08:52 a-onl Start_Run_7332,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334805 10/14/10 08:41 a-onl Start_Run_7331,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334804 10/14/10 08:05 jroche Scinitllator timing: things look better
  • 334803 10/14/10 07:52 a-onl Start_Run_7330,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334802 10/14/10 07:33 a-onl Start_Run_7329,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334801 10/14/10 07:33 yez LH2 Loop temperature was a little big unstable but fine now
  • 334800 10/14/10 07:07 a-onl Start_Run_7328,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334799 10/14/10 06:49 a-onl Start_Run_7327,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334798 10/14/10 06:34 jroche Cerenkov calibration constant: new values
  • 334797 10/14/10 06:33 jroche Cerenkov calibration constant: new values
  • 334796 10/14/10 06:33 a-onl Start_Run_7326,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334795 10/14/10 06:31 a-onl End_of_Run_7325
  • 334794 10/14/10 06:11 a-onl Start_Run_7325,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=trigger timing test (ps1=25 only trigger)
  • 334793 10/14/10 06:10 a-onl End_of_Run_7324
  • 334792 10/14/10 06:08 a-onl Start_Run_7324,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=trigger timing test (ps1 only the trigger)
  • 334791 10/14/10 05:38 adaq scaler data
  • 334790 10/14/10 05:37 a-onl Start_Run_7323,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334789 10/14/10 05:26 jroche Compton: hall C to 10 uA makes rates A-ok for our Comptom
  • 334788 10/14/10 04:58 a-onl Start_Run_7322,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.406,leftarm_theta=37.0,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334787 10/14/10 04:51 jroche LHRS to second elastic setting for calorimeter calibration.
  • 334786 10/14/10 04:50 jroche Trouble moving LHRS from angle 34.76 to 37.00 (TECHS)
  • 334785 10/14/10 04:48 yez target screen -- cryo
  • 334784 10/14/10 04:47 yez target screen -- temp and press
  • 334783 10/14/10 04:28 jroche Trouble getting the LHRS Dipole NMR to lock (TECHS)
  • 334782 10/14/10 03:50 adaq Beam Time Accounting
  • 334781 10/14/10 03:07 a-onl Start_Run_7321,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334780 10/14/10 02:48 auto_snap Radiation and Ozone snapshot
  • 334779 10/14/10 02:48 auto_snap Miscellaneous snapshot
  • 334778 10/14/10 02:48 auto_snap Hall A Tools snapshot
  • 334777 10/14/10 02:48 auto_snap Compton Polarimeter snapshot
  • 334776 10/14/10 02:47 auto_snap Beam Position and Current snapshot
  • 334775 10/14/10 02:33 a-onl Start_Run_7320,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334774 10/14/10 02:17 jroche Compton rates got higher: aborted Compton data taking
  • 334773 10/14/10 02:09 a-onl Start_Run_7319,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334772 10/14/10 01:52 a-onl Start_Run_7318,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334771 10/14/10 01:38 a-onl Start_Run_7317,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334770 10/14/10 01:19 a-onl Start_Run_7316,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334769 10/14/10 01:09 jroche left LHRS angle display in General tools.
  • 334768 10/14/10 01:02 a-onl Start_Run_7315,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334767 10/14/10 00:47 a-onl Start_Run_7314,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334766 10/14/10 00:46 a-onl Start_Run_7313,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334765 10/14/10 00:44 a-onl End_of_Run_7312
  • 334764 10/14/10 00:40 a-onl Start_Run_7312,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14 ROC18
  • 334763 10/14/10 00:31 a-onl Start_Run_7311,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC14
  • 334762 10/14/10 00:06 a-onl Start_Run_7310,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC18 ROC14
  • Wednesday

  • 334761 10/13/10 23:40 jroche LHRS angle in Hall A general Tool is way off
  • 334760 10/13/10 23:23 a-onl Start_Run_7309,Run_type=Elastic Calibration,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=500 V ROC17 ROC18 ROC14
  • 334759 10/13/10 23:17 a-onl Start_Run_7304,Run_type=Pedestal,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Pedestal run HV off - after one run
  • 334758 10/13/10 23:04 a-onl Start_Run_7308,Run_type=Elastic Calibration,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=400 V ROC17 and ROC14 only
  • 334757 10/13/10 23:03 ZAFAR AHMED Shift summary Wednesday October 13
  • 334756 10/13/10 23:03 a-onl End_of_Run_7307
  • 334755 10/13/10 23:03 a-onl Start_Run_7307,Run_type=Elastic Calibration,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=400 V ROC17 and ROC14 only
  • 334754 10/13/10 23:01 a-onl End_of_Run_7306
  • 334753 10/13/10 22:42 a-onl Start_Run_7306,Run_type=Elastic Calibration,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration - HV=400 V
  • 334752 10/13/10 22:40 a-onl End_of_Run_7305
  • 334751 10/13/10 22:40 a-onl Start_Run_7305,Run_type=Elastic Calibration,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Pedestal run HV off - after one run
  • 334750 10/13/10 22:38 a-onl End_of_Run_7304
  • 334749 10/13/10 22:37 a-onl End_of_Run_7303
  • 334748 10/13/10 22:34 a-onl Start_Run_7303,Run_type=Pedestal,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Pedestal run HV off
  • 334747 10/13/10 22:33 A. Camsonne Low rate ARS pedestal run 7303
  • 334746 10/13/10 22:25 A. Camsonne coda EB seg faulting at end of run
  • 334745 10/13/10 22:23 A. Camsonne coda EB seg faulting at end of run
  • 334744 10/13/10 22:22 a-onl End_of_Run_7300
  • 334743 10/13/10 22:22 a-onl Start_Run_7300,Run_type=Pedestal,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Pedestal run HV off - Junk
  • 334742 10/13/10 22:21 a-onl End_of_Run_7299
  • 334741 10/13/10 22:07 a-onl Start_Run_7298,Run_type=Elastic Calibration,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Elastic calibration ROC 14 17 18 HV = 400 V
  • 334740 10/13/10 22:01 a-onl Start_Run_7296,Run_type=Pedestal,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Pedestal run ROC 14 17 18
  • 334739 10/13/10 21:59 A. Camsonne DVCS ARS Calorimeter pedestal run 7296
  • 334738 10/13/10 21:49 a-onl Start_Run_7295,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=3 crates triggered by coincidence S2m and S0 - HV On 600 V
  • 334737 10/13/10 21:38 a-onl Start_Run_7294,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=3 crates triggered by coincidence S2m and S0 - HV On 600 V
  • 334736 10/13/10 21:32 mcanan loop 2 temperature (green)
  • 334735 10/13/10 21:23 a-onl Start_Run_7293,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=3 crates triggered by coincidence S2m and S0 - HV On 600 V
  • 334734 10/13/10 21:12 a-onl Start_Run_7292,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=3 crates triggered by coincidence S2m and S0 - HV On 600 V
  • 334733 10/13/10 20:49 A. Camsonne Access for DVCS
  • 334732 10/13/10 20:35 mcanan Loop 2 temperature change hal2_Cernox_2_R
  • 334731 10/13/10 20:31 M Friend Compton rates
  • 334730 10/13/10 20:29 a-onl Start_Run_7291,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=3 crates triggered by coincidence S2m and S0 - HV On 600 V
  • 334729 10/13/10 20:07 a-onl Start_Run_7290,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=3 crates triggered by coincidence S2m and S0 - HV On 600 V
  • 334728 10/13/10 20:06 Hisham HRS Singles
  • 334727 10/13/10 20:06 a-onl End_of_Run_7289
  • 334726 10/13/10 20:05 a-onl End_of_Run_7286
  • 334725 10/13/10 20:03 a-onl End_of_Run_7288
  • 334724 10/13/10 20:00 ZAFAR AHMED Beam permit
  • 334723 10/13/10 19:39 M Friend Compton photon beamline
  • 334722 10/13/10 19:14 ZAFAR AHMED Controlled access
  • 334721 10/13/10 18:57 ZAFAR AHMED Sive is out
  • 334720 10/13/10 18:51 a-onl Start_Run_7288,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=2.606,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS triggered by coincidence S2m and S0 - HV On 600 V
  • 334719 10/13/10 18:49 ZAFAR AHMED Beam current is 5uA
  • 334718 10/13/10 18:47 auto_snap Radiation and Ozone snapshot
  • 334717 10/13/10 18:47 auto_snap Miscellaneous snapshot
  • 334716 10/13/10 18:47 auto_snap Hall A Tools snapshot
  • 334715 10/13/10 18:47 auto_snap Compton Polarimeter snapshot
  • 334714 10/13/10 18:47 auto_snap Beam Position and Current snapshot
  • 334713 10/13/10 18:47 a-onl End_of_Run_7287
  • 334712 10/13/10 18:41 a-onl Start_Run_7287,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS triggered by coincidence S2m and S0 - HV On 600 V
  • 334711 10/13/10 18:38 brads Run Plan for Wed 13 SWING -> Thurs 14 DAY
  • 334710 10/13/10 18:26 a-onl Start_Run_7286,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ARS triggered by coincidence S2m and S0
  • 334709 10/13/10 18:19 ZAFAR AHMED Hall A tools
  • 334708 10/13/10 18:16 ZAFAR AHMED Target is LH2
  • 334707 10/13/10 18:16 mcanan Target Motion
  • 334706 10/13/10 18:07 mcanan Target Motion
  • 334705 10/13/10 18:04 ZAFAR AHMED Spot++
  • 334704 10/13/10 18:04 a-onl End_of_Run_7279
  • 334703 10/13/10 18:03 ZAFAR AHMED Beam is back.
  • 334702 10/13/10 17:56 a-onl End_of_Run_7278
  • 334701 10/13/10 17:45 a-onl Start_Run_7278,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ROC14 and ROC17
  • 334700 10/13/10 17:43 a-onl End_of_Run_7277
  • 334699 10/13/10 17:40 a-onl Start_Run_7277,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ROC14 and ROC17
  • 334698 10/13/10 17:39 a-onl End_of_Run_7276
  • 334697 10/13/10 17:34 a-onl Start_Run_7276,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ROC14 and ROC17
  • 334696 10/13/10 17:32 a-onl Start_Run_7274,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ROC14 and ROC17
  • 334695 10/13/10 17:32 a-onl Start_Run_7275,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ROC14 and ROC17
  • 334694 10/13/10 17:30 a-onl End_of_Run_7274
  • 334693 10/13/10 17:30 a-onl End_of_Run_7272
  • 334692 10/13/10 17:27 a-onl End_of_Run_7273
  • 334691 10/13/10 17:08 ZAFAR AHMED Beam Update
  • 334690 10/13/10 17:03 a-onl Start_Run_7273,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ROC14 and ROC17
  • 334689 10/13/10 16:45 ZAFAR AHMED Beam update
  • 334688 10/13/10 16:38 a-onl Start_Run_7272,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ROC14 and ROC17
  • 334687 10/13/10 16:34 brads Target GUI Labels are now correct
  • 334686 10/13/10 16:26 a-onl Start_Run_7268,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=ROC18
  • 334685 10/13/10 16:19 mcanan Target Screens
  • 334684 10/13/10 16:19 ZAFAR AHMED Beam permit
  • 334682 10/13/10 16:03 saha Shift Summary
  • 334683 10/13/10 16:03 adaq Beam Time Accounting
  • 334681 10/13/10 15:49 a-onl Start_Run_7267,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=ARS 3 ROCs
  • 334680 10/13/10 15:46 hyde Changes to trigger retiming
  • 334679 10/13/10 15:44 A. Camsonne pulser in ARS 0x8
  • 334678 10/13/10 15:30 A. Camsonne Left HRS set to 2.606 GeV/c for elastic
  • 334677 10/13/10 14:45 a-onl End_of_Run_7265
  • 334676 10/13/10 14:43 a-onl Start_Run_7265,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=ARS test roc 18 without trigger board
  • 334675 10/13/10 14:22 a-onl Start_Run_7263,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=HRS (Charles)
  • 334674 10/13/10 14:18 a-onl End_of_Run_7262
  • 334673 10/13/10 14:15 a-onl Start_Run_7262,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ROC 17 (DVCSBuff) (Carlos)
  • 334672 10/13/10 14:14 a-onl End_of_Run_7261
  • 334671 10/13/10 14:12 J. Zhang Moller has been refilled
  • 334670 10/13/10 14:12 J. Zhang Moller has been refilled
  • 334669 10/13/10 14:11 a-onl Start_Run_7261,Run_type=Other,target_type=No Target, Empty,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=34.76,leftarm_sieve=out,calo_theta=22.6,calo_distance=5.5,comment_text=Test ROC 18 (Carlos)
  • 334668 10/13/10 14:06 A. Camsonne iocha14 rebooted by itself ?
  • 334667 10/13/10 14:02 a-onl Start_Run_7259,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Retiming is 40 ns after L1A
  • 334666 10/13/10 14:01 a-onl End_of_Run_7258
  • 334665 10/13/10 14:00 a-onl Start_Run_7258,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text= Retiming is 40 ns after L1A
  • 334664 10/13/10 13:58 a-onl End_of_Run_7257
  • 334663 10/13/10 13:54 hsmith Left Dipole polarity change
  • 334662 10/13/10 13:45 R. Michaels problems with trigger downloading (and remedy)
  • 334661 10/13/10 13:42 a-onl Start_Run_7257,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=removed S1 from S1.or.S2 for Retiming ps3=1 checking retiming
  • 334660 10/13/10 13:38 a-onl Start_Run_7256,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=removed S1 from S1.or.S2 for Retiming ps3=1
  • 334659 10/13/10 13:38 a-onl End_of_Run_7255
  • 334658 10/13/10 13:36 a-onl Start_Run_7255,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=removed S1 from S1.or.S2 for Retiming ps3=1
  • 334657 10/13/10 13:36 Carlos HRS & calorimeter at elastic positions
  • 334656 10/13/10 13:35 a-onl End_of_Run_7254
  • 334655 10/13/10 13:28 a-onl Start_Run_7254,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=delayed S1 by 37ns in S1orS2 for Retiming ps3=1
  • 334654 10/13/10 13:28 a-onl End_of_Run_7253
  • 334653 10/13/10 13:25 a-onl Start_Run_7253,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=delayed S1 by 37ns in S1orS2 for Retiming ps3=1
  • 334652 10/13/10 13:23 a-onl End_of_Run_7252
  • 334651 10/13/10 13:12 a-onl End_of_Run_7252
  • 334650 10/13/10 13:01 a-onl Start_Run_7252,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test of RT circuit in hut ps2=1 S0
  • 334649 10/13/10 12:59 kelleher loop 2 temperature (sensor?)
  • 334648 10/13/10 12:59 a-onl End_of_Run_7251
  • 334647 10/13/10 12:58 a-onl Start_Run_7251,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test of RT circuit in hut ps1=1
  • 334646 10/13/10 12:57 a-onl End_of_Run_7250
  • 334645 10/13/10 12:52 a-onl Start_Run_7250,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test of RT circuit in hut
  • 334644 10/13/10 12:43 efuchey Calorimeter RG 213 cables tested
  • 334643 10/13/10 12:01 atrig Trigger-Download
  • 334642 10/13/10 11:58 atrig Trigger-Download
  • 334641 10/13/10 11:55 atrig Trigger-Download
  • 334640 10/13/10 11:51 atrig Trigger-Download
  • 334639 10/13/10 11:50 atrig Trigger-Download
  • 334638 10/13/10 11:49 atrig Trigger-Download
  • 334637 10/13/10 11:47 atrig Trigger-Download
  • 334636 10/13/10 11:46 atrig Trigger-Download
  • 334635 10/13/10 11:42 atrig Trigger-Download
  • 334634 10/13/10 11:36 kelleher Hall A status
  • 334633 10/13/10 11:20 atrig Trigger-Download
  • 334632 10/13/10 11:08 atrig Trigger-Download
  • 334631 10/13/10 11:05 a-onl Start_Run_7249,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ROC 18 - Firmware of October 12 for Trigger module and control
  • 334630 10/13/10 10:50 a-onl Start_Run_7248,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ROC 18
  • 334629 10/13/10 10:49 a-onl End_of_Run_7247
  • 334628 10/13/10 10:48 a-onl Start_Run_7247,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ROC 18
  • 334627 10/13/10 10:47 auto_snap Radiation and Ozone snapshot
  • 334626 10/13/10 10:47 auto_snap Miscellaneous snapshot
  • 334625 10/13/10 10:47 auto_snap Hall A Tools snapshot
  • 334624 10/13/10 10:47 auto_snap Compton Polarimeter snapshot
  • 334623 10/13/10 10:47 auto_snap Beam Position and Current snapshot
  • 334622 10/13/10 10:45 kelleher target motion -- empty
  • 334621 10/13/10 10:36 a-onl End_of_Run_7246
  • 334620 10/13/10 10:34 hyde S1 HV
  • 334619 10/13/10 10:31 hyde S2m timing
  • 334618 10/13/10 10:25 a-onl Start_Run_7246,Run_type=Optics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Retiming check S1 HV off T4 only
  • 334617 10/13/10 10:25 a-onl End_of_Run_7245
  • 334616 10/13/10 10:24 a-onl Start_Run_7245,Run_type=Optics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Retiming check S1 HV off T4/5 only
  • 334615 10/13/10 10:23 a-onl End_of_Run_7244
  • 334614 10/13/10 10:20 a-onl Start_Run_7244,Run_type=Optics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Retiming check S1 HV off T2/5 only
  • 334613 10/13/10 10:19 a-onl End_of_Run_7243
  • 334612 10/13/10 10:15 a-onl Start_Run_7243,Run_type=Optics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Retiming check S1 HV off T1/10 only
  • 334611 10/13/10 10:12 a-onl End_of_Run_7242
  • 334610 10/13/10 10:09 a-onl Start_Run_7242,Run_type=Optics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2.5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Retiming check S1 HV off T1 only
  • 334609 10/13/10 10:01 kelleher target motion -- LH2 15cm
  • 334608 10/13/10 09:57 Saha Arc Scanner Test
  • 334607 10/13/10 09:54 a-onl End_of_Run_7241
  • 334606 10/13/10 09:46 a-onl Start_Run_7241,Run_type=Optics,target_type=C12 1mm foil,beam_e0=5.5,beam_raster=off,beam_current=5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=harp scanning during this run data taking
  • 334605 10/13/10 09:29 kelleher target motion -- optics target
  • 334604 10/13/10 09:28 a-onl End_of_Run_7240
  • 334603 10/13/10 08:54 A. Camsonne Wrong order in FPGA chain
  • 334602 10/13/10 08:24 a-onl Start_Run_7240,Run_type=Optics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=50,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=
  • 334601 10/13/10 08:22 a-onl End_of_Run_7239
  • 334600 10/13/10 08:04 jroche owl shift summary (better formating )
  • 334599 10/13/10 08:02 jroche owl shift summary
  • 334598 10/13/10 08:00 adaq Beam Time Accounting
  • 334597 10/13/10 06:57 a-onl Start_Run_7239,Run_type=Optics,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=50,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=S0 in the trigger (ps2=25)
  • 334596 10/13/10 06:56 a-onl End_of_Run_7238
  • 334595 10/13/10 06:15 jroche target is LH2 not C foils
  • 334594 10/13/10 06:11 jroche Scintillator timing (S2 only in the trigger)
  • 334593 10/13/10 05:21 a-onl Start_Run_7238,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=off,beam_current=50,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=S0 in the trigger (ps2=25)
  • 334592 10/13/10 04:59 yez target move to LH2 15cm Loop2
  • 334591 10/13/10 04:44 yez target move to empty
  • 334590 10/13/10 04:12 yez target move to Loop2 LH2 15cm
  • 334589 10/13/10 04:03 yez target moves to LH2 15cm
  • 334588 10/13/10 03:46 a-onl End_of_Run_7237
  • 334587 10/13/10 03:28 yez target screem2 - cryo gui
  • 334586 10/13/10 03:27 yez target screen 1 - L1,L2
  • 334585 10/13/10 02:50 a-onl Start_Run_7237,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=off,beam_current=50,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=S0 in the trigger (ps2=25)
  • 334584 10/13/10 02:47 auto_snap Radiation and Ozone snapshot
  • 334583 10/13/10 02:47 auto_snap Miscellaneous snapshot
  • 334582 10/13/10 02:47 auto_snap Hall A Tools snapshot
  • 334581 10/13/10 02:47 auto_snap Compton Polarimeter snapshot
  • 334580 10/13/10 02:47 auto_snap Beam Position and Current snapshot
  • 334579 10/13/10 02:44 jroche sieve back in position
  • 334578 10/13/10 02:40 yez Cannot see Multi-foils partten with Seive Slit in, but it is clear when Sieve Slit is out
  • 334577 10/13/10 02:24 a-onl Start_Run_7236,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=off,beam_current=50,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=S0 in the trigger (ps2=25)
  • 334576 10/13/10 02:23 a-onl End_of_Run_7235
  • 334575 10/13/10 02:15 a-onl Start_Run_7235,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=off,beam_current=50,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=S0 in the trigger (ps2=25)
  • 334574 10/13/10 02:15 a-onl End_of_Run_7234
  • 334573 10/13/10 02:15 a-onl Start_Run_7234,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=off,beam_current=50,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=S0 in the trigger (ps2=10)
  • 334572 10/13/10 02:14 a-onl End_of_Run_7233
  • 334571 10/13/10 02:13 a-onl Start_Run_7233,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=off,beam_current=50,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=S0 in the trigger
  • 334570 10/13/10 02:05 a-onl End_of_Run_7232
  • 334569 10/13/10 01:20 a-onl Start_Run_7232,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=off,beam_current=50,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=
  • 334568 10/13/10 01:20 a-onl Start_Run_7231,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=off,beam_current=50,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=
  • 334567 10/13/10 01:14 jroche sieve out
  • 334566 10/13/10 01:07 a-onl End_of_Run_7231
  • 334565 10/13/10 00:11 a-onl End_of_Run_7230
  • 334564 10/13/10 00:06 adaq Beam Time Accounting
  • 334563 10/13/10 00:05 Solvignon Shift summary
  • Tuesday

  • 334562 10/12/10 23:21 a-onl Start_Run_7230,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=off,beam_current=50,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=
  • 334561 10/12/10 23:20 a-onl End_of_Run_7229
  • 334560 10/12/10 22:13 a-onl Start_Run_7229,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=off,beam_current=15,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=
  • 334559 10/12/10 22:00 mcanan Target motion
  • 334558 10/12/10 21:53 mcanan Target Motion
  • 334557 10/12/10 21:52 Solvignon Starting beam recovery procedure
  • 334556 10/12/10 21:47 mcanan Target motion
  • 334555 10/12/10 21:10 mcanan Target Screens 10_12 Swing
  • 334554 10/12/10 20:18 Solvignon Starting compton test with different beam currents
  • 334553 10/12/10 20:00 efuchey Compton Rates after tuning
  • 334552 10/12/10 19:04 Solvignon Compton tune in progress
  • 334551 10/12/10 18:46 auto_snap Radiation and Ozone snapshot
  • 334550 10/12/10 18:46 auto_snap Miscellaneous snapshot
  • 334549 10/12/10 18:46 auto_snap Hall A Tools snapshot
  • 334548 10/12/10 18:46 auto_snap Compton Polarimeter snapshot
  • 334547 10/12/10 18:46 auto_snap Beam Position and Current snapshot
  • 334546 10/12/10 17:45 brads Helicity flip at 943.3Hz
  • 334545 10/12/10 17:42 brads Run Plan for Tues 12 SWING -> Wed 13 DAY
  • 334544 10/12/10 17:37 a-onl Start_Run_7228,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Test roc 17 and 14
  • 334543 10/12/10 17:36 Solvignon cw beam
  • 334542 10/12/10 17:31 mcanan Hall C target
  • 334541 10/12/10 17:29 mcanan Loop 2 hp power and JT Valve
  • 334540 10/12/10 17:27 brads re: Carbon foils operational current limit: 25 uA
  • 334539 10/12/10 17:23 Solvignon Beam tuning
  • 334538 10/12/10 17:18 a-onl End_of_Run_7226
  • 334537 10/12/10 17:18 a-onl Start_Run_7226,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Test roc 17 and 14
  • 334536 10/12/10 17:07 Solvignon/Ca EPICS gateway problem
  • 334535 10/12/10 17:00 a-onl Start_Run_7224,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Test
  • 334534 10/12/10 16:50 A. Camsonne DVCS DAQ access
  • 334533 10/12/10 16:49 J. Zhang Moeller magnet has been filled.
  • 334532 10/12/10 16:42 a-onl End_of_Run_7223
  • 334531 10/12/10 16:38 Solvignon Helicity flipping rate is now 960Hz
  • 334530 10/12/10 16:36 M Friend Compton DAQ trigger back to normal
  • 334529 10/12/10 16:34 mcanan loop 2 minor alarm
  • 334528 10/12/10 16:32 a-onl Start_Run_7223,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Test roc 14
  • 334527 10/12/10 16:21 Solvignon Hall A in controlled access
  • 334526 10/12/10 16:15 a-onl End_of_Run_7222
  • 334525 10/12/10 16:06 adaq Beam Time Accounting
  • 334524 10/12/10 15:57 a-onl Start_Run_7222,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Test roc 17
  • 334523 10/12/10 15:54 a-onl End_of_Run_7221
  • 334522 10/12/10 15:46 a-onl Start_Run_7221,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Test roc 17
  • 334521 10/12/10 15:24 adev End_of_Run_20016
  • 334520 10/12/10 11:18 Carlos BCM calibration aborted
  • 334519 10/12/10 11:17 a-onl End_of_Run_7220
  • 334518 10/12/10 11:11 a-onl End_of_Run_7219
  • 334517 10/12/10 11:08 a-onl Start_Run_7219,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Test roc 17
  • 334516 10/12/10 10:57 a-onl Start_Run_7217,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Test roc 18
  • 334515 10/12/10 10:55 a-onl End_of_Run_7216
  • 334514 10/12/10 10:50 a-onl Start_Run_7216,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Test roc 14
  • 334513 10/12/10 10:50 a-onl End_of_Run_7202
  • 334512 10/12/10 10:47 a-onl End_of_Run_7215
  • 334511 10/12/10 10:46 auto_snap Radiation and Ozone snapshot
  • 334510 10/12/10 10:46 auto_snap Miscellaneous snapshot
  • 334509 10/12/10 10:46 auto_snap Hall A Tools snapshot
  • 334508 10/12/10 10:46 auto_snap Compton Polarimeter snapshot
  • 334507 10/12/10 10:46 auto_snap Beam Position and Current snapshot
  • 334506 10/12/10 10:44 adev Start_Run_20016,Run_type=Cosmics,target_type=No Target,beam_e0=0,beam_raster=on,beam_halfwaveplate=in,rightarm_p=0,rightarm_theta=-70,comment_text=cosmics
  • 334505 10/12/10 10:41 a-onl Start_Run_7215,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Test roc 14
  • 334504 10/12/10 10:41 a-onl End_of_Run_7214
  • 334503 10/12/10 10:33 a-onl Start_Run_7214,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Test roc 14
  • 334502 10/12/10 10:32 a-onl End_of_Run_7213
  • 334501 10/12/10 10:20 a-onl Start_Run_7213,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Test roc 14
  • 334500 10/12/10 08:58 adaq Beam Time Accounting
  • 334499 10/12/10 08:44 Carlos BCM calibration setup
  • 334498 10/12/10 08:40 Carlos BCM calibration starting
  • 334497 10/12/10 08:39 X. Zheng Target moved to empty
  • 334496 10/12/10 08:38 a-onl End_of_Run_7212
  • 334495 10/12/10 08:26 a-onl Start_Run_7212,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=elastic scan @ dp =0%, sieve in, as in the prvious run.
  • 334494 10/12/10 08:15 jroche Owl shift summary
  • 334493 10/12/10 07:54 jroche changing conditions: still looking for a reason
  • 334492 10/12/10 07:53 a-onl End_of_Run_7211
  • 334491 10/12/10 06:41 a-onl Start_Run_7211,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=elastic scan @ dp =0%, sieve in, as in the prvious run.
  • 334490 10/12/10 06:39 a-onl End_of_Run_7210
  • 334489 10/12/10 06:34 jroche beam operation autorization
  • 334488 10/12/10 05:57 yez Graphs were all the same. Please check the most updated ones
  • 334487 10/12/10 05:29 a-onl Start_Run_7210,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=elastic scan @ dp =0%, sieve in, as in the prvious run.
  • 334486 10/12/10 04:59 a-onl End_of_Run_7209
  • 334485 10/12/10 03:40 a-onl Start_Run_7209,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=elastic scan @ dp =0%, sieve in.
  • 334484 10/12/10 03:33 jroche sieve in position
  • 334483 10/12/10 03:21 florian itar target
  • 334482 10/12/10 03:19 a-onl End_of_Run_7208
  • 334481 10/12/10 03:16 jroche Carbon foils operational current limit: 25 uA
  • 334480 10/12/10 03:04 a-onl Start_Run_7208,Run_type=Elastic Calibration,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=on,beam_current=2-10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=elastic scan @ dp =0%, sieve out.
  • 334479 10/12/10 02:57 jroche Left HRS to 3.327 GeV
  • 334478 10/12/10 02:46 auto_snap Radiation and Ozone snapshot
  • 334477 10/12/10 02:46 auto_snap Miscellaneous snapshot
  • 334476 10/12/10 02:46 auto_snap Hall A Tools snapshot
  • 334475 10/12/10 02:46 auto_snap Compton Polarimeter snapshot
  • 334474 10/12/10 02:46 auto_snap Beam Position and Current snapshot
  • 334473 10/12/10 02:28 jroche Q1 mistmatch fixed
  • 334472 10/12/10 02:28 jroche Q1 mistmatch fixed
  • 334471 10/12/10 02:13 jroche Q1 magnet mismatch??
  • 334470 10/12/10 02:08 a-onl End_of_Run_7207
  • 334469 10/12/10 01:57 A. Camsonne Trigger VME interface stuck
  • 334468 10/12/10 01:52 a-onl Start_Run_7207,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=2-10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=elastic scan @ dp = -4%, sieve out. ps1=25 all others at 0
  • 334467 10/12/10 01:51 a-onl End_of_Run_7206
  • 334466 10/12/10 01:50 a-onl Start_Run_7206,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=2-10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=elastic scan @ dp = -4%, sieve out. ps1=1 all others at 0
  • 334465 10/12/10 01:47 a-onl End_of_Run_7205
  • 334464 10/12/10 01:07 M Friend Compton background rates
  • 334463 10/12/10 00:59 A. Camsonne VME bus analyzer unplugged
  • 334462 10/12/10 00:40 A. Camsonne FPGA order on USB2
  • 334461 10/12/10 00:01 adaq Beam Time Accounting
  • 334460 10/12/10 00:01 Jin Huang Elastic stripe, dp = -4%
  • Monday

  • 334459 10/11/10 23:57 Jin Huang Shift Summary : Oct 11 Swing
  • 334458 10/11/10 23:48 yez Proton Mass Spectrum for Elastic Data with different Center Momentum Setting
  • 334457 10/11/10 23:46 a-onl Start_Run_7205,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=2-10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=elastic scan @ dp = -4%, sieve out. Start with 2uA, ramping up to 10uA
  • 334456 10/11/10 23:45 Jin Huang Hall A tool & HV @ dp = -4% elastic scan
  • 334455 10/11/10 23:43 a-onl End_of_Run_7204
  • 334454 10/11/10 23:43 Jin Huang Spot++ @ LH2, 2uA CW
  • 334453 10/11/10 23:40 BoZhao screen of target charts
  • 334452 10/11/10 23:40 yez Beam Spot on BeO with Raster off
  • 334451 10/11/10 23:39 BoZhao target moved to LH2@15cm
  • 334450 10/11/10 23:38 a-onl Start_Run_7204,Run_type=Other,target_type=BeO,beam_e0=5.5,beam_raster=on,beam_current=5,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Tuning
  • 334449 10/11/10 23:37 a-onl End_of_Run_7203
  • 334448 10/11/10 23:36 Jin Huang Spot++, 5uA CW on BeO
  • 334447 10/11/10 23:35 a-onl Start_Run_7203,Run_type=Other,target_type=Empty loop,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Tuning
  • 334446 10/11/10 23:28 Jin Spot++ with tune beam on LH2
  • 334445 10/11/10 23:24 BoZhao target moved to BeO
  • 334444 10/11/10 23:22 a-onl Start_Run_7202,Run_type=Other,target_type=Empty loop,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Tuning
  • 334443 10/11/10 23:15 BoZhao Target moved to loop2 LH2@15cm
  • 334442 10/11/10 23:11 a-onl Start_Run_7201,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test top crate - ROC17
  • 334441 10/11/10 23:02 a-onl Start_Run_7200,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test top crate - ROC18
  • 334440 10/11/10 22:58 a-onl End_of_Run_7199
  • 334439 10/11/10 22:50 a-onl Start_Run_7199,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test top crate - ROC14
  • 334438 10/11/10 22:44 a-onl End_of_Run_7198
  • 334437 10/11/10 22:43 a-onl Start_Run_7198,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test top crate - ROC17
  • 334436 10/11/10 22:39 a-onl End_of_Run_7197
  • 334435 10/11/10 22:11 Jin Huang Update: Beam is back
  • 334434 10/11/10 21:40 BoZhao screen of target charts
  • 334433 10/11/10 20:11 Jin Huang Update : No Beam. Asked for beam permit
  • 334432 10/11/10 19:44 A. Camsonne Pulses in ARS
  • 334431 10/11/10 19:37 a-onl Start_Run_7196,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bootom crate - ROC14
  • 334430 10/11/10 19:20 Alex ARS boards in crates
  • 334429 10/11/10 19:20 Alex
  • 334428 10/11/10 19:02 A. Camsonne FPGA programmers assignement
  • 334427 10/11/10 18:46 auto_snap Radiation and Ozone snapshot
  • 334426 10/11/10 18:45 auto_snap Miscellaneous snapshot
  • 334425 10/11/10 18:45 auto_snap Hall A Tools snapshot
  • 334424 10/11/10 18:45 auto_snap Compton Polarimeter snapshot
  • 334423 10/11/10 18:45 auto_snap Beam Position and Current snapshot
  • 334422 10/11/10 18:29 hyde DVCS cables under door
  • 334421 10/11/10 17:55 M Friend Temporary Compton DAQ trigger change
  • 334420 10/11/10 17:52 BoZhao screen of target charts
  • 334419 10/11/10 17:19 a-onl End_of_Run_7195
  • 334418 10/11/10 17:13 a-onl Start_Run_7195,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test crate 2 crates - lower rate
  • 334417 10/11/10 16:41 riordan Latest update is beam at 7:00PM
  • 334416 10/11/10 16:04 adaq Beam Time Accounting
  • 334415 10/11/10 16:01 a-onl End_of_Run_7192
  • 334414 10/11/10 16:01 a-onl Start_Run_7192,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test crate 2 crates
  • 334413 10/11/10 16:01 a-onl Start_Run_7189,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test crate DVCS
  • 334412 10/11/10 16:00 a-onl End_of_Run_7191
  • 334411 10/11/10 16:00 a-onl Start_Run_7191,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test crate top crates
  • 334410 10/11/10 15:59 BoZhao Screen of target charts
  • 334409 10/11/10 15:55 a-onl End_of_Run_7190
  • 334408 10/11/10 15:54 a-onl Start_Run_7190,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test crate 2 crates DVCS
  • 334407 10/11/10 15:43 a-onl End_of_Run_7189
  • 334406 10/11/10 15:43 a-onl Start_Run_7187,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test crate DVCS
  • 334405 10/11/10 15:42 a-onl End_of_Run_7188
  • 334404 10/11/10 15:39 a-onl Start_Run_7188,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test crate DVCS
  • 334403 10/11/10 15:38 a-onl End_of_Run_7187
  • 334402 10/11/10 15:36 riordan Left arm crate HV off
  • 334401 10/11/10 14:59 Holmstrom Target CommunicationPrb?
  • 334400 10/11/10 14:39 Kalyan follow-up on lumi analysis
  • 334399 10/11/10 14:05 A. Camsonne New version of ARS firmware
  • 334398 10/11/10 13:31 J. Butler Moeller
  • 334397 10/11/10 13:24 Jack (Techs) Left dipole NMR stability
  • 334396 10/11/10 12:24 Kalyan Re: Correction: Lumi linearity
  • 334395 10/11/10 11:40 Kalyan Correction: Lumi linearity
  • 334394 10/11/10 10:45 auto_snap Radiation and Ozone snapshot
  • 334393 10/11/10 10:45 auto_snap Miscellaneous snapshot
  • 334392 10/11/10 10:45 auto_snap Hall A Tools snapshot
  • 334391 10/11/10 10:45 auto_snap Compton Polarimeter snapshot
  • 334390 10/11/10 10:45 auto_snap Beam Position and Current snapshot
  • 334389 10/11/10 10:35 A. Camsonne DVCS crates boot script modified
  • 334388 10/11/10 09:04 O.Glamazdin Moller magnet is back to operation
  • 334387 10/11/10 08:51 riordan Rebooted iochla14
  • 334386 10/11/10 08:40 riordan Gateway restored but p0 setting not reading
  • 334385 10/11/10 08:29 riordan Injector work during day shift
  • 334384 10/11/10 08:25 adaq Beam Time Accounting
  • 334383 10/11/10 05:10 Hisham Owl shift summary
  • 334382 10/11/10 04:43 Hisham Owl shift summary
  • 334381 10/11/10 04:28 A. Camsonne EPICS gateway issues affect magnets regulation - giving up on beam for now
  • 334380 10/11/10 04:19 riordan Stopping for tonight - hall to controlled access
  • 334379 10/11/10 04:07 a-onl Start_Run_7185,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test top crate - 0x12 47a925 #1
  • 334378 10/11/10 04:07 Hisham Spot++
  • 334377 10/11/10 04:04 a-onl End_of_Run_7184
  • 334376 10/11/10 04:04 Hisham Beam is back
  • 334375 10/11/10 04:03 a-onl Start_Run_7184,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test top crate - 0x12 47a925 #1
  • 334374 10/11/10 03:57 Hisham NMR is locked
  • 334373 10/11/10 03:20 A. Camsonne Swing shift DAQ work
  • 334372 10/11/10 03:17 A. Camsonne Hall A Epics gateway down
  • 334371 10/11/10 02:50 riordan IOCs down
  • 334370 10/11/10 02:45 auto_snap Radiation and Ozone snapshot
  • 334369 10/11/10 02:45 auto_snap Miscellaneous snapshot
  • 334368 10/11/10 02:45 auto_snap Hall A Tools snapshot
  • 334367 10/11/10 02:45 auto_snap Compton Polarimeter snapshot
  • 334366 10/11/10 02:45 auto_snap Beam Position and Current snapshot
  • 334365 10/11/10 02:43 Hisham NMR is not locked
  • 334364 10/11/10 02:09 hisham Controlled acces
  • 334363 10/11/10 02:08 hisham Late entry: Beam is back
  • 334362 10/11/10 02:07 a-onl End_of_Run_7183
  • 334361 10/11/10 02:06 hisham Late entry: Out of the Hall
  • 334360 10/11/10 02:03 a-onl Start_Run_7183,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test top crate - 0x12 47a925 #1
  • 334358 10/11/10 01:09 a-onl Start_Run_7182,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test top crate - 0x12 47a925 #1
  • 334357 10/11/10 01:02 a-onl End_of_Run_7181
  • 334356 10/11/10 00:55 a-onl Start_Run_7181,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test top crate - 0x1 0x12 47a925 #3
  • 334355 10/11/10 00:55 a-onl End_of_Run_7180
  • 334354 10/11/10 00:51 a-onl Start_Run_7180,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test top crate - 0x1 0x12 47a925
  • 334353 10/11/10 00:49 a-onl End_of_Run_7179
  • 334352 10/11/10 00:40 a-onl Start_Run_7179,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test top crate - 0x1 0x12 47a925
  • 334351 10/11/10 00:36 Hisham Screen snapshot: NMR
  • 334350 10/11/10 00:31 a-onl Start_Run_7174,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board 0x12 firmware E06 #3
  • 334349 10/11/10 00:16 a-onl End_of_Run_7177
  • 334348 10/11/10 00:12 a-onl Start_Run_7177,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board 0x12 firmware E06 #3 - board 16e not in crate
  • 334347 10/11/10 00:12 a-onl End_of_Run_7176
  • 334346 10/11/10 00:08 a-onl Start_Run_7176,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board 0x12 firmware E06 #2 - board 16e not in crate
  • 334345 10/11/10 00:06 a-onl End_of_Run_7175
  • 334344 10/11/10 00:03 adaq Beam Time Accounting
  • 334343 10/11/10 00:01 a-onl Start_Run_7175,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board 0x12 firmware E06 #1 - board 16e not in crate
  • 334342 10/11/10 00:01 a-onl End_of_Run_7172
  • 334341 10/11/10 00:00 LeRose Shift Summary (Swing)
  • Sunday

  • 334340 10/10/10 23:56 a-onl End_of_Run_7174
  • 334339 10/10/10 23:54 a-onl End_of_Run_7173
  • 334338 10/10/10 23:53 a-onl Start_Run_7173,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board 0x12 firmware E06 #2
  • 334337 10/10/10 23:51 a-onl Start_Run_7172,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board 0x12 firmware E06 #1
  • 334336 10/10/10 23:48 a-onl End_of_Run_7171
  • 334335 10/10/10 23:48 a-onl Start_Run_7171,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board 1 firmware E06 #1 stop on scope
  • 334334 10/10/10 23:47 jroche Vertex resolution with new t0 for vdc
  • 334333 10/10/10 23:40 a-onl End_of_Run_7170
  • 334332 10/10/10 23:37 a-onl Start_Run_7170,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board 8 firmware E06 #3
  • 334331 10/10/10 23:35 a-onl End_of_Run_7169
  • 334330 10/10/10 23:31 a-onl Start_Run_7169,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board 8 firmware E06 #2
  • 334329 10/10/10 23:31 a-onl Start_Run_7167,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board 8 firmware E06 Junk
  • 334328 10/10/10 23:31 a-onl End_of_Run_7168
  • 334327 10/10/10 23:27 a-onl Start_Run_7168,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board 8 firmware E06 #1
  • 334326 10/10/10 23:25 a-onl End_of_Run_7164
  • 334325 10/10/10 23:25 a-onl End_of_Run_7167
  • 334324 10/10/10 23:21 a-onl End_of_Run_7166
  • 334323 10/10/10 23:17 jroche Cerenkov PMT1 now showing a nice pe peak
  • 334322 10/10/10 23:13 a-onl Start_Run_7166,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board f firmware E06 #3
  • 334321 10/10/10 23:11 a-onl Start_Run_7163,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board e firmware E06 #3 junk
  • 334320 10/10/10 23:09 a-onl End_of_Run_7165
  • 334319 10/10/10 23:06 a-onl Start_Run_7165,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board f firmware E06 #2
  • 334318 10/10/10 22:59 a-onl Start_Run_7164,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board f firmware E06 #1
  • 334317 10/10/10 22:56 a-onl Start_Run_7158,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=test top crate
  • 334316 10/10/10 22:56 a-onl End_of_Run_7163
  • 334315 10/10/10 22:51 a-onl End_of_Run_7162
  • 334314 10/10/10 22:47 a-onl Start_Run_7162,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board e firmware E06 #3
  • 334313 10/10/10 22:42 a-onl End_of_Run_7161
  • 334312 10/10/10 22:37 a-onl Start_Run_7161,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board e firmware E06 #2
  • 334311 10/10/10 22:36 a-onl End_of_Run_7160
  • 334310 10/10/10 22:29 a-onl Start_Run_7160,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board e firmware E06 #1
  • 334309 10/10/10 22:08 a-onl End_of_Run_7159
  • 334308 10/10/10 22:04 a-onl Start_Run_7159,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Test bottom crate board d firmware E06 #3
  • 334307 10/10/10 22:03 a-onl End_of_Run_7158
  • 334306 10/10/10 21:54 a-onl Start_Run_7151,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board d firmware E06 #2 after reboot
  • 334305 10/10/10 21:53 a-onl End_of_Run_7148
  • 334304 10/10/10 21:51 a-onl Start_Run_7148,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board d firmware E06 #2
  • 334303 10/10/10 21:48 a-onl Start_Run_7150,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board d firmware E06 #2
  • 334302 10/10/10 21:45 a-onl Start_Run_7149,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board d firmware E06 #2
  • 334301 10/10/10 21:42 a-onl End_of_Run_7146
  • 334300 10/10/10 21:37 a-onl End_of_Run_7147
  • 334299 10/10/10 21:31 a-onl Start_Run_7147,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board d firmware E06 #3
  • 334298 10/10/10 21:30 a-onl Start_Run_7146,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board d firmware E06 #3
  • 334297 10/10/10 21:30 a-onl Start_Run_7145,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 6 firmware E06 #3
  • 334296 10/10/10 21:18 a-onl End_of_Run_7145
  • 334295 10/10/10 21:16 a-onl Start_Run_7136,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board e only #1
  • 334294 10/10/10 21:12 a-onl Start_Run_7144,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 6 firmware E06 #3
  • 334293 10/10/10 21:12 a-onl End_of_Run_7144
  • 334292 10/10/10 21:06 a-onl End_of_Run_7143
  • 334291 10/10/10 21:01 a-onl Start_Run_7143,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 6 firmware E06 #3
  • 334290 10/10/10 20:55 a-onl End_of_Run_7142
  • 334289 10/10/10 20:51 a-onl Start_Run_7142,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 1 firmware E06 #3
  • 334288 10/10/10 20:50 a-onl End_of_Run_7141
  • 334287 10/10/10 20:45 a-onl Start_Run_7141,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 1 firmware E06 #2
  • 334286 10/10/10 20:34 a-onl Start_Run_7140,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 1 firmware E06 #1
  • 334285 10/10/10 20:31 a-onl End_of_Run_7138
  • 334284 10/10/10 20:30 a-onl End_of_Run_7138
  • 334283 10/10/10 20:27 a-onl Start_Run_7138,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board d firmware E06
  • 334282 10/10/10 20:21 a-onl End_of_Run_7137
  • 334281 10/10/10 20:20 a-onl Start_Run_7137,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 16def8
  • 334280 10/10/10 20:18 A. Camsonne DVCS bottom crate firmware update from file E06.pof
  • 334279 10/10/10 19:58 Jin Huang Target: 2nd temperature sensor on loop2 just fixed itself
  • 334278 10/10/10 19:53 a-onl End_of_Run_7136
  • 334277 10/10/10 19:49 a-onl End_of_Run_7135
  • 334276 10/10/10 19:48 a-onl Start_Run_7133,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board d only #1
  • 334275 10/10/10 19:45 a-onl Start_Run_7135,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board d only #3
  • 334274 10/10/10 19:44 a-onl End_of_Run_7134
  • 334273 10/10/10 19:40 a-onl Start_Run_7134,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board d only #2
  • 334272 10/10/10 19:39 a-onl End_of_Run_7133
  • 334271 10/10/10 19:34 a-onl End_of_Run_7127
  • 334270 10/10/10 19:29 a-onl End_of_Run_7132
  • 334269 10/10/10 19:26 a-onl Start_Run_7132,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 6 only #3
  • 334268 10/10/10 19:26 a-onl End_of_Run_7131
  • 334267 10/10/10 19:22 a-onl Start_Run_7131,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 6 only #2
  • 334266 10/10/10 19:22 a-onl End_of_Run_7130
  • 334265 10/10/10 19:16 a-onl Start_Run_7130,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 6 only good prescale #1
  • 334264 10/10/10 19:10 a-onl End_of_Run_7129
  • 334263 10/10/10 19:07 a-onl Start_Run_7129,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 1 only good prescale #3
  • 334262 10/10/10 19:07 a-onl End_of_Run_7126
  • 334261 10/10/10 19:06 a-onl End_of_Run_7128
  • 334260 10/10/10 19:01 a-onl Start_Run_7128,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 1 only good prescale #2
  • 334259 10/10/10 18:56 a-onl Start_Run_7127,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 1 only good prescale
  • 334258 10/10/10 18:55 a-onl Start_Run_7126,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 1 only good prescale
  • 334257 10/10/10 18:52 a-onl End_of_Run_7125
  • 334256 10/10/10 18:49 a-onl Start_Run_7125,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS tests bottom crate board 1 only
  • 334255 10/10/10 18:45 auto_snap Radiation and Ozone snapshot
  • 334254 10/10/10 18:45 auto_snap Miscellaneous snapshot
  • 334253 10/10/10 18:45 auto_snap Hall A Tools snapshot
  • 334252 10/10/10 18:44 auto_snap Compton Polarimeter snapshot
  • 334251 10/10/10 18:44 auto_snap Beam Position and Current snapshot
  • 334250 10/10/10 18:14 L Controlled Access
  • 334249 10/10/10 17:51 riordan Run Plan for Swing 10/10 through Day 10/11
  • 334248 10/10/10 17:37 LeRose Controlled Access
  • 334247 10/10/10 17:11 Carlos ARS tests (run 7116)
  • 334246 10/10/10 17:09 Kalyan Current scan for lumi detector
  • 334245 10/10/10 17:08 Jin Huang Target -> Empty
  • 334244 10/10/10 17:01 a-onl End_of_Run_7124
  • 334243 10/10/10 16:53 Carlos ARS tests
  • 334242 10/10/10 16:48 A. Camsonne crl files copied in Oct-10-2010
  • 334241 10/10/10 16:46 A. Camsonne crl files copied in Sept9-2010
  • 334240 10/10/10 16:26 Jin Huang Target: Heater upper limit from 300W->500W for both loops
  • 334239 10/10/10 16:10 adaq Beam Time Accounting
  • 334238 10/10/10 16:00 jmammei Day shift summary
  • 334237 10/10/10 15:51 a-onl Start_Run_7124,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Elastic calibration - dp = -4%
  • 334236 10/10/10 15:48 a-onl End_of_Run_7123
  • 334235 10/10/10 15:46 a-onl Start_Run_7123,Run_type=Spot++,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Spot check
  • 334234 10/10/10 15:43 A. Camsonne Spot++ run 7123 on LH2 5uA
  • 334233 10/10/10 15:32 Bellini Target move to LH2 15cm
  • 334232 10/10/10 15:27 A. Camsonne Raster check run 7122
  • 334231 10/10/10 15:17 a-onl End_of_Run_7120
  • 334230 10/10/10 15:09 a-onl Start_Run_7120,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS pulser run - two crates ROC17 47a3925 ROC18 def6 - lower rate
  • 334229 10/10/10 14:42 jmammei raster was reset to 4x4mm^2
  • 334228 10/10/10 14:24 jmammei turned spectrometer HV on
  • 334227 10/10/10 14:23 O.Glamazdin Moller magnet
  • 334226 10/10/10 14:13 jmammei Sasha out of hall
  • 334225 10/10/10 14:08 jmammei controlled access for MOLLER
  • 334224 10/10/10 13:27 jpchen temperature spikes due to noise
  • 334223 10/10/10 13:06 a-onl End_of_Run_7115
  • 334222 10/10/10 12:55 a-onl Start_Run_7115,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS pulser run - two crates ROC17 47a3925 ROC18 def6 - lower rate
  • 334221 10/10/10 12:55 a-onl End_of_Run_7114
  • 334220 10/10/10 12:52 jmammei Carlos came out of hall
  • 334219 10/10/10 12:52 Carlos Pulse signal plugged into lower ARS crate
  • 334218 10/10/10 12:32 jmammei Jixie came out of Hall
  • 334217 10/10/10 12:31 J. Zhang Moller magnet has been refilled
  • 334216 10/10/10 12:28 jmammei raster decreased for MOLLER
  • 334215 10/10/10 12:27 jmammei Jixie and Carlos go in for escorted access
  • 334214 10/10/10 12:13 Bellini Target screens
  • 334213 10/10/10 12:12 jmammei cryo temp drop
  • 334212 10/10/10 12:03 a-onl Start_Run_7114,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS pulser run - two crates ROC17 47a3925 ROC18 def6 - lower rate
  • 334211 10/10/10 12:03 a-onl End_of_Run_7113
  • 334210 10/10/10 11:54 a-onl End_of_Run_7112
  • 334209 10/10/10 11:50 Bellini Transient red alarm on Loop1
  • 334208 10/10/10 11:47 Bellini Screen snapshot
  • 334207 10/10/10 11:43 a-onl Start_Run_7112,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=ARS pulser run - two crates ROC17 47a3925 ROC18 def6
  • 334206 10/10/10 11:36 A. Camsonne ROC18 crashed
  • 334205 10/10/10 11:27 jmammei calorimeter HV back off
  • 334204 10/10/10 11:25 a-onl End_of_Run_7111
  • 334203 10/10/10 11:24 a-onl Start_Run_7111,Run_type=Spot++,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334202 10/10/10 11:14 jmammei turned off spectrometer HV
  • 334201 10/10/10 11:13 a-onl Start_Run_7110,Run_type=Spot++,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334200 10/10/10 10:56 Magali run on ARS Board at the electronic Lab
  • 334199 10/10/10 10:48 jmammei preparing for MOLLER tests
  • 334198 10/10/10 10:44 auto_snap Radiation and Ozone snapshot
  • 334197 10/10/10 10:44 auto_snap Miscellaneous snapshot
  • 334196 10/10/10 10:44 auto_snap Hall A Tools snapshot
  • 334195 10/10/10 10:44 auto_snap Compton Polarimeter snapshot
  • 334194 10/10/10 10:44 auto_snap Beam Position and Current snapshot
  • 334193 10/10/10 10:32 Bellini Target moved to empty
  • 334192 10/10/10 09:37 Carlos Run pulser 7103
  • 334191 10/10/10 09:35 Carlos Run pulser 7103
  • 334190 10/10/10 08:58 jmammei NMR lock lost briefly - fixed itself
  • 334189 10/10/10 08:43 cshloka Cryo levels
  • 334188 10/10/10 08:00 adaq Beam Time Accounting
  • 334187 10/10/10 08:00 M Friend Owl shift summary
  • 334186 10/10/10 07:07 M Friend NMR lost lock temporarily
  • 334185 10/10/10 06:26 M Friend NMR temporarily lost lock again
  • 334184 10/10/10 05:42 M Friend NMR temporarily lost lock
  • 334183 10/10/10 05:13 D.Wang target moved Loop2, 15cm
  • 334182 10/10/10 05:05 itard spot++
  • 334181 10/10/10 05:05 D.Wang target moved to BeO
  • 334180 10/10/10 05:03 a-onl Start_Run_7109,Run_type=Spot++,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334179 10/10/10 04:31 D.Wang target moved to empty
  • 334178 10/10/10 04:31 a-onl End_of_Run_7108
  • 334177 10/10/10 03:47 a-onl Start_Run_7108,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334176 10/10/10 03:47 a-onl End_of_Run_7107
  • 334175 10/10/10 03:11 D.Wang target status
  • 334174 10/10/10 02:48 a-onl Start_Run_7107,Run_type=Cosmics,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334173 10/10/10 02:47 a-onl End_of_Run_7106
  • 334172 10/10/10 02:44 auto_snap Radiation and Ozone snapshot
  • 334171 10/10/10 02:44 auto_snap Miscellaneous snapshot
  • 334170 10/10/10 02:44 auto_snap Hall A Tools snapshot
  • 334169 10/10/10 02:44 auto_snap Compton Polarimeter snapshot
  • 334168 10/10/10 02:44 auto_snap Beam Position and Current snapshot
  • 334167 10/10/10 02:15 M Friend Beam down
  • 334166 10/10/10 01:48 a-onl Start_Run_7106,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334165 10/10/10 01:46 A. Camsonne DVCS DAQ status
  • 334164 10/10/10 01:34 D.Wang target status
  • 334163 10/10/10 01:32 a-onl End_of_Run_7103
  • 334162 10/10/10 01:31 M Friend Calorimeter HVs OFF
  • 334161 10/10/10 01:29 a-onl Start_Run_7103,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334160 10/10/10 01:12 M Friend Hall back to beam permit
  • 334159 10/10/10 01:11 M Friend Calorimeter HVs on
  • 334158 10/10/10 01:07 A. Camsonne Calorimeter VME interface board unplugged
  • 334157 10/10/10 01:03 M Friend Hall in Controlled access
  • 334156 10/10/10 01:01 a-onl End_of_Run_7102
  • 334155 10/10/10 00:56 a-onl Start_Run_7102,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334154 10/10/10 00:51 A. Camsonne Channels with pulser
  • 334153 10/10/10 00:50 a-onl End_of_Run_7101
  • 334152 10/10/10 00:40 M Friend Calorimeter HV GUI problem
  • 334151 10/10/10 00:36 a-onl Start_Run_7101,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334150 10/10/10 00:27 a-onl End_of_Run_7100
  • 334149 10/10/10 00:25 a-onl Start_Run_7100,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334148 10/10/10 00:24 M Friend GC ch. 1 HV to 1800V
  • 334147 10/10/10 00:05 adaq Beam Time Accounting
  • 334146 10/10/10 00:01 Jin Huang Shift Summary : Oct 09 Swing
  • Saturday

  • 334145 10/09/10 23:54 jroche new T0 for the Left VDC updated in the database
  • 334144 10/09/10 23:31 A. Camsonne DVCS calorimeter HV off
  • 334143 10/09/10 23:26 A. Camsonne Typo in onlana.odef
  • 334142 10/09/10 23:22 a-onl Start_Run_7099,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334141 10/09/10 23:22 a-onl End_of_Run_7098
  • 334140 10/09/10 23:21 a-onl Start_Run_7098,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334139 10/09/10 23:13 a-onl End_of_Run_7096
  • 334138 10/09/10 23:13 a-onl End_of_Run_7082
  • 334137 10/09/10 23:07 a-onl Start_Run_7096,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.194,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=p0 - 4%
  • 334136 10/09/10 23:02 a-onl End_of_Run_7095
  • 334135 10/09/10 23:00 Jin and Alex LHRS p0 -> 3.392GeV/c (-4%)
  • 334134 10/09/10 22:50 A. Camsonne NMR current setting
  • 334133 10/09/10 22:49 a-onl Start_Run_7095,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Calorimeter top crate boards 47a3925 - Calorimeter HV 400V - HRS Off - Pulser only
  • 334132 10/09/10 22:47 a-onl End_of_Run_7094
  • 334131 10/09/10 22:43 a-onl Start_Run_7094,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=Calorimeter top crate boards 47a3925 - Calorimeter HV 400V
  • 334130 10/09/10 22:42 A. Camsonne Left HRS HV off for run 7095
  • 334129 10/09/10 22:39 A. Camsonne Calorimeter random test top crate Run 7094
  • 334128 10/09/10 22:38 Jin & Alex CaF2 cal. ADC readback
  • 334127 10/09/10 22:31 A. Camsonne Calorimeter HV changed to 400 V and turned on for random data taking on calorimeter
  • 334126 10/09/10 22:28 Jin & Alex PbF2 Cal. ADC Read back
  • 334125 10/09/10 22:19 a-onl End_of_Run_7093
  • 334124 10/09/10 21:33 salvatore Cherenkov
  • 334123 10/09/10 21:14 a-onl Start_Run_7093,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=HV on gas cer PMT#2(count start from 1) was increased from -1750V tp -1775V. raster 5.5x4 sieve out p0=+0% Beam current=10.0uA
  • 334122 10/09/10 21:14 Jin Huang HV change on Gas Cer
  • 334121 10/09/10 21:12 a-onl End_of_Run_7092
  • 334120 10/09/10 21:09 jroche Cerenkov (partial) calibration
  • 334119 10/09/10 21:06 Jin Huang Update on run database (online replay)
  • 334118 10/09/10 20:54 Min Huang warning of haL1_cernox_3_R
  • 334117 10/09/10 20:33 Jin Huang Elastic Kinematics Check Plots
  • 334116 10/09/10 20:22 Jin Huang Beam current -> 10uA
  • 334115 10/09/10 20:11 a-onl Start_Run_7092,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+0% Beam current=5.0uA
  • 334114 10/09/10 20:06 Jin Huang Cryo flow back. LH2 stablized. Continue taking data
  • 334113 10/09/10 20:06 a-onl End_of_Run_7091
  • 334112 10/09/10 20:03 Jin Huang Stop beam due to drop of 15K cryo supply
  • 334111 10/09/10 20:00 Jin Huang JT valve change, beam current adjustment
  • 334110 10/09/10 19:53 a-onl Start_Run_7091,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=5,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+0% Beam current=5.0uA
  • 334109 10/09/10 19:52 a-onl End_of_Run_7090
  • 334108 10/09/10 19:49 a-onl Start_Run_7090,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+0% Beam current=10.0uA
  • 334107 10/09/10 19:47 a-onl End_of_Run_7089
  • 334106 10/09/10 18:48 A. Camsonne DVCS logic module updated
  • 334105 10/09/10 18:45 a-onl Start_Run_7089,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.327,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+0% Beam current=10.0uA
  • 334104 10/09/10 18:44 auto_snap Radiation and Ozone snapshot
  • 334103 10/09/10 18:44 auto_snap Miscellaneous snapshot
  • 334102 10/09/10 18:44 auto_snap Hall A Tools snapshot
  • 334101 10/09/10 18:44 auto_snap Compton Polarimeter snapshot
  • 334100 10/09/10 18:44 auto_snap Beam Position and Current snapshot
  • 334099 10/09/10 18:40 Jin Huang LHRS p0->3.327GeV/c, (+0% to elastic)
  • 334098 10/09/10 18:33 a-onl End_of_Run_7088
  • 334097 10/09/10 17:42 Jin Huang Open up JT valve of Loop2 (LH2) to 47.8%(+2%)
  • 334096 10/09/10 17:27 a-onl Start_Run_7088,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.427,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+3% Beam current=10.0uA
  • 334095 10/09/10 17:27 a-onl End_of_Run_7087
  • 334094 10/09/10 17:09 mhuang target screen shot
  • 334093 10/09/10 17:09 salvatore Cherenkov counter6
  • 334092 10/09/10 16:48 Kalyan Online lumi monitoring
  • 334091 10/09/10 16:31 adaq Beam Time Accounting
  • 334090 10/09/10 16:30 A. Camsonne Gas cerenkov channel 6 no pedestal or single photoelectron peak
  • 334089 10/09/10 16:26 a-onl Start_Run_7087,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.427,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+3% Beam current=10.0uA Lumi test
  • 334088 10/09/10 16:23 a-onl End_of_Run_7086
  • 334087 10/09/10 16:14 a-onl Start_Run_7086,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.427,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+3% Beam current=15.0uA Lumi test
  • 334086 10/09/10 16:12 a-onl End_of_Run_7085
  • 334085 10/09/10 16:05 Carlos Day shift summary
  • 334084 10/09/10 15:58 a-onl Start_Run_7085,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.427,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+3% Beam current=10.0uA Lumi test
  • 334083 10/09/10 15:56 a-onl End_of_Run_7084
  • 334082 10/09/10 15:46 a-onl Start_Run_7084,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.427,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+3% Beam current=5.0uA Lumi test
  • 334081 10/09/10 15:45 a-onl End_of_Run_7083
  • 334080 10/09/10 15:35 a-onl Start_Run_7083,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.427,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+3% Beam current=2.0uA Lumi test
  • 334079 10/09/10 15:33 a-onl End_of_Run_7081
  • 334078 10/09/10 15:32 adaq The beam current for run7082 is:1.0uA, not 0.5uA
  • 334077 10/09/10 15:23 a-onl Start_Run_7082,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.427,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+3% Beam current=0.5uA Lumi test
  • 334076 10/09/10 15:21 a-onl End_of_Run_7080
  • 334075 10/09/10 15:18 efuchey Hall A tools scrennshots
  • 334074 10/09/10 15:10 a-onl Start_Run_7081,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.427,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+3% Beam current=0.5uA Lumi test
  • 334073 10/09/10 15:09 Bellini Twin red alarm for Loop1
  • 334072 10/09/10 15:08 a-onl End_of_Run_7074
  • 334071 10/09/10 14:55 chench Are the missing pion rejection ADCs out of acceptance?
  • 334070 10/09/10 14:55 a-onl Start_Run_7080,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.427,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+3%
  • 334069 10/09/10 14:29 a-onl Start_Run_7079,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.427,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+3%
  • 334068 10/09/10 14:27 a-onl End_of_Run_7078
  • 334067 10/09/10 14:10 efuchey HRS NMR lock unstable
  • 334066 10/09/10 13:36 Kalyan Lumi HV settings for 10uA on H2 15cm (loop2)
  • 334065 10/09/10 13:30 a-onl Start_Run_7078,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.427,leftarm_theta=27.5,leftarm_sieve=out,calo_theta=22.0,calo_distance=5,comment_text=raster 5.5x4 sieve out p0=+3%
  • 334064 10/09/10 13:18 Bellini Move back to LH2 target
  • 334063 10/09/10 13:08 a-onl End_of_Run_7077
  • 334062 10/09/10 13:03 a-onl Start_Run_7077,Run_type=Spot++,target_type=Empty loop,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.427,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=27.5,calo_distance=5,comment_text=raster 5.5x4 sieve out
  • 334061 10/09/10 12:49 efuchey Issues with beam
  • 334060 10/09/10 12:46 Bellini Moving to empty target
  • 334059 10/09/10 12:26 efuchey HRS-L momentum change
  • 334058 10/09/10 11:54 efuchey HRS-L change
  • 334057 10/09/10 11:40 O.Glamazdin Moller SC magnet has been filled
  • 334056 10/09/10 11:32 Bellini HI and HIHI limits Change in Gas Panels
  • 334055 10/09/10 10:57 riordan Going to Controlled Access
  • 334054 10/09/10 10:54 Bellini / Ca Yellow persistent alarm on LH2
  • 334053 10/09/10 10:43 auto_snap Radiation and Ozone snapshot
  • 334052 10/09/10 10:43 auto_snap Miscellaneous snapshot
  • 334051 10/09/10 10:43 auto_snap Hall A Tools snapshot
  • 334050 10/09/10 10:43 auto_snap Compton Polarimeter snapshot
  • 334049 10/09/10 10:43 auto_snap Beam Position and Current snapshot
  • 334048 10/09/10 10:35 Bellini Red transient alarm on L1
  • 334047 10/09/10 10:35 Magali state of the ARS test in the Electronic Lab
  • 334046 10/09/10 10:34 Magali state of the ARS test in the Electronic Lab
  • 334045 10/09/10 10:33 riordan HRS to p0 = 3.775 GeV/c
  • 334044 10/09/10 09:54 a-onl End_of_Run_7076
  • 334043 10/09/10 09:39 a-onl Start_Run_7076,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.932,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster 5.5x4 sieve out
  • 334042 10/09/10 09:38 A. Camsonne Scaler rates LH2 10 uA after trigger download
  • 334041 10/09/10 09:30 A. Camsonne Trigger for Left HRS redownloaded
  • 334040 10/09/10 09:29 a-onl End_of_Run_7075
  • 334039 10/09/10 09:28 atrig Trigger-Download
  • 334038 10/09/10 09:11 a-onl Start_Run_7075,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.932,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster 5.5x4 sieve out
  • 334037 10/09/10 09:10 a-onl End_of_Run_7073
  • 334036 10/09/10 09:01 a-onl Start_Run_7074,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=3.932,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster 5.5x4 sieve out
  • 334035 10/09/10 08:57 Carlos Rates at p0=3.932 GeV sieve OUT
  • 334034 10/09/10 08:53 Bellini Target Move to LH2 15cm
  • 334033 10/09/10 08:50 Carlos Beam's back
  • 334032 10/09/10 08:40 Carlos New HRS setting
  • 334031 10/09/10 08:23 Carlos Sieve pattern
  • 334030 10/09/10 08:17 Bellini Target moved to empty
  • 334029 10/09/10 08:11 adaq Beam Time Accounting
  • 334028 10/09/10 08:10 Carlos Owl shift summary
  • 334027 10/09/10 08:09 Carlos Beam off for 30 min
  • 334026 10/09/10 07:53 Carlos Email test
  • 334025 10/09/10 07:27 riordan Too many factors to consider resolution right now
  • 334024 10/09/10 07:22 a-onl Start_Run_7073,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=4.,leftarm_theta=21.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=raster 5.5x4 sieve in
  • 334023 10/09/10 07:21 a-onl End_of_Run_7072
  • 334022 10/09/10 06:29 Carlos Rates with sieve IN
  • 334021 10/09/10 06:25 Carlos Rates at 10uA on 15cm LH2
  • 334020 10/09/10 06:22 Carlos Elastic scattering on LH2
  • 334019 10/09/10 06:21 D.Wang target status
  • 334018 10/09/10 06:13 a-onl Start_Run_7072,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=4.,leftarm_theta=21.5,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=raster 5.5x4
  • 334017 10/09/10 06:09 a-onl End_of_Run_7071
  • 334016 10/09/10 06:09 a-onl Start_Run_7065,Run_type=Spot++,target_type=Cross hair,beam_e0=,beam_raster=on,beam_current=5,leftarm_p=4.,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster check
  • 334015 10/09/10 05:39 A. Camsonne Vacuum stil pumping in HRS ?
  • 334014 10/09/10 05:08 a-onl Start_Run_7071,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=4.,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster 5.5x4
  • 334013 10/09/10 05:07 a-onl End_of_Run_7070
  • 334012 10/09/10 05:02 a-onl Start_Run_7070,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=4.,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster 5.5x4
  • 334011 10/09/10 04:55 a-onl End_of_Run_7069
  • 334010 10/09/10 04:51 Carlos Raster 7x7 on LH2
  • 334009 10/09/10 04:41 a-onl Start_Run_7069,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=4.,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster 7x7
  • 334008 10/09/10 04:37 a-onl End_of_Run_7068
  • 334007 10/09/10 04:35 Carlos spot++ on LH2
  • 334006 10/09/10 04:29 Carlos Vertex resolution
  • 334005 10/09/10 04:23 D.Wang Target status
  • 334004 10/09/10 04:20 a-onl Start_Run_7068,Run_type=Elastic Calibration,target_type=LH2,beam_e0=5.5,beam_raster=on,beam_current=10,leftarm_p=4.,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster 5.5x4
  • 334003 10/09/10 04:15 Carlos Optics reconstruction
  • 334002 10/09/10 04:08 a-onl End_of_Run_7067
  • 334001 10/09/10 04:07 D.Wang Target moved to Loop2(LH2), 15cm
  • 334000 10/09/10 03:08 a-onl Start_Run_7067,Run_type=Spot++,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=off,beam_current=10,leftarm_p=4.,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=optics check
  • 333999 10/09/10 03:00 a-onl End_of_Run_7066
  • 333998 10/09/10 02:57 D.Wang Target moved to Optics
  • 333997 10/09/10 02:57 Carlos Raster 5.5x4
  • 333996 10/09/10 02:43 auto_snap Radiation and Ozone snapshot
  • 333995 10/09/10 02:43 auto_snap Miscellaneous snapshot
  • 333994 10/09/10 02:43 auto_snap Hall A Tools snapshot
  • 333993 10/09/10 02:43 auto_snap Compton Polarimeter snapshot
  • 333992 10/09/10 02:43 auto_snap Beam Position and Current snapshot
  • 333991 10/09/10 02:30 a-onl Start_Run_7066,Run_type=Spot++,target_type=Cross hair,beam_e0=5.5,beam_raster=on,beam_current=5,leftarm_p=4.,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster check
  • 333990 10/09/10 02:28 a-onl End_of_Run_7065
  • 333989 10/09/10 02:16 Carlos spot++ 2x2
  • 333988 10/09/10 02:07 Carlos Calorimeter cosmics runs
  • 333987 10/09/10 02:02 a-onl Start_Run_7064,Run_type=Spot++,target_type=Cross hair,beam_e0=,beam_raster=on,beam_current=5,leftarm_p=4.,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster check
  • 333986 10/09/10 01:54 D.Wang Target moved to Carbon Hole
  • 333985 10/09/10 01:51 a-onl End_of_Run_7063
  • 333984 10/09/10 01:48 D.Wang target status
  • 333983 10/09/10 01:47 D.Wang target moved from empty to BeO
  • 333982 10/09/10 01:44 Carlos Calorimeter HV off
  • 333981 10/09/10 01:39 Carlos Beam ON
  • 333980 10/09/10 00:40 a-onl Start_Run_7063,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS cosmics test ROC 17 only
  • 333979 10/09/10 00:39 Carlos Sweep in progress
  • 333978 10/09/10 00:30 a-onl End_of_Run_7062
  • 333977 10/09/10 00:29 jpchen GUIs may still have wrong lables
  • 333976 10/09/10 00:26 jpchen target screens
  • 333975 10/09/10 00:24 a-onl Start_Run_7062,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS cosmics test ROC 17 only
  • 333974 10/09/10 00:21 jpchen cryotarget cooldown complete
  • 333973 10/09/10 00:13 adaq Beam Time Accounting
  • 333972 10/09/10 00:01 jroche for c S2 ADC basically already matched
  • Friday

  • 333971 10/08/10 23:37 jpchen Instruction for Switch between LH2 (loop2) and LD2 (loop1)
  • 333970 10/08/10 23:23 A. Camsonne DVCS calorimeter cosmics run 7061
  • 333969 10/08/10 23:15 A. Camsonne DVCS calorimeter cosmics run 7061
  • 333968 10/08/10 23:13 a-onl Start_Run_7061,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS cosmics test
  • 333967 10/08/10 23:09 a-onl End_of_Run_7060
  • 333966 10/08/10 23:06 A. Camsonne Calorimeter HV on for cosmics tests
  • 333965 10/08/10 23:05 a-onl Start_Run_7060,Run_type=Other,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test
  • 333964 10/08/10 23:02 a-onl Start_Run_7059,Run_type=Other,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test
  • 333963 10/08/10 22:58 a-onl End_of_Run_7057
  • 333962 10/08/10 22:52 A. Camsonne ARS Firmware E06 loaded
  • 333961 10/08/10 22:46 A. Camsonne FPGA programmers
  • 333960 10/08/10 21:53 riordan Sieve is OUT
  • 333959 10/08/10 21:47 riordan HRS to p0 = 4.050 GeV
  • 333958 10/08/10 21:07 Bo D2 pressure vs Loop1 temperature
  • 333957 10/08/10 20:40 jpchen Loop 1 is D2 and is already cold now.
  • 333956 10/08/10 19:34 Bo An alarm of D2 low Pressure
  • 333955 10/08/10 19:17 Bo screen of target temperature
  • 333954 10/08/10 18:43 auto_snap Radiation and Ozone snapshot
  • 333953 10/08/10 18:43 auto_snap Miscellaneous snapshot
  • 333952 10/08/10 18:43 auto_snap Hall A Tools snapshot
  • 333951 10/08/10 18:43 auto_snap Compton Polarimeter snapshot
  • 333950 10/08/10 18:43 auto_snap Beam Position and Current snapshot
  • 333949 10/08/10 18:03 A. Camsonne Laptop in HRS
  • 333948 10/08/10 17:57 adaq Beam Time Accounting
  • 333947 10/08/10 17:55 riordan Plans for Tonight
  • 333946 10/08/10 17:52 jpchen cryotarget status: waiting for Qweak target cool-down with uncertainty
  • 333945 10/08/10 17:06 a-onl End_of_Run_7057
  • 333944 10/08/10 17:04 jroche Left TSETTLE plugged in trigger 9
  • 333943 10/08/10 17:02 jroche helicity frequency test
  • 333942 10/08/10 16:55 a-onl Start_Run_7057,Run_type=Other,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=helicity trigger check, Helicity flipping rate at 240 Hz
  • 333941 10/08/10 16:54 a-onl End_of_Run_7056
  • 333940 10/08/10 16:47 a-onl Start_Run_7056,Run_type=Other,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=helicity trigger check, Helicity flipping rate at 658 Hz
  • 333939 10/08/10 16:32 a-onl Start_Run_7055,Run_type=Other,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=helicity trigger check
  • 333938 10/08/10 16:31 a-onl End_of_Run_7054
  • 333937 10/08/10 16:29 gomez Left HRS -- Q2/Q3 degaussed & momentum set to 3.5 GeV
  • 333936 10/08/10 16:25 a-onl Start_Run_7054,Run_type=Other,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=helicity trigger check
  • 333935 10/08/10 16:20 andrew/tech checklist
  • 333934 10/08/10 16:08 yez target screen -- loop1,loop2
  • 333933 10/08/10 16:07 yez target screen - main screen
  • 333932 10/08/10 16:06 yez target screen - temp,press. and flow
  • 333931 10/08/10 15:33 a-onl Start_Run_7053,Run_type=Other,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=helicity trigger check
  • 333930 10/08/10 15:32 jroche plugged a copy of TSETTLE in TS9
  • 333929 10/08/10 15:31 a-onl End_of_Run_7052
  • 333928 10/08/10 15:29 a-onl Start_Run_7052,Run_type=Other,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=helicity trigger check
  • 333927 10/08/10 15:28 a-onl End_of_Run_7051
  • 333926 10/08/10 15:27 a-onl Start_Run_7051,Run_type=Other,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=helicity trigger check
  • 333925 10/08/10 15:25 a-onl End_of_Run_7050
  • 333924 10/08/10 15:18 a-onl Start_Run_7050,Run_type=Other,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=helicity trigger check
  • 333923 10/08/10 15:15 a-onl End_of_Run_7049
  • 333922 10/08/10 15:13 a-onl Start_Run_7049,Run_type=Other,target_type=No Target, Empty,beam_e0=,beam_raster=on,beam_current=0,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=helicity trigger check
  • 333921 10/08/10 15:00 Carlos Calorimeter at 5m
  • 333920 10/08/10 11:34 O.Glamazdin Moller SC magnet has been filled
  • 333919 10/08/10 10:42 auto_snap Radiation and Ozone snapshot
  • 333918 10/08/10 10:42 auto_snap Miscellaneous snapshot
  • 333917 10/08/10 10:42 auto_snap Hall A Tools snapshot
  • 333916 10/08/10 10:42 auto_snap Compton Polarimeter snapshot
  • 333915 10/08/10 10:42 auto_snap Beam Position and Current snapshot
  • 333914 10/08/10 09:58 jpchen Qweak and Hall A target status and our target cooldown plan
  • 333913 10/08/10 08:05 adaq Beam Time Accounting
  • 333912 10/08/10 08:02 adaq Beam Time Accounting
  • 333911 10/08/10 07:44 Bellini Target report
  • 333910 10/08/10 02:42 auto_snap Radiation and Ozone snapshot
  • 333909 10/08/10 02:42 auto_snap Miscellaneous snapshot
  • 333908 10/08/10 02:42 auto_snap Hall A Tools snapshot
  • 333907 10/08/10 02:42 auto_snap Compton Polarimeter snapshot
  • 333906 10/08/10 02:42 auto_snap Beam Position and Current snapshot
  • 333905 10/08/10 00:31 Bellini screen snapshot
  • Thursday

  • 333904 10/07/10 23:49 Jin Huang Target Screen 6
  • 333903 10/07/10 23:19 Jin Huang Target Screen 5
  • 333902 10/07/10 21:35 Jin Huang Target Screen 4
  • 333901 10/07/10 21:06 riordan Run Plan for Swing 10/7 through Day 10/8, plans into next week
  • 333900 10/07/10 19:34 Jin Huang Target Screen 3
  • 333899 10/07/10 18:50 Jin Huang Target Screen 2.1
  • 333898 10/07/10 18:42 auto_snap Radiation and Ozone snapshot
  • 333897 10/07/10 18:42 auto_snap Miscellaneous snapshot
  • 333896 10/07/10 18:42 auto_snap Hall A Tools snapshot
  • 333895 10/07/10 18:42 auto_snap Compton Polarimeter snapshot
  • 333894 10/07/10 18:42 auto_snap Beam Position and Current snapshot
  • 333893 10/07/10 18:38 jpchen cryotargets status
  • 333892 10/07/10 18:34 Jin Huang Target Screen 2/2
  • 333891 10/07/10 18:34 Jin Huang Target Screen 1/2
  • 333890 10/07/10 17:16 J. Zhang Moller magnet is full
  • 333889 10/07/10 17:11 adaq Beam Time Accounting
  • 333888 10/07/10 16:13 M Friend Compton data from Oct 1st
  • 333887 10/07/10 16:12 Andrew/tech Checklist
  • 333886 10/07/10 16:03 hsmith Q1 magnet power supply
  • 333885 10/07/10 15:53 hsmith Q2 magnet power supply
  • 333884 10/07/10 15:04 adev Start_Run_20015,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=,beam_e0=0,beam_raster=on,beam_halfwaveplate=in,rightarm_p=0,rightarm_theta=-70,comment_text=cosmics
  • 333883 10/07/10 14:07 jpchen Reduced coolant flow/fan speeds
  • 333882 10/07/10 12:31 jpchen Hall C cryotarget cooldown started
  • 333881 10/07/10 10:35 A. Camsonne ARS test runs in DAQ room list
  • 333880 10/07/10 10:24 riordan Late Entry: Hall in Restricted Access
  • 333879 10/07/10 09:44 adaq Beam Time Accounting
  • 333878 10/07/10 09:36 yez cryotarget: yellow arlarm shows up-> haL1_hp_power_R
  • 333877 10/07/10 09:29 yez Cryotarget RED alarm shows up again: halTC502_1_Sensor_3_R
  • 333876 10/07/10 09:13 yez cryotarget screen 2
  • 333875 10/07/10 09:13 yez cryotarget screens
  • 333874 10/07/10 08:01 Bellini Errata Corrige
  • 333873 10/07/10 06:29 Carlos Owl shift summary
  • 333872 10/07/10 06:22 Carlos Dipole ramped down
  • 333871 10/07/10 06:17 Carlos Controlled access requested
  • 333870 10/07/10 06:12 Carlos Calorimeter HV reduced
  • 333869 10/07/10 06:10 Carlos Lumi tests
  • 333868 10/07/10 06:10 a-onl End_of_Run_7048
  • 333867 10/07/10 06:10 Carlos DC current with beam OFF
  • 333866 10/07/10 06:08 Carlos Lumi tests
  • 333865 10/07/10 06:05 Carlos PMT at 600V DC current
  • 333864 10/07/10 05:59 a-onl Start_Run_7048,Run_type=Spot++,target_type=LH2,beam_e0=5.562,beam_raster=on,beam_current=5,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster on 2x2, check
  • 333863 10/07/10 05:58 a-onl End_of_Run_7047
  • 333862 10/07/10 05:58 a-onl Start_Run_7047,Run_type=Spot++,target_type=LH2,beam_e0=5.562,beam_raster=on,beam_current=5,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster on 2x2, check
  • 333861 10/07/10 05:56 brads re: Raster 4x4 on C hole (shouldn't worry about non-centered hole)
  • 333860 10/07/10 05:44 Bellini Red Alarm On Target Loop 1
  • 333859 10/07/10 05:44 Carlos 7x6 raster on C hole
  • 333858 10/07/10 05:41 a-onl End_of_Run_7046
  • 333857 10/07/10 05:26 Carlos Raster 4x4 on C hole
  • 333856 10/07/10 05:25 a-onl Start_Run_7046,Run_type=Spot++,target_type=Cross hair,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster on 7x6, check
  • 333855 10/07/10 05:22 a-onl End_of_Run_7045
  • 333854 10/07/10 05:04 a-onl Start_Run_7045,Run_type=Spot++,target_type=Cross hair,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster on 4x4, check
  • 333853 10/07/10 05:03 a-onl End_of_Run_7044
  • 333852 10/07/10 04:58 a-onl Start_Run_7044,Run_type=Spot++,target_type=Cross hair,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster on 4x4, check
  • 333851 10/07/10 04:56 a-onl End_of_Run_7043
  • 333850 10/07/10 04:47 a-onl Start_Run_7043,Run_type=Spot++,target_type=Cross hair,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0.5,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=raster on 2x2, check
  • 333849 10/07/10 04:36 a-onl Start_Run_7042,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test top crate - Board 4 7 a 3 9 2 5
  • 333848 10/07/10 04:18 Carlos Target to Carbon hole
  • 333847 10/07/10 04:18 Carlos Moved target to BeO
  • 333846 10/07/10 04:05 riordan Ramping up Dipole to 200A
  • 333845 10/07/10 04:04 M Friend Compton tune not successful
  • 333844 10/07/10 04:00 Carlos Compton tuning stopped
  • 333843 10/07/10 03:41 Carlos Run 7041, event 23000
  • 333842 10/07/10 03:27 Carlos Calorimeter DAQ analysis
  • 333841 10/07/10 00:42 Carlos Beam back
  • 333840 10/07/10 00:15 adaq Beam Time Accounting
  • 333839 10/07/10 00:08 adaq Target Screen Shot
  • Wednesday

  • 333838 10/06/10 23:21 M Friend Compton tuning begun
  • 333837 10/06/10 21:39 M Friend Beam problems
  • 333836 10/06/10 21:10 A. Camsonne Issues to be checked from run 7041
  • 333835 10/06/10 20:15 M Friend Hall A energy stripchart
  • 333834 10/06/10 19:34 efuchey PbF2 Calorimeter HV Off
  • 333833 10/06/10 19:21 a-onl End_of_Run_7041
  • 333832 10/06/10 19:03 Solvignon Beam tuning starting
  • 333831 10/06/10 19:00 Jack (Techs) Magnet testing status
  • 333830 10/06/10 18:54 A. Camsonne Event Dump Run 7041 event 88
  • 333829 10/06/10 18:45 A. Camsonne ARS Cosmics run
  • 333828 10/06/10 18:33 a-onl Start_Run_7041,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test top crate - Board 4 7 a 3 9 2 5
  • 333827 10/06/10 18:32 a-onl End_of_Run_7040
  • 333826 10/06/10 18:27 a-onl Start_Run_7040,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test top crate - Board 4 7 a 3 9 2 5
  • 333825 10/06/10 18:22 a-onl End_of_Run_7039
  • 333824 10/06/10 18:14 a-onl Start_Run_7039,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test top crate - Board 5 9 2
  • 333823 10/06/10 18:09 a-onl Start_Run_7038,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test top crate - Board 5 9 2
  • 333822 10/06/10 17:52 A. Camsonne crl directory
  • 333821 10/06/10 17:49 a-onl Start_Run_7036,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test top crate
  • 333820 10/06/10 17:45 A. Camsonne HRS HV off for tests
  • 333819 10/06/10 17:26 adaq Beam Time Accounting
  • 333818 10/06/10 16:41 riordan Run Plan for Swing 10/6 through Day 10/7
  • 333817 10/06/10 16:27 adaq Beam Time Accounting
  • 333816 10/06/10 16:01 andrew/tech checklist
  • 333815 10/06/10 14:36 R. Michaels end-run scripts
  • 333814 10/06/10 14:26 a-onl End_of_Run_7035
  • 333813 10/06/10 14:18 a-onl End_of_Run_7035
  • 333812 10/06/10 14:15 yez cryo warning sound is gone and the presure is going back.
  • 333811 10/06/10 14:15 techs Q1
  • 333810 10/06/10 14:13 a-onl End_of_Run_7035
  • 333809 10/06/10 14:08 yez cryotarget low presure warning!
  • 333808 10/06/10 13:45 a-onl End_of_Run_7035
  • 333807 10/06/10 13:43 a-onl Start_Run_7035,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333806 10/06/10 13:39 O.Glamazdin Moller SC magnet has been filled
  • 333805 10/06/10 13:08 hsmith Q2 Over temperature faults
  • 333804 10/06/10 11:56 jpchen LD2 pressure
  • 333803 10/06/10 11:45 a-onl End_of_Run_7034
  • 333802 10/06/10 11:41 a-onl Start_Run_7034,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333801 10/06/10 11:28 a-onl End_of_Run_7033
  • 333800 10/06/10 11:26 a-onl Start_Run_7033,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333799 10/06/10 11:25 a-onl End_of_Run_7032
  • 333798 10/06/10 11:25 a-onl Start_Run_7032,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333797 10/06/10 11:25 adev End_of_Run_20014
  • 333796 10/06/10 11:23 a-onl Start_Run_7033,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333795 10/06/10 11:22 a-onl End_of_Run_7030
  • 333794 10/06/10 11:20 a-onl Start_Run_7030,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333793 10/06/10 11:14 yez L2Temps went high
  • 333792 10/06/10 11:09 a-onl Start_Run_7029,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333791 10/06/10 10:59 hyde Compton eDet noise
  • 333790 10/06/10 10:55 yez cryotarget screens-Loop1&Loop2
  • 333789 10/06/10 10:54 yez cryotarget control screen
  • 333788 10/06/10 10:49 a-onl End_of_Run_7029
  • 333787 10/06/10 10:47 a-onl Start_Run_7029,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333786 10/06/10 10:45 a-onl End_of_Run_7028
  • 333785 10/06/10 10:38 a-onl Start_Run_7028,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333784 10/06/10 10:36 a-onl End_of_Run_7026
  • 333783 10/06/10 10:30 a-onl Start_Run_7026,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333782 10/06/10 10:27 yez Cryotarget screen
  • 333781 10/06/10 09:46 yez hacsbc2 automatically rebooted
  • 333780 10/06/10 08:05 adaq Beam Time Accounting
  • 333779 10/06/10 02:42 auto_snap Radiation and Ozone snapshot
  • 333778 10/06/10 02:42 auto_snap Miscellaneous snapshot
  • 333777 10/06/10 02:41 auto_snap Hall A Tools snapshot
  • 333776 10/06/10 02:41 auto_snap Compton Polarimeter snapshot
  • 333775 10/06/10 02:41 auto_snap Beam Position and Current snapshot
  • 333774 10/06/10 01:02 a-onl Start_Run_7025,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333773 10/06/10 01:01 a-onl End_of_Run_7024
  • 333772 10/06/10 01:01 a-onl Start_Run_7024,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333771 10/06/10 00:58 a-onl End_of_Run_7023
  • 333770 10/06/10 00:56 a-onl Start_Run_7023,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333769 10/06/10 00:54 a-onl End_of_Run_7022
  • 333768 10/06/10 00:51 a-onl Start_Run_7022,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333767 10/06/10 00:44 a-onl End_of_Run_7021
  • 333766 10/06/10 00:42 a-onl Start_Run_7021,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333765 10/06/10 00:13 jroche, cshl BPMs have been calibrated
  • 333764 10/06/10 00:10 adaq cryotarget screens
  • 333763 10/06/10 00:06 jpchen cryotarget cooldown completed
  • Tuesday

  • 333762 10/05/10 23:26 a-onl End_of_Run_7020
  • 333761 10/05/10 23:22 a-onl Start_Run_7020,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333760 10/05/10 23:16 a-onl End_of_Run_7019
  • 333759 10/05/10 23:14 a-onl Start_Run_7019,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333758 10/05/10 23:03 a-onl End_of_Run_7018
  • 333757 10/05/10 22:39 a-onl Start_Run_7018,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333756 10/05/10 22:26 adaq Beam Time Accounting
  • 333755 10/05/10 22:02 a-onl Start_Run_7017,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333754 10/05/10 21:54 Magali Magne Timing on bus reduction for ARS
  • 333753 10/05/10 21:44 a-onl End_of_Run_7016
  • 333752 10/05/10 20:24 jpchen switched to cold-return
  • 333751 10/05/10 18:41 auto_snap Radiation and Ozone snapshot
  • 333750 10/05/10 18:41 auto_snap Miscellaneous snapshot
  • 333749 10/05/10 18:41 auto_snap Hall A Tools snapshot
  • 333748 10/05/10 18:41 auto_snap Compton Polarimeter snapshot
  • 333747 10/05/10 18:41 auto_snap Beam Position and Current snapshot
  • 333746 10/05/10 18:19 M Friend Compton FADC DAQ clock change
  • 333745 10/05/10 18:06 jroche Pedestals for s1,s2,cer and pir updated
  • 333744 10/05/10 17:37 jpchen J-T valve for loop 3
  • 333743 10/05/10 16:45 R. Michaels epicsLogger script split in two
  • 333742 10/05/10 16:38 a-onl Start_Run_7016,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333741 10/05/10 16:23 riordan Update to run plan
  • 333740 10/05/10 16:18 jpchen Re-start target cool-down
  • 333738 10/05/10 15:50 a-onl End_of_Run_7015
  • 333737 10/05/10 15:44 a-onl Start_Run_7015,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics - test 2 crates - Some error messages at init
  • 333736 10/05/10 15:41 a-onl End_of_Run_7013
  • 333735 10/05/10 15:40 Carlos Calorimeter anode current check
  • 333734 10/05/10 15:31 J. Zhang Moller solenoid has been refilled
  • 333733 10/05/10 15:22 a-onl End_of_Run_7013
  • 333732 10/05/10 14:42 jroche Calorimeter DC monitor GUI now reliable
  • 333731 10/05/10 12:51 A. Camsonne Top crate first registers read comparison
  • 333730 10/05/10 12:51 A. Camsonne Top crate first registers read comparison
  • 333729 10/05/10 12:31 a-onl Start_Run_7012,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics
  • 333728 10/05/10 12:29 a-onl End_of_Run_7011
  • 333727 10/05/10 12:28 a-onl Start_Run_7011,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Calorimeter cosmics
  • 333726 10/05/10 12:21 a-onl End_of_Run_7010
  • 333725 10/05/10 12:15 A. Camsonne DVCS calorimeter HV back on
  • 333724 10/05/10 12:15 hisham ADC synch checker cables are installed
  • 333723 10/05/10 12:11 a-onl Start_Run_7010,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test - pulser trigger - Test alignement - All boards top crate 59247a3
  • 333722 10/05/10 12:00 a-onl End_of_Run_7009
  • 333721 10/05/10 12:00 jpchen cryotarget cooldown status
  • 333720 10/05/10 11:58 a-onl Start_Run_7009,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test - pulser trigger - Test alignement - All boards top crate 59247a3
  • 333719 10/05/10 11:54 a-onl End_of_Run_7008
  • 333718 10/05/10 11:54 a-onl Start_Run_7008,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test - pulser trigger - Test alignement
  • 333717 10/05/10 11:49 a-onl Start_Run_7000,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test
  • 333716 10/05/10 11:49 a-onl End_of_Run_7007
  • 333715 10/05/10 11:46 a-onl End_of_Run_7006
  • 333714 10/05/10 11:39 A. Camsonne ARS Firmware D26 loaded
  • 333713 10/05/10 11:39 riordan Run Plan for Swing 10/5 through Day 10/6
  • 333712 10/05/10 11:33 a-onl Start_Run_7006,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test - pulser trigger
  • 333711 10/05/10 11:30 a-onl End_of_Run_7005
  • 333710 10/05/10 11:30 a-onl Start_Run_7005,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test
  • 333709 10/05/10 11:28 a-onl End_of_Run_7004
  • 333708 10/05/10 11:23 M Friend Moved Compton FADC DAQ BCM
  • 333707 10/05/10 11:20 a-onl Start_Run_7004,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test
  • 333706 10/05/10 11:20 hyde HRS Left Cerenkov Threshold
  • 333705 10/05/10 11:10 a-onl Start_Run_7003,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test
  • 333704 10/05/10 11:05 a-onl End_of_Run_7003
  • 333703 10/05/10 11:03 a-onl Start_Run_7003,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test
  • 333702 10/05/10 11:02 a-onl End_of_Run_7002
  • 333701 10/05/10 11:01 a-onl Start_Run_7002,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test
  • 333700 10/05/10 10:54 a-onl End_of_Run_7001
  • 333699 10/05/10 10:53 a-onl Start_Run_7001,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=ARS test
  • 333698 10/05/10 10:51 a-onl End_of_Run_7000
  • 333697 10/05/10 10:41 auto_snap Radiation and Ozone snapshot
  • 333696 10/05/10 10:41 auto_snap Miscellaneous snapshot
  • 333695 10/05/10 10:41 auto_snap Hall A Tools snapshot
  • 333694 10/05/10 10:41 auto_snap Compton Polarimeter snapshot
  • 333693 10/05/10 10:41 auto_snap Beam Position and Current snapshot
  • 333692 10/05/10 09:55 Hisham Helicity cables
  • 333691 10/05/10 09:34 adaq Beam Time Accounting
  • 333690 10/05/10 09:22 adev Start_Run_20014,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=,beam_e0=0,beam_raster=on,beam_halfwaveplate=in,rightarm_p=0,rightarm_theta=-70,comment_text=cosmics
  • 333689 10/05/10 09:16 adev End_of_Run_20013
  • 333688 10/05/10 06:04 riordan Requested Restricted Access
  • 333687 10/05/10 04:01 Hisham Owl shift summary
  • 333686 10/05/10 04:00 a-onl Start_Run_6990,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Anode current epics test
  • 333685 10/05/10 03:59 Carlos BPM calibration (cont)
  • 333684 10/05/10 03:44 a-onl Start_Run_6999,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration Beam Position 0 0 mm
  • 333683 10/05/10 03:44 a-onl End_of_Run_6998
  • 333682 10/05/10 03:33 a-onl Start_Run_6998,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration Beam Position 0 0 mm
  • 333681 10/05/10 03:23 a-onl End_of_Run_6997
  • 333680 10/05/10 03:05 a-onl Start_Run_6997,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration Beam Position 0 -3 mm
  • 333679 10/05/10 02:52 a-onl End_of_Run_6996
  • 333678 10/05/10 02:41 auto_snap Radiation and Ozone snapshot
  • 333677 10/05/10 02:41 auto_snap Miscellaneous snapshot
  • 333676 10/05/10 02:41 auto_snap Hall A Tools snapshot
  • 333675 10/05/10 02:40 auto_snap Compton Polarimeter snapshot
  • 333674 10/05/10 02:40 auto_snap Beam Position and Current snapshot
  • 333673 10/05/10 02:36 a-onl Start_Run_6996,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration Beam Position -3 0 mm
  • 333672 10/05/10 02:22 a-onl End_of_Run_6995
  • 333671 10/05/10 02:04 a-onl Start_Run_6995,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration Beam Position +3 0 mm
  • 333670 10/05/10 01:45 a-onl End_of_Run_6994
  • 333669 10/05/10 01:42 Carlos Calorimeter anode current
  • 333668 10/05/10 01:32 a-onl Start_Run_6994,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration Beam Position 0 +3 mm
  • 333667 10/05/10 01:13 a-onl End_of_Run_6993
  • 333666 10/05/10 00:58 a-onl Start_Run_6993,Run_type=Check out,target_type=Optics: multi 1mm C foils,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration Beam Position -3 +3 mm
  • 333665 10/05/10 00:18 Carlos BPM calibration (cont)
  • 333664 10/05/10 00:01 adaq Beam Time Accounting
  • Monday

  • 333663 10/04/10 23:53 Jin Huang Shift Summary : Oct 04 Swing
  • 333662 10/04/10 23:52 a-onl End_of_Run_6992
  • 333661 10/04/10 23:52 Jin Huang Target -> empty
  • 333660 10/04/10 23:39 a-onl Start_Run_6992,Run_type=Check out,target_type=Optics: multi 1mm C foils,beam_e0=5.562,beam_raster=off,beam_current=2,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Kalyan's test run, Trigger on T8 (ps8=1)
  • 333659 10/04/10 23:36 a-onl End_of_Run_6991
  • 333658 10/04/10 23:29 Jin Huang Moving -> Optics target
  • 333657 10/04/10 23:18 Jin Huang Moving to BeO
  • 333656 10/04/10 23:15 a-onl Start_Run_6991,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Anode current epics test
  • 333655 10/04/10 23:12 Jin Huang Screen shot: Tune beam in hall
  • 333654 10/04/10 23:11 a-onl End_of_Run_6990
  • 333653 10/04/10 23:11 a-onl End_of_Run_6989
  • 333652 10/04/10 22:54 a-onl Start_Run_6989,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Anode current epics test
  • 333651 10/04/10 22:14 Jin Huang Tuning beam into hall
  • 333650 10/04/10 22:06 R. Michaels EPICS truncation due to ET insertions
  • 333649 10/04/10 21:44 R. Michaels re: EPICS problem : how to know if it's a problem with Podd or not
  • 333648 10/04/10 21:10 Jin Huang Update: Hall remain Controled Access due to injector issues
  • 333647 10/04/10 21:02 Jin Huang Sweep finished
  • 333646 10/04/10 20:12 Carlos Calorimeter HV off
  • 333645 10/04/10 20:01 Carlos Hall A locking up in 15min
  • 333644 10/04/10 19:56 yez L-HRS Pion Rejectors Plots for cosmic ray run6988
  • 333643 10/04/10 19:53 yez L-HRS Gas Cerenkov Plots for cosmic ray run6988
  • 333642 10/04/10 19:52 yez L-HRS Scin Plots for cosmic ray run6988 (Same plots, but with names)
  • 333641 10/04/10 19:51 meekins Target Pump Purge
  • 333640 10/04/10 19:48 yez L-HRS Scin Plots for cosmic ray run6988
  • 333639 10/04/10 19:36 yez L-HRS VDC plots for cosmic ray run6988
  • 333638 10/04/10 18:40 auto_snap Radiation and Ozone snapshot
  • 333637 10/04/10 18:40 auto_snap Miscellaneous snapshot
  • 333636 10/04/10 18:40 auto_snap Hall A Tools snapshot
  • 333635 10/04/10 18:40 auto_snap Compton Polarimeter snapshot
  • 333634 10/04/10 18:40 auto_snap Beam Position and Current snapshot
  • 333633 10/04/10 18:32 jroche Pedestals need adjusting: will do that tomorrow
  • 333632 10/04/10 18:01 Carlos EPICS truncation (cont)
  • 333631 10/04/10 17:33 riordan Run Plan for Swing 10/4 through Day 10/5
  • 333630 10/04/10 17:26 a-onl Start_Run_6988,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=HRS cosmics runs
  • 333629 10/04/10 17:24 a-onl End_of_Run_6987
  • 333628 10/04/10 16:18 Kalyan lumi channel 7 and 8
  • 333627 10/04/10 16:09 adaq Beam Time Accounting
  • 333626 10/04/10 15:59 A. Camsonne ARS cosmics run 6987 firmware D24
  • 333625 10/04/10 15:46 A. Camsonne Signal tap - second file working
  • 333624 10/04/10 15:43 a-onl Start_Run_6987,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test ARS calorimeter - cosmics
  • 333623 10/04/10 15:42 andrewl Cryo target pre beam checklist
  • 333622 10/04/10 15:38 a-onl End_of_Run_6986
  • 333621 10/04/10 15:36 A. Camsonne DVCS calorimeter cosmics scintillators paddle HV on
  • 333620 10/04/10 15:29 R. Michaels new helicity patterns (thanks, Qweak)
  • 333619 10/04/10 15:24 a-onl Start_Run_6986,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test ARS calorimeter - cosmics
  • 333618 10/04/10 15:22 a-onl End_of_Run_6985
  • 333617 10/04/10 15:21 a-onl Start_Run_6985,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test ARS calorimeter
  • 333616 10/04/10 15:18 A. Camsonne HRS HV off for DAQ test
  • 333615 10/04/10 15:16 a-onl End_of_Run_6984
  • 333614 10/04/10 15:16 a-onl Start_Run_6983,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test ARS calorimeter
  • 333612 10/04/10 15:07 a-onl End_of_Run_6983
  • 333611 10/04/10 14:52 hisham Helicity signals
  • 333610 10/04/10 14:50 A. Camsonne Calorimeter HV on for cosmics tests
  • 333609 10/04/10 14:02 R. Michaels ADC synch checker installed
  • 333608 10/04/10 13:37 a-onl Start_Run_6982,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=
  • 333607 10/04/10 12:32 a-onl End_of_Run_6981
  • 333606 10/04/10 12:15 a-onl Start_Run_6981,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test, HRS (calo)T5 only
  • 333605 10/04/10 12:14 a-onl End_of_Run_6980
  • 333604 10/04/10 12:13 a-onl Start_Run_6980,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test, HRS T4 only
  • 333603 10/04/10 12:12 a-onl End_of_Run_6979
  • 333602 10/04/10 11:37 a-onl Start_Run_6979,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test, HRS T3 only
  • 333601 10/04/10 11:37 a-onl End_of_Run_6978
  • 333600 10/04/10 11:34 a-onl Start_Run_6978,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test, HRS T2 only
  • 333599 10/04/10 11:32 a-onl End_of_Run_6977
  • 333598 10/04/10 11:30 a-onl Start_Run_6977,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test, HRS T1 only
  • 333597 10/04/10 10:58 riordan Hall in Restricted Access
  • 333596 10/04/10 10:40 auto_snap Radiation and Ozone snapshot
  • 333595 10/04/10 10:40 auto_snap Miscellaneous snapshot
  • 333594 10/04/10 10:40 auto_snap Hall A Tools snapshot
  • 333593 10/04/10 10:40 auto_snap Compton Polarimeter snapshot
  • 333592 10/04/10 10:40 auto_snap Beam Position and Current snapshot
  • 333591 10/04/10 09:47 jmammei Day Shift Summary
  • 333590 10/04/10 09:35 adev Start_Run_20013,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=,beam_e0=0,beam_raster=on,beam_halfwaveplate=in,rightarm_p=0,rightarm_theta=-70,comment_text=cosmics
  • 333589 10/04/10 09:35 adev End_of_Run_20012
  • 333588 10/04/10 08:16 adaq Beam Time Accounting
  • 333587 10/04/10 06:57 Carlos Owl shift summary
  • 333586 10/04/10 06:46 Carlos 2uA on optics target
  • 333585 10/04/10 06:43 Carlos MCC will run multi-harp scan
  • 333584 10/04/10 06:09 a-onl End_of_Run_6976
  • 333583 10/04/10 06:07 Carlos Target back to empty
  • 333582 10/04/10 06:06 a-onl Start_Run_6976,Run_type=Calibration,target_type=Optics: multi 1mm C foils,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test
  • 333581 10/04/10 05:58 Carlos Move target to optics
  • 333580 10/04/10 05:55 a-onl End_of_Run_6975
  • 333579 10/04/10 05:53 Carlos Raster 7x5
  • 333578 10/04/10 05:53 a-onl Start_Run_6975,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Raster Calibration: 7x5 @ (0,0)
  • 333577 10/04/10 05:51 Carlos Raster calibration
  • 333576 10/04/10 05:47 a-onl End_of_Run_6974
  • 333575 10/04/10 05:46 a-onl Start_Run_6974,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Raster Calibration: 3x2 @ (0,0)
  • 333574 10/04/10 05:45 Carlos 3x2 raster
  • 333573 10/04/10 05:43 a-onl End_of_Run_6973
  • 333572 10/04/10 05:42 a-onl Start_Run_6973,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Raster Calibration: 3x2 @ (0,0)
  • 333571 10/04/10 05:41 a-onl End_of_Run_6972
  • 333570 10/04/10 05:27 a-onl Start_Run_6972,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Raster Calibration: 3x2 @ (0,0)
  • 333569 10/04/10 05:20 adaq Beam Time Accounting
  • 333568 10/04/10 04:43 Carlos Raster calibration
  • 333567 10/04/10 04:43 Carlos BPM calibration runlist
  • 333566 10/04/10 04:42 Carlos BPM calibration aborted
  • 333565 10/04/10 04:30 a-onl End_of_Run_6971
  • 333564 10/04/10 04:14 a-onl Start_Run_6971,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration: pt (0,+3)
  • 333563 10/04/10 03:58 a-onl End_of_Run_6970
  • 333562 10/04/10 03:39 a-onl Start_Run_6970,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration: pt (-3,-3)
  • 333561 10/04/10 02:53 a-onl End_of_Run_6969
  • 333560 10/04/10 02:40 auto_snap Radiation and Ozone snapshot
  • 333559 10/04/10 02:40 auto_snap Miscellaneous snapshot
  • 333558 10/04/10 02:40 auto_snap Hall A Tools snapshot
  • 333557 10/04/10 02:40 auto_snap Compton Polarimeter snapshot
  • 333556 10/04/10 02:39 auto_snap Beam Position and Current snapshot
  • 333555 10/04/10 02:36 a-onl Start_Run_6969,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration: pt (-3,-3)
  • 333554 10/04/10 02:04 a-onl End_of_Run_6968
  • 333553 10/04/10 01:49 a-onl Start_Run_6968,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration: pt (+3,-3)
  • 333552 10/04/10 01:48 a-onl End_of_Run_6967
  • 333551 10/04/10 01:47 a-onl Start_Run_6967,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration: pt (+3,-3)
  • 333550 10/04/10 01:39 a-onl End_of_Run_6966
  • 333549 10/04/10 01:13 a-onl End_of_Run_6965
  • 333548 10/04/10 00:55 a-onl Start_Run_6965,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration: pt (+3,+3)
  • 333547 10/04/10 00:41 a-onl End_of_Run_6964
  • 333546 10/04/10 00:24 a-onl Start_Run_6964,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration: pt (0,0)
  • 333545 10/04/10 00:18 a-onl End_of_Run_6963
  • 333544 10/04/10 00:15 a-onl Start_Run_6963,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=5,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM Calibration: pt (0,0)
  • 333543 10/04/10 00:13 a-onl End_of_Run_6962
  • 333542 10/04/10 00:02 efuchey compton rates for tonight
  • Sunday

  • 333540 10/03/10 23:58 Carlos Compton rates still extremely high
  • 333538 10/03/10 23:21 a-onl Start_Run_6962,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test epics readout
  • 333537 10/03/10 23:14 Carlos Moller commissioning finished
  • 333536 10/03/10 23:07 O.Glamazdin Moller results
  • 333535 10/03/10 23:03 O.Glamazdin Moller results
  • 333534 10/03/10 21:33 O.Glamazdin Moller solenoid quench
  • 333533 10/03/10 19:24 jroche bpms are working: need harps scan for calibration
  • 333532 10/03/10 18:56 adaq Beam Time Accounting
  • 333531 10/03/10 18:39 auto_snap Radiation and Ozone snapshot
  • 333530 10/03/10 18:39 auto_snap Miscellaneous snapshot
  • 333529 10/03/10 18:39 auto_snap Hall A Tools snapshot
  • 333528 10/03/10 18:39 auto_snap Compton Polarimeter snapshot
  • 333527 10/03/10 18:39 auto_snap Beam Position and Current snapshot
  • 333526 10/03/10 17:18 riordan Run Plan for Swing 10/3 through Day 10/4
  • 333525 10/03/10 16:21 a-onl Start_Run_6961,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test D24 - All prescale
  • 333524 10/03/10 16:13 a-onl End_of_Run_6960
  • 333523 10/03/10 16:02 a-onl Start_Run_6960,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test D24
  • 333522 10/03/10 16:00 a-onl End_of_Run_6959
  • 333521 10/03/10 15:58 a-onl Start_Run_6959,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Test D24
  • 333520 10/03/10 15:48 a-onl End_of_Run_6958
  • 333519 10/03/10 15:38 a-onl Start_Run_6958,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=run 5 uA - changing beam positions to BPMB 2.4, -2.4 BPMA (2.,-2)
  • 333518 10/03/10 15:34 a-onl End_of_Run_6957
  • 333517 10/03/10 15:25 a-onl Start_Run_6957,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=run 10 uA - changing beam positions to BPMB 2.5, 3.24 BPMA (2.1,1.6)
  • 333516 10/03/10 15:23 A. Camsonne Rough BCM calibration
  • 333515 10/03/10 15:22 a-onl End_of_Run_6956
  • 333514 10/03/10 15:20 a-onl Start_Run_6956,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=junk
  • 333513 10/03/10 15:17 a-onl End_of_Run_6955
  • 333512 10/03/10 15:10 a-onl Start_Run_6955,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Beam run 10 uA - changing beam positions to BPMB 0,-2.5 BPMA (-2,-2)
  • 333511 10/03/10 15:09 a-onl End_of_Run_6954
  • 333510 10/03/10 14:44 a-onl Start_Run_6954,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Beam run 5 uA - changing beam positions to BPMB 0,-2.5 BPMA (-2,-2)
  • 333509 10/03/10 14:23 a-onl End_of_Run_6953
  • 333508 10/03/10 14:18 a-onl Start_Run_6953,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Beam run 5 uA - changing beam positions to -2,-2 and then +2,+2 (roughly)
  • 333507 10/03/10 14:11 a-onl End_of_Run_6952
  • 333506 10/03/10 14:07 a-onl Start_Run_6952,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Beam run 5 uA - EPICS position BPM4A (-0.29,0.24) BPM4B (-0.1,0.12)
  • 333505 10/03/10 14:05 A. Camsonne Spot BPM pedestals updated
  • 333504 10/03/10 14:00 A. Camsonne Spot Left DB backed up
  • 333503 10/03/10 13:53 A. Camsonne BPM pedestals from get_bpm_pedestals.C
  • 333502 10/03/10 13:49 a-onl End_of_Run_6951
  • 333501 10/03/10 13:48 a-onl Start_Run_6951,Run_type=Calibration,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM pedestal run - Gains set to 0 for BPM 4A and 4B
  • 333500 10/03/10 13:46 A. Camsonne BPM pedestal runs 6950,6951
  • 333499 10/03/10 13:45 a-onl Start_Run_6950,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=BPM pedestal run
  • 333498 10/03/10 13:36 A. Camsonne Requested BPM gains set to 0 for BPM pedestal run
  • 333497 10/03/10 13:21 a-onl Start_Run_6948,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=test
  • 333496 10/03/10 13:18 a-onl Start_Run_6947,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=test
  • 333495 10/03/10 13:11 Magali aquisition rack status
  • 333494 10/03/10 12:51 A. Camsonne BPM offset and attenuation adjusted
  • 333493 10/03/10 12:49 A. Camsonne BPM offset and attenuation adjusted
  • 333492 10/03/10 12:30 benesch EES checked out their piece of the 4A/4B BPMs
  • 333491 10/03/10 10:57 Fansler Cryo
  • 333490 10/03/10 10:39 auto_snap Radiation and Ozone snapshot
  • 333489 10/03/10 10:39 auto_snap Miscellaneous snapshot
  • 333488 10/03/10 10:39 auto_snap Hall A Tools snapshot
  • 333487 10/03/10 10:39 auto_snap Compton Polarimeter snapshot
  • 333486 10/03/10 10:39 auto_snap Beam Position and Current snapshot
  • 333485 10/03/10 09:44 adaq Beam Time Accounting
  • 333484 10/03/10 06:10 Carlos Going to controlled access
  • 333483 10/03/10 05:47 Carlos Owl shift summary
  • 333482 10/03/10 05:43 a-onl End_of_Run_6945
  • 333481 10/03/10 05:42 a-onl Start_Run_6945,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333480 10/03/10 05:39 a-onl End_of_Run_6944
  • 333479 10/03/10 05:39 Carlos Done for tonight
  • 333478 10/03/10 05:38 a-onl Start_Run_6944,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333477 10/03/10 05:38 Carlos Calorimeter anode current study
  • 333476 10/03/10 05:37 D.Wang target moved to Empty
  • 333475 10/03/10 05:35 a-onl Start_Run_6943,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333474 10/03/10 05:32 Carlos Calorimeter with HV off
  • 333473 10/03/10 05:30 Carlos Calorimeter 3uA on optics target
  • 333472 10/03/10 05:26 a-onl End_of_Run_6942
  • 333471 10/03/10 05:24 a-onl Start_Run_6942,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333470 10/03/10 05:23 D.Wang Target moved to Optics
  • 333469 10/03/10 05:22 Carlos 1 uA on optics target
  • 333468 10/03/10 05:21 Carlos Optics target in
  • 333467 10/03/10 05:19 a-onl End_of_Run_6941
  • 333466 10/03/10 05:16 a-onl Start_Run_6941,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333465 10/03/10 05:14 Carlos Calorimeter at 2uA in C
  • 333464 10/03/10 05:09 Carlos Calorimeter at 1uA on C
  • 333463 10/03/10 05:01 Carlos Carbon target in
  • 333462 10/03/10 04:53 Carlos BPM studies
  • 333461 10/03/10 04:49 Carlos BPMs at 10uA
  • 333460 10/03/10 04:46 a-onl Start_Run_6939,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333459 10/03/10 04:37 Carlos BPM signals at 2uA
  • 333458 10/03/10 04:09 a-onl Start_Run_6937,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333457 10/03/10 04:08 a-onl End_of_Run_6936
  • 333456 10/03/10 04:02 a-onl Start_Run_6936,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333455 10/03/10 04:01 a-onl End_of_Run_6935
  • 333454 10/03/10 03:51 a-onl Start_Run_6935,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333453 10/03/10 03:43 a-onl Start_Run_6934,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333452 10/03/10 03:43 a-onl End_of_Run_6933
  • 333451 10/03/10 03:26 a-onl Start_Run_6933,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333450 10/03/10 03:26 a-onl End_of_Run_6932
  • 333449 10/03/10 03:25 a-onl End_of_Run_6919
  • 333448 10/03/10 03:19 a-onl Start_Run_6932,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333447 10/03/10 03:18 a-onl End_of_Run_6931
  • 333446 10/03/10 03:15 Carlos BPM working at extremely low current
  • 333445 10/03/10 03:07 a-onl Start_Run_6931,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333444 10/03/10 03:07 a-onl End_of_Run_6930
  • 333443 10/03/10 02:58 a-onl Start_Run_6930,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333442 10/03/10 02:58 a-onl End_of_Run_6929
  • 333441 10/03/10 02:52 Carlos Compton studies aborted
  • 333440 10/03/10 02:52 a-onl Start_Run_6929,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333439 10/03/10 02:51 a-onl End_of_Run_6928
  • 333438 10/03/10 02:51 a-onl End_of_Run_6928
  • 333437 10/03/10 02:45 efuchey Compton Rates
  • 333436 10/03/10 02:39 auto_snap Radiation and Ozone snapshot
  • 333435 10/03/10 02:39 auto_snap Miscellaneous snapshot
  • 333434 10/03/10 02:39 auto_snap Hall A Tools snapshot
  • 333433 10/03/10 02:39 auto_snap Compton Polarimeter snapshot
  • 333432 10/03/10 02:39 auto_snap Beam Position and Current snapshot
  • 333431 10/03/10 02:36 a-onl Start_Run_6928,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333430 10/03/10 02:35 a-onl End_of_Run_6927
  • 333429 10/03/10 02:32 Carlos Beam trips
  • 333428 10/03/10 02:18 a-onl Start_Run_6927,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333427 10/03/10 02:18 a-onl End_of_Run_6926
  • 333426 10/03/10 01:59 a-onl Start_Run_6926,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333425 10/03/10 01:58 a-onl End_of_Run_6925
  • 333424 10/03/10 01:45 a-onl Start_Run_6925,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333423 10/03/10 01:44 a-onl End_of_Run_6924
  • 333422 10/03/10 01:33 Carlos Beam restoration...
  • 333421 10/03/10 01:08 a-onl Start_Run_6924,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333420 10/03/10 01:08 a-onl End_of_Run_6923
  • 333419 10/03/10 00:54 a-onl Start_Run_6923,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333418 10/03/10 00:54 a-onl End_of_Run_6922
  • 333417 10/03/10 00:53 efuchey Compton Rates too high tonight
  • 333416 10/03/10 00:33 a-onl Start_Run_6922,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333415 10/03/10 00:33 a-onl End_of_Run_6921
  • 333413 10/03/10 00:24 a-onl Start_Run_6921,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333414 10/03/10 00:24 adaq Beam Time Accounting
  • 333412 10/03/10 00:23 a-onl End_of_Run_6919
  • Saturday

  • 333411 10/02/10 23:39 a-onl Start_Run_6919,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333410 10/02/10 23:36 a-onl End_of_Run_6918
  • 333409 10/02/10 23:20 a-onl Start_Run_6918,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333408 10/02/10 23:19 a-onl End_of_Run_6917
  • 333407 10/02/10 23:18 a-onl Start_Run_6917,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333406 10/02/10 23:09 a-onl End_of_Run_6916
  • 333405 10/02/10 23:04 O.Glamazdin Moller commissioning
  • 333404 10/02/10 23:00 a-onl Start_Run_6916,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only) - Check on flag
  • 333403 10/02/10 22:35 a-onl End_of_Run_6915
  • 333402 10/02/10 22:26 a-onl Start_Run_6915,Run_type=Other,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D23 ROC17 592 (T5 only)
  • 333401 10/02/10 22:24 Roman FADC Data Trigger Time Distribution
  • 333400 10/02/10 22:19 a-onl End_of_Run_6914
  • 333398 10/02/10 20:40 riordan Shift Update
  • 333397 10/02/10 18:40 Carlos Work done during controlled access
  • 333396 10/02/10 18:38 auto_snap Radiation and Ozone snapshot
  • 333395 10/02/10 18:38 auto_snap Miscellaneous snapshot
  • 333394 10/02/10 18:38 auto_snap Hall A Tools snapshot
  • 333393 10/02/10 18:38 auto_snap Compton Polarimeter snapshot
  • 333392 10/02/10 18:38 auto_snap Beam Position and Current snapshot
  • 333391 10/02/10 16:27 riordan Run Plan for Swing 10/2 - Day 10/3
  • 333390 10/02/10 16:16 adaq Beam Time Accounting
  • 333389 10/02/10 16:11 a-onl Start_Run_6914,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D22 ROC17 592 (T5 only)
  • 333388 10/02/10 16:08 a-onl Start_Run_6913,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=D22 ROC17 592
  • 333387 10/02/10 16:06 A. Camsonne Run 6914 Top crate 592 D22
  • 333386 10/02/10 16:03 a-onl End_of_Run_6912
  • 333385 10/02/10 15:57 A. Camsonne ARS board 592 firmware D22 loaded
  • 333384 10/02/10 15:56 a-onl Start_Run_6912,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Testing BPM
  • 333383 10/02/10 15:55 a-onl End_of_Run_6911
  • 333382 10/02/10 15:43 a-onl Start_Run_6911,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Testing BPM
  • 333381 10/02/10 15:00 riordan beam permit for Moller
  • 333380 10/02/10 14:25 fansler LQ2
  • 333379 10/02/10 13:42 Fansler Cryo
  • 333378 10/02/10 12:59 a-onl End_of_Run_6910
  • 333377 10/02/10 12:22 a-onl Start_Run_6910,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Testing BPM gate
  • 333376 10/02/10 12:22 a-onl End_of_Run_6909
  • 333375 10/02/10 12:10 fansler Moeller
  • 333374 10/02/10 11:29 a-onl Start_Run_6909,Run_type=Check out,target_type=No Target, Empty,beam_e0=5.562,beam_raster=off,beam_current=0,leftarm_p=0,leftarm_theta=21.5,leftarm_sieve=out,calo_theta=22,calo_distance=5,comment_text=Testing BPM gate
  • 333373 10/02/10 11:27 a-onl End_of_Run_6902
  • 333372 10/02/10 10:59 adaq Beam Time Accounting
  • 333371 10/02/10 10:58 adaq Beam Time Accounting
  • 333370 10/02/10 10:38 auto_snap Radiation and Ozone snapshot
  • 333369 10/02/10 10:38 auto_snap Miscellaneous snapshot
  • 333368 10/02/10 10:38 auto_snap Hall A Tools snapshot
  • 333367 10/02/10 10:38 auto_snap Compton Polarimeter snapshot
  • 333366 10/02/10 10:38 auto_snap Beam Position and Current snapshot
  • 333365 10/02/10 10:22 Fansler Moeller N2
  • 333364 10/02/10 10:18 fansler Cryo
  • 333363 10/02/10 02:38 auto_snap Radiation and Ozone snapshot
  • 333362 10/02/10 02:38 auto_snap Miscellaneous snapshot
  • 333361 10/02/10 02:38 auto_snap Hall A Tools snapshot
  • 333360 10/02/10 02:38 auto_snap Compton Polarimeter snapshot
  • 333359 10/02/10 02:38 auto_snap Beam Position and Current snapshot
  • 333358 10/02/10 00:37 riordan Requested Controlled Access
  • 333357 10/02/10 00:09 Xiaohui Zhan Xiaohui Zhan
  • Friday

  • 333356 10/01/10 23:24 efuchey Compton Scanning
  • 333355 10/01/10 22:37 Carlos BMP wires 3 and 4 showing wrong readings
  • 333354 10/01/10 21:23 Kalyan Lumi test runs 6908 6907
  • 333353 10/01/10 20:55 a-onl End_of_Run_6908
  • 333352 10/01/10 20:51 Kalyan lumi test runs
  • 333351 10/01/10 20:48 Kalyan Target empty
  • 333350 10/01/10 20:39 a-onl Start_Run_6908,Run_type=Optics,target_type=C12 1mm foil,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=lumi test 2uA on optics target
  • 333349 10/01/10 20:38 a-onl End_of_Run_6907
  • 333348 10/01/10 20:33 Kalyan Lumi high voltage screen shot
  • 333347 10/01/10 20:31 Kalyan Target move to optics
  • 333346 10/01/10 20:30 Kalyan target moved to carbon
  • 333345 10/01/10 20:10 a-onl Start_Run_6907,Run_type=Optics,target_type=C12 1mm foil,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=lumi test 2uA on carbon target
  • 333344 10/01/10 20:08 a-onl End_of_Run_6906
  • 333343 10/01/10 19:52 A. Camsonne One missing histogram for pion rejector
  • 333342 10/01/10 19:51 riordan BeO Spot on Target for BPM (0,0)
  • 333341 10/01/10 19:47 Carlos BPM readout problem
  • 333340 10/01/10 19:46 a-onl Start_Run_6906,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Cosmics run T3 and T1 for spot++ (4x4 raster)
  • 333339 10/01/10 19:45 a-onl End_of_Run_6905
  • 333338 10/01/10 19:45 a-onl Start_Run_6905,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Cosmics run T3 and T1 for spot++
  • 333337 10/01/10 19:45 riordan Target to Single Carbon Foil
  • 333336 10/01/10 19:44 a-onl End_of_Run_6904
  • 333335 10/01/10 19:43 a-onl Start_Run_6904,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Cosmics run T3 and T1 for spot++
  • 333334 10/01/10 19:42 a-onl End_of_Run_6903
  • 333333 10/01/10 19:40 a-onl Start_Run_6903,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Cosmics run T3 and T1 for spot++
  • 333332 10/01/10 19:39 riordan Target to BeO
  • 333331 10/01/10 19:24 a-onl Start_Run_6902,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Cosmics run T3
  • 333330 10/01/10 19:23 a-onl End_of_Run_6901
  • 333329 10/01/10 18:49 a-onl Start_Run_6901,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=test
  • 333328 10/01/10 18:44 a-onl End_of_Run_6900
  • 333327 10/01/10 18:38 auto_snap Radiation and Ozone snapshot
  • 333326 10/01/10 18:38 auto_snap Miscellaneous snapshot
  • 333325 10/01/10 18:37 auto_snap Hall A Tools snapshot
  • 333324 10/01/10 18:37 auto_snap Compton Polarimeter snapshot
  • 333323 10/01/10 18:37 auto_snap Beam Position and Current snapshot
  • 333322 10/01/10 18:23 jroche,hisha Width of discriminated cerenkov signals
  • 333321 10/01/10 17:44 riordan RC Experiment Summary, Oct 1, 2010
  • 333320 10/01/10 17:09 a-onl Start_Run_6900,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=test
  • 333319 10/01/10 17:07 a-onl End_of_Run_6899
  • 333318 10/01/10 17:05 a-onl Start_Run_6893,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Lumi test
  • 333317 10/01/10 17:02 riordan Target Motion Operational
  • 333316 10/01/10 16:51 jpchen solid target can be used for checkout with low current
  • 333315 10/01/10 16:31 a-onl Start_Run_6899,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=bmp test
  • 333314 10/01/10 16:07 adaq Beam Time Accounting
  • 333313 10/01/10 14:16 a-onl End_of_Run_6898
  • 333312 10/01/10 14:14 a-onl Start_Run_6898,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=T1 only
  • 333311 10/01/10 14:13 a-onl End_of_Run_6897
  • 333310 10/01/10 13:51 a-onl Start_Run_6897,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=T5 only
  • 333309 10/01/10 13:50 a-onl End_of_Run_6896
  • 333308 10/01/10 13:30 Techs Cryo Target Pre Beam Checklist
  • 333307 10/01/10 12:31 a-onl Start_Run_6896,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=T5 only
  • 333306 10/01/10 12:30 a-onl End_of_Run_6895
  • 333305 10/01/10 12:10 a-onl Start_Run_6895,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=HRS triggers
  • 333304 10/01/10 12:10 a-onl End_of_Run_6894
  • 333303 10/01/10 11:54 a-onl Start_Run_6894,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=HRS triggers
  • 333302 10/01/10 11:54 a-onl End_of_Run_6893
  • 333301 10/01/10 11:39 hyde bpm in ROC 4
  • 333300 10/01/10 11:32 hyde roc 4 slot 23 gate cable
  • 333299 10/01/10 11:06 a-onl End_of_Run_6892
  • 333298 10/01/10 11:03 Carlos 2nd DVCS HV crate moved to port 6 of dvcstlab1
  • 333297 10/01/10 10:47 adev Start_Run_20012,Run_type=Cosmics,target_type=No Target,targetpol_poldir=Transverse-in-plane,targetpol_pol=,beam_e0=0,beam_raster=on,beam_halfwaveplate=in,rightarm_p=0,rightarm_theta=-70,comment_text=cosmics
  • 333296 10/01/10 10:37 auto_snap Radiation and Ozone snapshot
  • 333295 10/01/10 10:37 auto_snap Miscellaneous snapshot
  • 333294 10/01/10 10:37 auto_snap Hall A Tools snapshot
  • 333293 10/01/10 10:37 auto_snap Compton Polarimeter snapshot
  • 333292 10/01/10 10:37 auto_snap Beam Position and Current snapshot
  • 333291 10/01/10 10:17 M Friend Compton EPICS variable change
  • 333290 10/01/10 10:05 a-onl Start_Run_6892,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=2.261,beam_raster=on,beam_current=2,leftarm_p=1.131,leftarm_theta=5.0,leftarm_sieve=sieve,calo_theta=22,calo_distance=5,comment_text=Lumi test
  • 333289 10/01/10 10:05 Carlos Lumi "pedestal" run 6892
  • 333288 10/01/10 09:22 R. Michaels roc4 missing a gate AGAIN
  • 333287 10/01/10 08:43 adaq Beam Time Accounting
  • 333286 10/01/10 07:00 riordan Hall to Restricted Access
  • 333285 10/01/10 02:37 auto_snap Radiation and Ozone snapshot
  • 333284 10/01/10 02:37 auto_snap Miscellaneous snapshot
  • 333283 10/01/10 02:37 auto_snap Hall A Tools snapshot
  • 333282 10/01/10 02:37 auto_snap Compton Polarimeter snapshot
  • 333281 10/01/10 02:37 auto_snap Beam Position and Current snapshot
  • 333280 10/01/10 01:55 efuchey Compton rates: beam in chicane 2
  • 333279 10/01/10 01:44 adaq Controlled access requested
  • 333278 10/01/10 01:32 efuchey compton rates: beam in chicane
  • 333277 10/01/10 00:48 Carlos Beam established
  • 333276 10/01/10 00:28 Carlos Beam through chicane
  • 333275 10/01/10 00:11 adaq Beam Time Accounting

  • Earlier months at hallaweb.jlab.org/adaq (archive).  The HALOG was generated by version HA1.48
    If you have any comments or problems, please contact:
    Robert Michaels
    rom@jlab.org
    (757) 269-7410